### v1

# Power allocation for cell-free MIMO integrated sensing and communication

Abstract
In this paper, we investigate integrated sensing and communication (ISAC) in a cell-free (CF) multiple-input multiple-output (MIMO) network with single-antenna access points (APs), where each AP functions either as a transmitter for both sensing and communication or as a receiver for target-reflected signals. We derive closed-form Cramér–Rao lower bounds (CRLBs) for location and velocity estimation under arbitrary power allocation ratios, assuming the radar cross-section (RCS) is deterministic and unknown over the observation interval. A power allocation optimization problem is formulated to maximize the communication signal-to-interference-plus-noise ratio (SINR), subject to CRLB-based sensing constraints and per-transmitter power limits. To solve the resulting nonlinear and non-convex problem, we propose a penalty function and projection-based modified conjugate gradient algorithm with inexact line search (PP-MCG-ILS), and an alternative method based on a modified steepest descent approach (PP-MSD-ILS). Additionally, for power minimization in pure sensing scenarios, we introduce a penalty function-based normalized conjugate gradient algorithm (P-NCG-ILS). We analyze the convergence behavior and qualitatively compare the computational complexity of the proposed algorithms. Simulation results confirm the accuracy of the derived CRLBs and demonstrate the effectiveness of the proposed power allocation strategies in enhancing both sensing and overall ISAC performance.
Index Terms:
Cell-free MIMO ISAC, CRLB, Power allocation, Penalty function, Projection operator, Conjugate gradient.
I Introduction
Integrated sensing and communication (ISAC) has emerged as a key enabling technology for sixth-generation (6G) applications, such as smart cities, autonomous driving, and remote healthcare [1, 2]. Unlike traditional systems with separate designs for communication and sensing, ISAC unifies these functionalities within shared hardware platforms and spectral resources, thereby significantly enhancing spectral, energy, and hardware efficiency [3, 4].
Depending on the geometric deployment of transmitters and receivers, various ISAC system architectures have been proposed. In mono-static ISAC systems—where the transmitter and receiver are co-located—advanced beamforming techniques are employed to mitigate self-interference in full-duplex operations [5, 6]. In contrast, cell-free (CF) or multi-static ISAC systems involve widely distributed transmitters and receivers. These systems inherit advantages from both CF communication and distributed multi-static radar due to their architectural similarities [7]. MIMO (multiple-input, multiple-output) technology plays a central role by employing orthogonal waveforms across transmitters to achieve multiplexing and diversity gains for both communication and sensing [8]. Owing to these benefits, CF MIMO-empowered ISAC has been actively explored for unified communication-sensing operations [9, 10, 11, 12, 13, 14].
ISAC systems address a range of sensing tasks, including target detection [15, 16, 17], and parameter estimation for radar cross-section (RCS) [18, 19], location [20, 21, 22], velocity [23, 24], and angle [25, 26]. These capabilities support diverse applications from navigation to environmental monitoring. A fundamental challenge in ISAC system design is the formulation of robust performance metrics, which are critical for resource optimization and balancing communication-sensing trade-offs [3, 4]. While unified metrics based on mutual information [27] and Kullback–Leibler divergence [28] have been proposed, tailoring such metrics to specific sensing objectives remains challenging. Consequently, task-specific metrics for sensing and communication continue to be widely adopted.
For communication, Shannon capacity underpins metrics such as spectral efficiency [29], achievable rate [30], and signal-to-interference-plus-noise ratio (SINR). Other metrics like coverage probability and energy efficiency have also been addressed in scenarios with stringent requirements [31]. On the sensing side, metrics such as sensing mutual information [32, 33], sensing SINR [9, 10], and the Cramér–Rao lower bound (CRLB) [34, 35, 36, 37, 38, 39, 40, 41] have been explored extensively, with CRLB-based designs proving particularly effective.
CRLB-based approaches fall into two categories depending on the RCS model:
Random RCS assumption: RCS values are treated as random variables, and likelihood functions are derived by averaging over their distributions [34, 35, 36, 37]. These studies established CRLBs for joint location and velocity estimation, and later extended them to multi-target scenarios.
Deterministic RCS assumption: RCS values are assumed to be deterministic but unknown and spatially diverse [38, 40, 41, 39]. This model is particularly relevant in slowly varying environments or when prior RCS knowledge is unavailable. Under this setting, CRLBs for both single-target and multiple-target location estimation have been analyzed, with [39] showing convergence to single-target performance as target separations increase.
However, most of the above CRLB formulations rely on continuous-time signal models, while practical radar systems process discrete-time signals, where the sampling rate significantly impacts sensing accuracy. Our prior work [42] introduces a discrete-time likelihood function over Gaussian channels as a more realistic model. Building on this, we derive CRLBs for joint location and velocity estimation under the deterministic RCS assumption in both single- and multiple-target scenarios.
Another critical challenge lies in resource allocation, involving time [43], frequency [44, 45], space [46, 47], and power [9, 10, 12, 11, 13, 14]. While recent works have addressed power allocation for maximizing sensing SINR or minimizing power under CRLB and communication constraints [9, 10, 12, 11, 13], they typically consider localization only, neglecting velocity or angular estimation, and are based on continuous-time models that overlook the impact of sampling rate.
In contrast, this paper focuses on joint location and velocity estimation using discrete-time signal models. Even with simplified CRLBs, the resulting CRLB-constrained power allocation problems are highly nonlinear and non-convex, posing significant challenges to analytical optimization. To address this, we leverage advanced optimization techniques including penalty methods, projection operations, and gradient-based search, building algorithms suitable for such complex constrained problems.
I-A Contribution
This paper addresses the power allocation problem in a CF MIMO-ISAC system with widely distributed single-antenna access points (APs), where each AP operates as either an ISAC transmitter or sensing receiver. We derive CRLBs for joint location and velocity estimation using discrete received signals under the deterministic RCS assumption. These CRLBs are adopted as sensing metrics to guide power allocation under communication SINR maximization objectives. The main contributions are as follows.
1) We derive the approximate CRLBs for joint location and velocity estimation of a single target under arbitrary power allocations. The CRLBs are validated against MLE results, confirming their effectiveness as performance metrics. With the sensing requirements constrained by the CRLBs of both location and velocity estimation, we propose a Penalty-function based Normalized Modified Conjugate Gradient algorithm with Inexact Line Search (P-NCG-ILS) to generate a power allocation scheme for the total power minimization. Simulation results verify its effectiveness.
2) We formulate the CF MIMO-ISAC power allocation problem as a communication SINR maximization task, subject to constraints on both location and velocity CRLBs and total/individual power budgets. To tackle the non-convexity and nonlinearity, we propose a Penalty-function and Projection-based Modified Conjugate Gradient algorithm with Inexact Line Search (PP-MCG-ILS) together with PP-MSD-ILS algorithm based on the modified steepest descent algorithm (MSD).
3) The proposed PA algorithms for sensing and CF MIMO-ISAC can be broadly applicable to constrained non-convex optimization problems involving affine equality constraints (handled via projection) and nonlinear inequality constraints (handled via penalties). We analyze their convergence and computational complexity. Simulation results demonstrate the robust performance, fast convergence, and effectiveness of the proposed algorithms (e.g., PP-MCG-ILS) in optimizing joint sensing and communication performance.
The remainder of the following parts of this paper is organized as follows. Section II describes the received signal model for ISAC. Section IV formulates the power allocation problem for ISAC. Section V details the proposed ISAC power allocation algorithms. Section VII gives the convergence and computational complexity analysis. Section VIII illustrates the simulation results. Section IX concludes this paper.
[…]
II Signal Model
As shown in Fig. 1, we consider a CF MIMO-ISAC system with widely separated APs. Each AP either serves as an ISAC transmitter or a sensing receiver, similar to the setup in [9, 10]. Assume there are $N$ single-antenna transmitters and $K$ single-antenna receivers, with the positions of the transmitter $n$ and receiver $k$ denoted as ${\bm{l}}_{n}=[x_{n},y_{n}]^{\rm T}$ and ${\bm{l}}_{k}=[x_{k},y_{k}]^{\rm T}$, respectively. All transmitters and receivers are scheduled and managed by the edge cloud and the core network, e.g., the fifth generation (5G) communication networks. Denote the location and velocity vectors of the $q$th sensing target as ${\bm{l}}_{q}=[x_{q},y_{q}]^{\rm T}$ and ${\bm{v}}_{q}=[v_{x,q},v_{y,q}]^{\rm T}$, respectively.
Figure 1: CF MIMO system for ISAC.
The Doppler frequency intorduced by the relative radial velocity of target $q$ versus any transmitter $n$ and receiver $k$, is computed as
$\small f_{n,k,q}=\dfrac{1}{\lambda}{\bm{v}}_{q}^{\rm T}\left(\frac{{\bm{l}}_{n}-{\bm{l}}_{q}}{\|{\bm{l}}_{n}-{\bm{l}}_{q}\|_{2}}+\frac{{\bm{l}}_{k}-{\bm{l}}_{q}}{\|{\bm{l}}_{k}-{\bm{l}}_{q}\|_{2}}\right),$ (1)
where $\lambda$ denotes the wavelength. Denote $\tau_{n,k,q}$ as the propagation delay between the $n$th transmitter and the $k$th receiver, as reflected by the $q$th target. Then, we have
$\small\tau_{n,k,q}$ $=\tau_{n,q}+\tau_{k,q},$ (2a)
$\tau_{n,q}$ $={\|{\bm{l}}_{n}-{\bm{l}}_{q}\|_{2}}/{c},$ (2b)
$\tau_{k,q}$ $={\|{\bm{l}}_{k}-{\bm{l}}_{q}\|_{2}}/{c},$ (2c)
where $\tau_{n,q}$ and $\tau_{q,k}$ represent the propagation delays from the $n$th transmitter to the $q$th target, and from the target to the $k$th receiver, respectively, and $c$ denotes the speed of light. In this paper, we consider the single-target scenario, and thus the subscript $q$ can be omitted.
[…]
${\bm{x}}_{n}(t)=\sqrt{P\rho_{n}}{b}_{n}\varsigma s_{n}(t),$ (3)
where $P$ denotes the total available energy, $\rho_{n}$ is the power allocation (PA) factor with $\sum_{n=1}^{N}\rho_{n}=1$, ${b}_{n}$ is the normalized transmit beamforming weight with $|{b}_{n}|^{2}=1$, and $\varsigma$ is the complex data symbol with $\mathbb{E}\{|\varsigma|^{2}\}=1$. Denote ${\bm{\rho}}=[\rho_{1},\rho_{2},\cdots,\rho_{N}]^{\rm T}$ by the vector of the PA coefficient of the $N$ transmitters.
For target-reflected reception, we refer the term “RCS” as encompassing both channel fading and radar RCS. Then, we define the RCS as ${\alpha}_{n,k}$ over transmitter $n$ and receiver $k$. We consider the Swerling-I sensing model, i.e. the deterministic RCS assumption where RCSs are assumed to be deterministic but unknown during the observation time [48, 38, 40, 41, 49]. The lowpass equivalent of the reflected signal from the $n$th transmitter by a target received at receiver $k$ is given by [^1] The transmitted signals over different transmitters are orthogonal and maintain approximate orthogonality over different delays and Doppler frequencies such that the signals contributed from different transmitters can be separated at each receiver [50, 8, 34, 35, 36].
$r_{n,k}(t)=\alpha_{n,k}y_{n,k}(t)+z_{n,k}(t),$ (4)
where $z_{n,k}(t)$ denotes the clutter-plus-noise and $y_{n,k}(t)$ is the loss-free and noise-free received signal,
${y}_{n,k}(t)=\sqrt{P\rho_{n}}b_{n}s_{n}(t-\tau_{n,k})e^{j2\pi f_{n,k}t}.$ (5)
Note that clutters can initially be mitigated by Doppler processing or space-time adaptive processing [51] and whitened through a whitening filter [34]. It is reasonable to assume the clutter-plus-noise to be additive white Gaussian noise with zero mean and variance $\sigma_{\rm cn}^{2}$ [8, 52, 53, 34, 35, 36]. Define $\delta_{\rm cn}=\frac{P}{T_{\rm eff}\sigma_{\rm cn}^{2}}$ as the signal-to-clutter-plus-noise ratio (SCNR).
The received signal of the communication user is given by,
${r}_{u}(t)$ $=\sqrt{P}\sum_{n=1}^{N}g_{n}{b}_{n}\sqrt{\rho_{n}}s_{n}(t)+{z}(t),$ (6)
where $g_{n}$ is the channel fading between the transmitter $n$ and the user, and the interference-plus-noise ${z}(t)$ is assumed to be the additive circular Gaussian white with zero mean and variance $\sigma_{z}^{2}$. Then, the SINR at the user terminal can be given by
$\gamma$ $=\sum_{n=1}^{N}|g_{n}{b}_{n}|^{2}\rho_{n}\delta\overset{(\text{i})}{=}\delta{\bm{\rho}}^{\rm T}{\bm{g}}$ (7)
where $\delta=\frac{P}{T_{\rm eff}\sigma_{z}^{2}}$ is the total SINR and ${\bm{g}}$ denotes the channel gain [^2] The communication channels are assumed to be estimated a priori using pilot signals [9, 10, 11, 12, 13]. with its $n$th element being $|g_{n}|^{2}$. Note that $(\mathrm{i})$ holds when using a normalized conjugate beamforming weight with ${b}_{n}={g}^{*}_{n}/|g_{n}|$.
III Sensing Metric
In this section, we present an approximate CRLB for CF MIMO radar under arbitrary power allocations.
First define the squared effective bandwidth (SEBW) and squared effective pulse time width (SETW) as
[…]
and
[…]
respectively. The cross-term between time and frequency is further defined as
[…]
We further define the average frequency as
[…]
and average time as
[…]
With a relatively large bandwidth and pulse time width, such that $\bar{f}_{n}^{2}\ll\bar{f_{n}^{2}}$ and $(\bar{t}_{n}+\tau_{n,k})^{2}\ll\bar{t_{n}^{2}}$, we have derived the approximate CRLBs for both location and velocity [42], respectively given by
${\bm{C}}_{\rm L}$ $=({\bm{P}}-{\bm{V}}{\bm{Y}}^{-1}{\bm{V}}^{\rm T})^{-1},$ (13)
and
${\bm{C}}_{\rm V}$ $=({\bm{Y}}-{\bm{V}}^{\rm T}{\bm{P}}^{-1}{\bm{V}})^{-1}.$ (14)
The corresponding $2\times 2$ sub-matrix ${\bm{P}}$ is given in (15), ${\bm{V}}$ given in (16),
${\bm{P}}=\sum_{n=1}^{N}\sum_{k=1}^{K}\begin{bmatrix}\beta_{n,k}^{2}a_{n,k}+2\beta_{n,k}\eta_{n,k}b_{n,k}+\eta_{n,k}^{2}d_{n,k}&\beta_{n,k}\zeta_{n,k}a_{n,k}+(\beta_{n,k}\kappa_{n,k}+\zeta_{n,k}\eta_{n,k})b_{n,k}+\eta_{n,k}\kappa_{n,k}d_{n,k}\\ \zeta_{n,k}\beta_{n,k}a_{n,k}+(\zeta_{n,k}\eta_{n,k}+\beta_{n,k}\kappa_{n,k})b_{n,k}+\kappa_{n,k}\eta_{n,k}d_{n,k}&\zeta_{n,k}^{2}a_{n,k}+2\zeta_{n,k}\kappa_{n,k}b_{n,k}+\kappa_{n,k}^{2}d_{n,k}\end{bmatrix},$ (15)
${\bm{V}}=\sum_{n=1}^{N}\sum_{k=1}^{K}\begin{bmatrix}\beta_{n,k}\xi_{n,k}b_{n,k}+\eta_{n,k}\xi_{n,k}d_{n,k}&\beta_{n,k}\varrho_{n,k}b_{n,k}+\eta_{n,k}\varrho_{n,k}d_{n,k}\\ \zeta_{n,k}\xi_{n,k}b_{n,k}+\kappa_{n,k}\xi_{n,k}d_{n,k}&\zeta_{n,k}\varrho_{n,k}b_{n,k}+\kappa_{n,k}\varrho_{n,k}d_{n,k}\end{bmatrix},$ (16)
and
${\bm{Y}}=\sum_{n=1}^{N}\sum_{k=1}^{K}\begin{bmatrix}\xi_{n,k}^{2}d_{n,k}&\xi_{n,k}\varrho_{n,k}d_{n,k}\\ \varrho_{n,k}\xi_{n,k}d_{n,k}&\varrho_{n,k}^{2}d_{n,k}\end{bmatrix}.$ (17)
The involved waveform-related parameters, e.g., $a_{n,k}$, $b_{n,k}$ and $d_{n,k}$, are respectively given by,
$a_{n,k}$ $=\frac{8\pi^{2}|\alpha_{n,k}|^{2}f_{s}P\rho_{n}\bar{f_{n}^{2}}}{\sigma_{z}^{2}},$ (18a)
$b_{n,k}$ $=\frac{4\pi|\alpha_{n,k}|^{2}f_{s}P\rho_{n}}{\sigma_{z}^{2}}\mathfrak{T}\{\sigma_{tf}\},$ (18b)
$d_{n,k}$ $\approx\frac{8\pi^{2}|\alpha_{n,k}|^{2}f_{s}P\rho_{n}\bar{t_{n}^{2}}}{\sigma_{z}^{2}},$ (18c)
where the approximation holds in (18c) when the effective time width meets $\tau_{n,k}\ll T_{\rm eff}$. The relevant geometric spread parameters, including $\beta_{n,k}$, $\zeta_{n,k}$, $\xi_{n,k}$, $\varrho_{n,k}$, $\eta_{n,k}$ and $\kappa_{n,k}$, are omitted here and can be referred to [34, 42].
To visualize the relationship between the CRLBs and the PA coefficients, we now reformulate the CRLB expressions. Given the equivalent location FIM Let $p_{ij}$ denote the $\{i,j\}$th element of matrix ${\bm{P}}$, similar for $v_{ij}$ and $y_{ij}$. The elements $p_{ij}$, $v_{ij}$ and $y_{ij}$ can be expressed as the weighted sum of $\rho_{n}$, i.e.,
$\begin{cases}p_{ij}=\sum_{n=1}^{N}w_{n}^{p_{ij}}\rho_{n}={\bm{\rho}}^{\rm T}{\bm{w}}_{p_{ij}},\\ v_{ij}=\sum_{n=1}^{N}w_{n}^{v_{ij}}\rho_{n}={\bm{\rho}}^{\rm T}{\bm{w}}_{v_{ij}},\\ y_{ij}=\sum_{n=1}^{N}w_{n}^{y_{ij}}\rho_{n}={\bm{\rho}}^{\rm T}{\bm{w}}_{y_{ij}},\end{cases}$ (19)
where $w_{n}^{p_{ij}}$ is the $n$th element of the weight vector ${\bm{w}}_{p_{ij}}$, similar for $w_{n}^{v_{ij}}$ and $w_{n}^{y_{ij}}$. These weights can be readily obtained from (15)-(17) by extracting the PA vector ${\bm{\rho}}$.
By substituting (19) into (13), the CRLB matrix for location estimation can be rewritten as,
${\bm{C}}_{\rm L}=\frac{1}{|{\bm{U}}|}\begin{bmatrix}u_{22}&-u_{12}\\ -u_{21}&u_{11}\end{bmatrix},$ (20)
where $|{\bm{U}}|=u_{11}u_{22}-u_{12}u_{21}$ and
$u_{11}=\frac{1}{|{\bm{Y}}|}\big{[}{\bm{\rho}}^{\rm T}({\bm{w}}_{p_{11}}{\bm{\rho}}^{\rm T}{\bm{W}}_{y}-{\bm{w}}_{v_{11}}{\bm{\rho}}^{\rm T}{\bm{W}}_{12}^{y}-{\bm{w}}_{v_{12}}{\bm{\rho}}^{\rm T}\breve{{\bm{W}}}_{11}^{y}){\bm{\rho}}\big{]},$ (21a)
$u_{12}=\frac{1}{|{\bm{Y}}|}\big{[}{\bm{\rho}}^{\rm T}({\bm{w}}_{p_{12}}{\bm{\rho}}^{\rm T}{\bm{W}}_{y}-{\bm{w}}_{v_{21}}{\bm{\rho}}^{\rm T}{\bm{W}}_{12}^{y}-{\bm{w}}_{v_{22}}{\bm{\rho}}^{\rm T}\breve{{\bm{W}}}_{11}^{y}){\bm{\rho}}\big{]},$ (21b)
$u_{22}=\frac{1}{|{\bm{Y}}|}\big{[}{\bm{\rho}}^{\rm T}({\bm{w}}_{p_{22}}{\bm{\rho}}^{\rm T}{\bm{W}}_{y}-{\bm{w}}_{v_{21}}{\bm{\rho}}^{\rm T}{\bm{W}}_{22}^{y}-{\bm{w}}_{v_{22}}{\bm{\rho}}^{\rm T}\breve{{\bm{W}}}_{21}^{y}){\bm{\rho}}\big{]},$ (21c)
where $|{\bm{Y}}|={\bm{\rho}}^{\rm T}{\bm{W}}_{y}{\bm{\rho}}$. Elements $u_{ij}$ are respectively given by,
$\begin{cases}{\bm{W}}_{y}={\bm{w}}_{y_{11}}{\bm{w}}_{y_{22}}^{\rm T}-{\bm{w}}_{y_{12}}{\bm{w}}_{y_{21}}^{\rm T},\\ {\bm{W}}^{y}_{12}={\bm{w}}_{v_{11}}{\bm{w}}_{y_{22}}^{\rm T}-{\bm{w}}_{v_{12}}{\bm{w}}_{y_{21}}^{\rm T},\\ {\bm{W}}^{y}_{22}={\bm{w}}_{v_{21}}{\bm{w}}_{y_{22}}^{\rm T}-{\bm{w}}_{v_{22}}{\bm{w}}_{y_{21}}^{\rm T},\\ \breve{\bm{W}}^{y}_{11}={\bm{w}}_{v_{12}}{\bm{w}}_{y_{11}}^{\rm T}-{\bm{w}}_{v_{11}}{\bm{w}}_{y_{12}}^{\rm T},\\ \breve{\bm{W}}^{y}_{21}={\bm{w}}_{v_{22}}{\bm{w}}_{y_{11}}^{\rm T}-{\bm{w}}_{v_{21}}{\bm{w}}_{y_{12}}^{\rm T}.\end{cases}$ (22)
Similarly, the velocity CRLB matrix can be rewritten as,
${\bm{C}}_{\rm V}=\frac{1}{|{\bm{H}}|}\begin{bmatrix}h_{22}&-h_{12}\\ -h_{21}&h_{11}\end{bmatrix},$ (23)
where $|{\bm{H}}|=h_{11}h_{22}-h_{12}h_{21}$. Elements $h_{ij}$ are respectively given by,
$h_{11}=\frac{1}{|{\bm{P}}|}\big{[}{\bm{\rho}}^{\rm T}({\bm{w}}_{y_{11}}{\bm{\rho}}^{\rm T}{\bm{W}}_{p}-{\bm{w}}_{v_{11}}{\bm{\rho}}^{\rm T}{\bm{W}}_{12}^{p}-{\bm{w}}_{v_{21}}{\bm{\rho}}^{\rm T}\breve{{\bm{W}}}_{11}^{p}){\bm{\rho}}\big{]},$ (24a)
$h_{12}=\frac{1}{|{\bm{P}}|}\big{[}{\bm{\rho}}^{\rm T}({\bm{w}}_{y_{12}}{\bm{\rho}}^{\rm T}{\bm{W}}_{p}-{\bm{w}}_{v_{12}}{\bm{\rho}}^{\rm T}{\bm{W}}_{12}^{p}-{\bm{w}}_{v_{22}}{\bm{\rho}}^{\rm T}\breve{{\bm{W}}}_{11}^{p}){\bm{\rho}}\big{]},$ (24b)
$h_{22}=\frac{1}{|{\bm{P}}|}\big{[}{\bm{\rho}}^{\rm T}({\bm{w}}_{y_{22}}{\bm{\rho}}^{\rm T}{\bm{W}}_{p}-{\bm{w}}_{v_{12}}{\bm{\rho}}^{\rm T}{\bm{W}}_{22}^{p}-{\bm{w}}_{v_{22}}{\bm{\rho}}^{\rm T}\breve{{\bm{W}}}_{21}^{p}){\bm{\rho}}\big{]},$ (24c)
where $|{\bm{P}}|={\bm{\rho}}^{\rm T}{\bm{W}}_{p}{\bm{\rho}}$, and
$\begin{cases}{\bm{W}}_{p}={\bm{w}}_{p_{11}}{\bm{w}}_{p_{22}}^{\rm T}-{\bm{w}}_{p_{12}}{\bm{w}}_{p_{21}}^{\rm T},\\ {\bm{W}}^{p}_{12}={\bm{w}}_{v_{11}}{\bm{w}}_{p_{22}}^{\rm T}-{\bm{w}}_{v_{21}}{\bm{w}}_{p_{21}}^{\rm T},\\ {\bm{W}}^{p}_{22}={\bm{w}}_{v_{12}}{\bm{w}}_{p_{22}}^{\rm T}-{\bm{w}}_{v_{22}}{\bm{w}}_{p_{21}}^{\rm T},\\ \breve{\bm{W}}^{p}_{11}={\bm{w}}_{v_{21}}{\bm{w}}_{p_{11}}^{\rm T}-{\bm{w}}_{v_{11}}{\bm{w}}_{p_{12}}^{\rm T},\\ \breve{\bm{W}}^{p}_{21}={\bm{w}}_{v_{22}}{\bm{w}}_{p_{11}}^{\rm T}-{\bm{w}}_{v_{12}}{\bm{w}}_{p_{12}}^{\rm T}.\end{cases}$ (25)
IV Problem Formulation
In general, a higher power budget leads to improved sensing and communication performance. This section investigates a novel PA scheme aimed at maximizing the overall performance of CF MIMO-ISAC systems. Depending on specific application requirements, the optimization objective can be either sensing or communication performance, while the quality of service (QoS) requirement of the other function is incorporated as a constraint in the optimization model.
Let ${\bm{\rho}}_{\min}=[\rho_{\min,1},\rho_{\min,2},\cdots,\rho_{\min,N}]^{\rm T}$ denote the minimum PA coefficient to satisfy the so-called larger SNR for tight CRLBs, and ${\bm{\rho}}_{\max}=[\rho_{\max,1},\rho_{\max,2},\cdots,\rho_{\max,N}]^{\rm T}$ denote the maximum PA coefficient due to the power budgets of individual transmitters [40, 13]. In this paper, we consider maximizing the communication data rate or SINR under the constraint of the sensing performance, i.e.,
$\arg\min\limits_{{\bm{\rho}}}$ $\ -{\bm{\rho}}^{\rm T}{\bm{g}},$ (26a)
${\rm s.t.}$ ${\bm{\rho}}^{\rm T}{\bm{1}}=1,$ (26b)
${\bm{\rho}}_{\min}\leq{\bm{\rho}}\leq{\bm{\rho}}_{\max},$ (26c)
${\rm tr}\{{\bm{C}}_{\rm L}\}\leq\delta_{l}^{2},$ (26d)
${\rm tr}\{{\bm{C}}_{\rm V}\}\leq\delta_{v}^{2},$ (26e)
where $\delta_{l}^{2}$ and $\delta_{v}^{2}$ denote the corresponding performance thresholds  [^3] The thresholds can be set by referring to the CRLBs under a uniform power allocation..
Similar to the approaches in [40, 49, 11, 13], we assume that the target’s RCS, location, and velocity have been coarsely estimated by the ISAC system and are periodically updated over time [^4] The previously estimated target parameters can be reused for PA optimization at the current time step, particularly in consecutive estimation or target tracking scenarios.. Therefore, the PA coefficient vector $\bm{\rho}$ becomes the sole optimization variable, while the other relevant parameters are implicitly embedded through their coupling in equations (22) and (25).
V Algorithm Derivation
In this section, we investigate the optimal PA strategy for maximizing the ISAC performance, as formulated in (57a)-(26e). However, the strong nonlinearity and non-convexity of the CRLB-based constraints in (26d) and (26e) render the problem analytically intractable. To overcome this challenge, we adopt a conjugate gradient (CG) iterative algorithm, enhanced with a penalty function and projection operator, to provide an efficient numerical solution.
We first introduce a penalty (function) for violating the nonlinear inequality constraints in (26c)-(26e)
$\alpha({\bm{\rho}})$ $=\tilde{f}_{l}({\bm{\rho}})+\tilde{f}_{v}({\bm{\rho}})+\sum_{n=1}^{N}\left(f_{\min,n}({\bm{\rho}})+f_{\max,n}({\bm{\rho}})\right),$ (27)
where
$\begin{cases}\tilde{f}_{l}({\bm{\rho}})=(\max\{0,{f}_{l}({\bm{\rho}})\})^{q},\\ \tilde{f}_{v}({\bm{\rho}})=(\max\{0,{f}_{v}({\bm{\rho}})\})^{q},\\ f_{\min,n}({\bm{\rho}})=(\max\{0,\rho_{\min,n}-\rho_{n}\})^{q},\\ f_{\max,n}({\bm{\rho}})=(\max\{0,\rho_{n}-\rho_{\max,n}\})^{q},\end{cases}$ (28)
with ${f}_{l}({\bm{\rho}})\triangleq{\rm tr}\{{\bm{C}}_{\rm L}\}-\delta_{l}^{2}$ and ${f}_{v}({\bm{\rho}})\triangleq{\rm tr}\{{\bm{C}}_{\rm V}\}-\delta_{v}^{2}$. We consider $q=2$ to guarantee the continuous differentiable penalty functions. Then, the objective function for the penalty problem can be written as
$L({\bm{\rho}})=-{\bm{\rho}}^{\rm T}{\bm{g}}+\mu\alpha({\bm{\rho}}).$ (29)
As mentioned in [54], as the penalty factor $\mu$ is made large, the points generated for minimizing (29) approach an optimal solution of the original objective function (57a) from outside the feasible region and the iteration process terminate with $\mu\alpha({\bm{\rho}})\leq\epsilon_{\mu}$ for a small threshold $\epsilon_{\mu}>0$. This issue can be solved by setting $\epsilon_{\mu}$ small enough or employing much tighter thresholds than the required bounds to some extent to mitigate the loss due to the non-zero threshold $\epsilon_{\mu}$.
On the other hand, intuitively, when orthogonal waveforms are employed, increasing the transmission power—or equivalently, the signal-to-noise ratio (SNR)—at any individual transmitter enhances both sensing and communication performance. Therefore, it is desirable to fully utilize the total power budget, as specified in (26b). To enforce this equality constraint and reduce both computational complexity and the nonlinearity of the optimization problem, we adopt a projection operator instead of incorporating another penalty sub-function
${\bm{\rho}}({\bm{\varpi}})={\bm{\Theta}}{\bm{\varpi}}+{\bm{\rho}}_{0},$ (30)
where ${\bm{\Theta}}\triangleq{\bm{I}}_{N\times N}-\frac{{\bm{1}}_{N}{\bm{1}}_{N}^{\rm T}}{N}$ is a projection matrix onto the zero space of ${\bm{1}}_{N}$. It is evident that ${\bm{1}}_{N}^{\rm T}{\bm{\rho}}({\bm{\varpi}})=1$ holds with ${\bm{1}}_{N}^{\rm T}{\bm{\rho}}_{0}=1$. Thus, ${\bm{\rho}}({\bm{\varpi}})$ is the sum of a constant PA vector ${\bm{\rho}}_{0}$ and a projection of any $N\times 1$ vector ${\bm{\varpi}}$ onto the zero space of ${\bm{1}}_{N}$. According to (30), the objective function can be further written as,
$\mathcal{L}({\bm{\varpi}})=-{\bm{\rho}}^{\rm T}({\bm{\varpi}}){\bm{g}}+\mu\alpha({\bm{\rho}}({\bm{\varpi}})).$ (31)
This objective function makes all the constraints satisfied and thus convert the original constrained optimization regarding ${\bm{\rho}}$ into an unconstrained one regarding ${\bm{\varpi}}$.
V-A Modified Conjugate Gradient Algorithm
We now propose a projection based modified CG (MCG) algorithm to minimize (31). Taking the derivative $\mathcal{L}({\bm{\varpi}})$ regarding ${\bm{\varpi}}$ yields
(ϖ)=bold-italic-$\nabla\mathcal{L}({\bm{\varpi}})=$ ${\bm{\Theta}}\nabla L({\bm{\rho}})={\bm{\Theta}}\big{[}{\bm{g}}+\nabla\alpha({\bm{\rho}})\big{]}.$ (32)
where
$\nabla\alpha({\bm{\rho}})=$ $\mu\big{(}\nabla\tilde{f}_{l}({\bm{\rho}})+\nabla\tilde{f}_{v}({\bm{\rho}})$
$+\sum_{n=1}^{N}(\nabla f_{\min,n}({\bm{\rho}})+\nabla f_{\max,n}({\bm{\rho}}))\big{)},$ (33)
with components
$\begin{cases}\nabla\tilde{f}_{l}({\bm{\rho}})=2\max\{f_{l}({\bm{\rho}}),0\}\nabla f_{l}({\bm{\rho}})\\ \nabla\tilde{f}_{v}({\bm{\rho}})=2\max\{f_{v}({\bm{\rho}}),0\}\nabla f_{v}({\bm{\rho}})\\ \nabla f_{\min,n}({\bm{\rho}})=2\min\{\rho_{n}-\rho_{\min,n},0\}{\bm{e}}_{n}\\ \nabla f_{\max,n}({\bm{\rho}})=2\max\{\rho_{n}-\rho_{\max,n},0\}{\bm{e}}_{n}.\end{cases}$ (34)
Note that ${\bm{e}}_{n}$ is an index vector with its element at index $n$ being $1$ and the others being zeros.
Solving (31) using an iterative method yields the update
${\bm{\varpi}}_{i+1}={\bm{\varpi}}_{i}+\upsilon_{i}{\bm{d}}_{i},$ (35)
where ${\bm{d}}_{i}$ is the descent direction of $\mathcal{L}({\bm{\varpi}}_{i})$ and $\upsilon_{i}$ is the step size. The descent direction using the MCG is updated by
${\bm{d}}_{i+1}$ $=-{\bm{\Theta}}\nabla L({\bm{\rho}}_{i+1})+\varsigma_{i}{\bm{d}}_{i},$ (36)
where $\varsigma_{i}$ denotes the deflection factor that characterizes a particular CG algorithm. Note that $-{\bm{\Theta}}\nabla L({\bm{\rho}}_{i})$ is still the descent direction of $L({\bm{\rho}}_{i})$ with $-\nabla L({\bm{\rho}}_{i})^{\rm T}{\bm{\Theta}}\nabla L({\bm{\rho}}_{i})\leq 0$ since the projection operator ${\bm{\Theta}}$ is a symmetric positive definite matrix.
Given the initial direction ${\bm{d}}_{1}=-{\bm{\Theta}}\nabla L({\bm{\rho}}_{1})$, and using the idempotent property of the projection matrix, i.e., ${\bm{\Theta}}^{2}={\bm{\Theta}}$, we have ${\bm{\Theta}}{\bm{d}}_{i}={\bm{d}}_{i}$, $\forall{i>0}$. Then, taking the projection of (35) leads to
${\bm{\rho}}_{i+1}={\bm{\rho}}_{i}+\upsilon_{i}{\bm{d}}_{i}.$ (37)
which has the same form as (35). Therefore, we hereafter directly adopt the update in (37). Moreover, note that the equality constraint ${\bm{1}}_{N}^{\rm T}{\bm{\rho}}_{i}=1$ holds for all $i>1$ if and only if ${\bm{1}}_{N}^{\rm T}{\bm{\rho}}_{1}=1$. This property follows from the iteration rules in (36) and (37).
To enable the convergence, it is necessary to develop a suitable step size $\upsilon_{i}$ and search direction ${\bm{d}}_{i}$ in each iteration to guarantee the non-increasing objective function i.e.,
$L({\bm{\rho}}_{i}+\upsilon_{i}{\bm{d}}_{i})-L({\bm{\rho}}_{i})\leq 0.$ (38)
Firstly, for the non-increasing search direction for each iteration $i$, it requires
${\bm{d}}_{i+1}^{\rm T}\nabla L({\bm{\rho}}_{i+1})\leq 0,$ (39)
where the equality holds if and only if $\nabla L({\bm{\rho}}_{i+1})={\bm{0}}$. Using (36), we have,
$\begin{cases}\varsigma_{i}\leq\varsigma_{i}^{\rm b},\ \varsigma_{i}^{\rm b}>0\\ \varsigma_{i}\geq\varsigma_{i}^{\rm b},\ \varsigma_{i}^{\rm b}<0\end{cases}$ (40)
where $\varsigma_{i}^{\rm b}$ is the bound for $\varsigma_{i}$
$\varsigma_{i}^{\rm b}=\dfrac{\nabla L({\bm{\rho}}_{i+1})^{\rm T}{\bm{\Theta}}\nabla L({\bm{\rho}}_{i+1})}{{\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i+1})}.$ (41)
Note that $\nabla L({\bm{\rho}}_{i+1})^{\rm T}{\bm{\Theta}}\nabla L({\bm{\rho}}_{i+1})>0$ for any non-zero gradient $\nabla L({\bm{\rho}}_{i+1})$ due to the positive definite ${\bm{\Theta}}$.
When considering inexact line searches, Hestenes and Stiefel’s [55, 56] choice for $\varsigma_{i}$ can be preferable. We have the modified HS deflection,
$\varsigma_{i}^{\rm HS}=\dfrac{\nabla L({\bm{\rho}}_{i+1})^{\rm T}{\bm{q}}_{i}}{{\bm{d}}_{i}^{\rm T}{\bm{q}}_{i}},$ (42)
which is obtained from ${\bm{d}}_{i+1}^{\rm T}{\bm{q}}_{i}=0$ with ${\bm{q}}_{i}\triangleq{\bm{\Theta}}[\nabla L({\bm{\rho}}_{i+1})-\nabla L({\bm{\rho}}_{i})]$. In fact, when using the MCG algorithm, the orthogonality is desired between the (conjugate) gradients in adjacent iterations, i.e.,
$|{\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i+1})|\ll|{\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i})|,$ (43)
$|\nabla L({\bm{\rho}}_{i+1})^{\rm T}{\bm{\Theta}}\nabla L({\bm{\rho}}_{i})|\ll|\nabla L({\bm{\rho}}_{i+1})^{\rm T}{\bm{\Theta}}\nabla L({\bm{\rho}}_{i+1})|.$ (44)
Equations (43) and (44) ensure the positivity of both ${\bm{d}}_{i}^{\rm T}{\bm{q}}i$ and $\nabla L({\bm{\rho}}{i+1})^{\rm T}{\bm{q}}_{i}$, respectively. This allow for a positive conjugate gradient scaling factor $\varsigma_{i}^{\rm HS}$. Furthermore, (43) and (44) also imply that $|{\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i+1})|\ll|{\bm{d}}_{i}^{\rm T}{\bm{q}}_{i}|$ and $|\nabla L({\bm{\rho}}_{i+1})^{\rm T}{\bm{q}}_{i}|\approx|\nabla L({\bm{\rho}}_{i+1})^{\rm T}{\bm{\Theta}}\nabla L({\bm{\rho}}_{i+1})|$, respectively. As a result, $\varsigma_{i}^{\rm HS}$ is expected to be smaller than $|\varsigma_{i}^{\rm b}|$, thereby satisfying the deflection constraint defined in (40).
Rather than strictly relying on the potentially tight and unquantifiable conjugacy conditions in (43) and (44), we instead enforce the positivity of both the numerator and denominator in (42), along with the inequality $\varsigma_{i}^{\rm HS}<|\varsigma_{i}^{\rm b}|$. When these conditions hold, we set $\varsigma_{i}=\varsigma_{i}^{\rm HS}$. Otherwise, we reset the MCG iteration by letting $\varsigma_{i}=0$. To sum up, the deflection factor is updated by
${\varsigma_{i}=}\begin{cases}\varsigma_{i}^{\rm HS},\ \min\{\nabla L({\bm{\rho}}_{i+1})^{\rm T}{\bm{q}}_{i},{\bm{d}}_{i}^{\rm T}{\bm{q}}_{i},|\varsigma_{i}^{\rm b}|-\varsigma_{i}^{\rm HS}\}>0,\\ 0,\ {\rm else}.\end{cases}$ (45)
Note that, after every consecutive $N$ MCG steps with $\varsigma_{i}=\varsigma_{i}^{\rm HS}$, we will also let $\varsigma_{i}=0$ to restart the new loop of MCG iteration. A proper reboot is potential in mitigating the accumulation of numerical errors, restoring the orthogonality and conjugacy of direction vectors, and avoiding local minima in nonlinear and non-convex optimization.
To optimize the step size selection, we define $f_{i}(\upsilon_{i})\triangleq L({\bm{\rho}}_{i}+\upsilon_{i}{\bm{d}}_{i})$. Then, we have $f_{i}(0)=L({\bm{\rho}}_{i})$. The second-order Taylor approximation of $f_{i}(\upsilon_{i})$ at $\upsilon_{i}=0$ is given by
$f_{i}(\upsilon_{i})$ $\thickapprox f_{i}(0)+\upsilon_{i}\nabla f_{i}(0)+\upsilon_{i}^{2}\nabla^{2}f_{i}(0)/2$
$=f_{i}(0)+\upsilon_{i}{\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i})+\upsilon_{i}^{2}/2{\bm{d}}_{i}^{\rm T}\nabla^{2}L({\bm{\rho}}_{i}){\bm{d}}_{i},$ (46)
where $\nabla^{2}L({\bm{\rho}}_{i})$ is the Hessian matrix of $L({\bm{\rho}})$ at ${\bm{\rho}}_{i}$. According to (39), we have ${\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i})\leq 0$. If $\|\nabla^{2}L({\bm{\rho}}_{i})\|$ is locally bounded, there exists a non-negative step size bound $\upsilon_{b}$ for given $\epsilon_{L}\in[0,1)$ such that, when $\upsilon_{i}\leq\upsilon_{b}$, we have
$f_{i}(\upsilon_{i})-f_{i}(0)\leq\epsilon_{L}\upsilon_{i}{\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i}).$ (47)
When ${\bm{d}}_{i}^{\rm T}\nabla^{2}L({\bm{\rho}}_{i}){\bm{d}}_{i}\leq 0$, (47) always holds. In contrast, we have $\upsilon_{b}=\dfrac{2(\epsilon_{L}-1){\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i})}{{\bm{d}}_{i}^{\rm T}\nabla^{2}L({\bm{\rho}}_{i}){\bm{d}}_{i}}$ when ${\bm{d}}_{i}^{\rm T}\nabla^{2}L({\bm{\rho}}_{i}){\bm{d}}_{i}>0$.
Due to the inherent high computational complexity of the Hessian matrix $\nabla^{2}L({\bm{\rho}}_{i})$ in calaulating $\upsilon_{b}$, we can apply an inexact line search (ILS) for the step size to enable
$L({\bm{\rho}}_{i}+\upsilon_{i}{\bm{d}}_{i})-L({\bm{\rho}}_{i})\leq\epsilon_{L}\upsilon_{i}{\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i}).$ (48)
According to (39), the term on the right-hand side of (48) is always non-positive. Thus, (48) contributes to (38). The ILS controlling factor $\epsilon_{L}$ is used to control the decrement of the function $L({\bm{\rho}})$ by moving a step $\upsilon_{i}$. When (48) is unmet for given $\upsilon_{i}$, it is scaled by,
$\upsilon_{i}=\sigma_{\upsilon}\upsilon_{i}$ (49)
where the scaling factor $\sigma_{\upsilon}\in(0,1)$ and an initial step size is denoted as $\upsilon_{i}=\upsilon_{0,i}$.
On the other hand, the penalty function does not inherently guarantee that the individual power budget constraints at each transmitter are satisfied during the optimization process. As a result, negative PA values may emerge during intermediate iterations. These infeasible PA values can lead to misleading CRLB evaluations, such as yielding negative CRLBs instead of infinite ones, thereby artificially satisfying the CRLB constraints. To prevent unnecessary iterations within the infeasible region and to ensure physically meaningful results, it is essential to enforce non-negative PAs at each iteration step. To this end, the step size $\upsilon_{i}$ needs to satisfy,
$0<\upsilon_{i}\leq\upsilon_{i}^{\rm b}$ (50)
where $\upsilon_{i}^{\rm b}$ is a step size upper bound given by
$\upsilon_{i}^{\rm b}=\min_{d_{n,i}<0}\{\rho_{n,i}/|{d}_{n,i}|\},$ (51)
with $\rho_{n,i}$ denoting the $n$th element of ${\bm{\rho}}_{i}$ and similar for ${d}_{n,i}$. Thus, the initial step size for ILS can be selected as,
$\upsilon_{0,i}=\epsilon_{\upsilon}\upsilon_{i}^{\rm b}$ (52)
with $\epsilon_{\upsilon}\in(0,1]$. Note that there must exist ${d}_{n,i}<0$ at each iteration $i$ unless ${\bm{d}}_{i}={\bm{0}}$ due to $\sum_{n=1}^{N}{d}_{n,i}=0$. The ILS method characterized in (48)-(52) can be seen as the extension and application of the Armijo’s Rule [57, 56] in our work.
A termination condition is significant for the iterative algorithm. A natural choice, based on the ILS condition in (48), is that the absolute difference in objective function values between consecutive iterations satisfies $|L({\bm{\rho}}_{i+1})-L({\bm{\rho}}_{i})|\leq\epsilon$ where $\epsilon$ is a small positive threshold. To avoid numerical instability arising from the magnitude of the objective function itself, we apply a normalized termination condition, i.e., $|L({\bm{\rho}}_{i+1})-L({\bm{\rho}}_{i})|/|L({\bm{\rho}}_{i+1})|\leq\epsilon_{\rm th}$, as employed in both the MCG-based and MSD-based algorithms.
We now present the Penalty-function and Projection-based Modified Conjugate Gradient algorithm with inexact line search (PP-MCG-ILS). A large penalty factor may lead to computational issues, such as an ill-conditioned Hessian matrix, which can adversely affect the convergence of the MCG algorithm. To mitigate this, we adopt a strategy of progressively increasing the penalty factor across iterations. For given penalty factor $\mu_{k}$, the MCG algorithm is executed until its termination condition is satisfied. Then, the penalty factor is updated as $\mu_{k+1}=\varphi\mu_{k}$ for the next loop of penalty-based iteration, where $\varphi>1$ is a scaling constant. This process continues until the convergence criterion $\mu_{k}\alpha(\bm{\rho})\leq\epsilon_{\mu}$ is met, with $\epsilon_{\mu}$ being a small, user-defined threshold [54]. This approach ensures numerical stability and gradual enforcement of the constraint. The complete PP-MCG-ILS algorithm is summarized in Algorithm 1.
Algorithm 1 The PP-MCG-ILS algorithm
Input: The positive scalar termination factor $\epsilon_{\mu}$ and $\epsilon_{\rm th}$, small positive constant $\sigma_{\upsilon}$, $\epsilon_{L}$ and $\epsilon_{\upsilon}$, initial penalty factor $\mu_{1}$, scaling factor $\varphi>1$; initial PA parameter ${\bm{\rho}}_{1}$ satisfying ${\bm{1}}^{\rm T}{\bm{\rho}}_{1}=1$, projection operator ${\bm{\Theta}}$, ${\bm{y}}_{1}={\bm{\rho}}_{1}$, and ${\bm{d}}_{1}=-{\bm{\Theta}}\nabla L({\bm{\rho}}_{1})$; a counter $k=0$ and $i=j=1$.
Output: The PA vector ${\bm{y}}_{j}$.
1:  Step size: Calculate $\upsilon_{i}^{\rm b}$ according to (51); set $\upsilon_{i}=\epsilon_{\upsilon}\upsilon_{i}^{\rm b}$ , while $L({\bm{\rho}}_{i}+\upsilon_{i}{\bm{d}}_{i})-L({\bm{\rho}}_{i})>\epsilon_{L}\upsilon_{i}{\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i})$, let $\upsilon_{i}=\sigma_{\upsilon}\upsilon_{i}$.
2:  Parameter update: ${\bm{\rho}}_{i+1}={\bm{\rho}}_{i}+\upsilon_{i}{\bm{d}}_{i}$.
3:  Convergence condition: If $|L({\bm{\rho}}_{i+1})-L({\bm{\rho}}_{i})|/|L({\bm{\rho}}_{i+1})|\leq\epsilon_{\rm th}$, ${\bm{y}}_{j}={\bm{\rho}}_{i+1}$, and go to step 5; Otherwise, go to next step.
4:  Direction update: According to (45), if $\varsigma_{i}=\varsigma_{i}^{\rm HS}$, $k=k+1$, or $k=0$ otherwise; when $k=N$, let $\varsigma_{i}=0$ and $k=0$; update ${\bm{d}}_{i+1}=-{\bm{\Theta}}\nabla L({\bm{\rho}}_{i+1})+\varsigma_{i}{\bm{d}}_{i}$; let $i=i+1$, and go back to step 1.
5:  Penalty update: If $\mu_{j}\alpha({\bm{y}}_{j})<\epsilon_{\mu}$, stop; Otherwise, let $\mu_{j+1}=\varphi\mu_{j}$, and ${\bm{d}}_{i+1}=-{\bm{\Theta}}\nabla L({\bm{\rho}}_{i+1})$; let $j=j+1$ and $i=i+1$, and go to step 1.
V-B Modified steepest descent algorithm
In fact, alternative unconstrained optimization methods can also be employed to solve the proposed projection-based penalty problem in (31). For comparison purposes, we also develop a Modified Steepest Descent (MSD) algorithm, adapted from the proposed MCG framework. The primary distinction between the MSD and MCG algorithms lies in the choice of the descent direction. Specifically, in the MSD method, the descent direction at each iteration is directly set as the negative gradient of the objective function
${\bm{d}}_{i}$ $=-{\bm{\Theta}}\nabla L({\bm{\rho}}_{i}),$ (53)
which can be directly obtained from MCG by setting the deflection factor to zero. We herein omit the detailed implementation steps of PP-MSD-ILS since they can be obtained from the PP-MCG-ILS by simply revising step 4 of Algorithm 1.
V-C Benchmarks
We now introduce two benchmark algorithms, derived from the PP-MCG-ILS and PP-MSD-ILS frameworks, respectively. These benchmark variants adopt fixed step sizes at each iteration instead of employing ILS. To mitigate potential instability arising from overly large or small gradient magnitudes under fixed step updates, we normalize the update directions using their $l_{2}$-norm. For the Penalty-function and Projection driven Normalized Steepest Descent (PP-NSD) algorithm, the descent direction is defined as
${\bm{d}}_{i}$ $=-{\bm{g}}_{i}/(\|{\bm{g}}_{i}\|_{2}+\epsilon_{d}),$ (54)
where ${\bm{g}}_{i}={\bm{\Theta}}\nabla L({\bm{\rho}}_{i})$, and $\epsilon_{d}$ is a small constant to prevent division by zero. Similarly, for the penalty function and projection driven normalized CG (PP-NCG) algorithm, the descent direction is given by,
${\bm{d}}_{i+1}$ $=-({\bm{g}}_{i+1}+\varsigma_{i}{\bm{d}}_{i})/(\|({\bm{g}}_{i+1}+\varsigma_{i}{\bm{d}}_{i})\|_{2}+\epsilon_{d}),$ (55)
where the deflection factor $\varsigma_{i}$ is computed by (45).
The step size selection in both benchmark algorithms is
$\upsilon_{i}=\min\{\epsilon_{\upsilon}\upsilon_{i}^{\rm b},\upsilon_{0}\},$ (56)
with $\upsilon_{i}^{\rm b}$ provided by (51) and $\upsilon_{0}$ being a predetermined step size. Additionally, the descent function (38) may not hold when utilizing the fixed step sizes. Therefore, it is necessary to provide a sufficiently large integer $I$ as the upper bound of the number of iterations. In the same vain, due to the non-monotonic penalty function, the minimum may not occur at the final iteration but within a few steps before it. To address this, we introduce an alternative penalty stopping criterion given by $\mu_{j}\min\limits_{i+1-S\leq q\leq i+1}\{\alpha({\bm{\rho}}_{q})\}<\epsilon_{\mu}$ where $S$ represents the window size for searching the minimum penalty value over recent iterations. The detailed implementation steps of the PP-NSD algorithm are omitted here, as it can be readily constructed by modifying step 4 of the PP-NCG algorithm described in Algorithm 2.
Algorithm 2 The PP-NCG algorithm
Input: Positive scalar termination factor $\epsilon_{\mu}$ and $\epsilon_{\rm th}$, a large integer $I$ and a small integer $S$, small positive constants $\sigma_{\upsilon}$, $\epsilon_{L}$, $\epsilon_{d}$ and $\epsilon_{\upsilon}$, initial step size $\upsilon_{0}$, initial penalty factor $\mu_{1}$ and its scaling factor $\varphi>1$; initial PA parameter ${\bm{\rho}}_{1}$ satisfying ${\bm{1}}^{\rm T}{\bm{\rho}}_{1}=1$, projection operator ${\bm{\Theta}}$, ${\bm{y}}_{1}={\bm{\rho}}_{1}$, and ${\bm{d}}_{1}=-{\bm{\Theta}}\nabla L({\bm{\rho}}_{1})/(\|{\bm{\Theta}}\nabla L({\bm{\rho}}_{1})\|_{2}+\epsilon_{d})$; a counter $k=0$ and $i=j=1$.
Output: The PA vector ${\bm{y}}_{j}$.
1:  Step size: $\upsilon_{i}=\min\{\epsilon_{\upsilon}\upsilon_{i}^{\rm b},\upsilon_{0}\}$.
2:  Parameter update: ${\bm{\rho}}_{i+1}={\bm{\rho}}_{i}+\upsilon_{i}{\bm{d}}_{i}$.
3:  Convergence condition: If $|L({\bm{\rho}}_{i+1})-L({\bm{\rho}}_{i})|/|L({\bm{\rho}}_{i+1})|\leq\epsilon_{\rm th}$ or $i+1=I$, ${\bm{y}}_{j}={\bm{\rho}}_{i+1}$, go to step 5; Otherwise, go to next step.
4:  Direction update: According to (45), if $\varsigma_{i}=\varsigma_{i}^{\rm HS}$, $k=k+1$, or $k=0$ otherwise; when $k=N$, let $\varsigma_{i}=0$ and $k=0$; update ${\bm{d}}_{i+1}$ through (55); let $i=i+1$, and go back to step 1.
5:  Penalty update: If $\mu_{j}\min\limits_{i+1-S\leq q\leq i+1}\{\alpha({\bm{\rho}}_{q})\}<\epsilon_{\mu}$, stop; Otherwise, let $\mu_{j+1}=\varphi\mu_{j}$, and ${\bm{d}}_{i+1}=-{\bm{\Theta}}\nabla L({\bm{\rho}}_{i+1})/(\|{\bm{\Theta}}\nabla L({\bm{\rho}}_{i+1})\|_{2}+\epsilon_{d})$; let $j=j+1$ and $i=i+1$, and go to step 1.
VI Extension to pure sensing
In this section, we propose a Penalty-function based Normalized Conjugate Gradient algorithm with inexact line search (P-NCG-ILS) for the PA in pure sensing scenario. We consider the power minimization under power budget constraints and CRLB constraints, i.e.,
$\arg\min\limits_{{\bm{\rho}}}$ $\ {\bm{\rho}}^{\rm T}{\bm{1}},$ (57a)
${\rm s.t.}$ ${\bm{\rho}}^{\rm T}{\bm{1}}\leq 1,$ (57b)
[…]
We present the penalty function as follows
$\alpha_{\rm s}({\bm{\rho}})=$ ${f}_{\rm s}({\bm{\rho}})+\tilde{f}_{l}({\bm{\rho}})+\tilde{f}_{v}({\bm{\rho}})$
$+\sum_{n=1}^{N}\left(f_{\min,n}({\bm{\rho}})+f_{\max,n}({\bm{\rho}})\right),$ (58)
with the penalty function for the total power budget
${f}_{\rm s}({\bm{\rho}})=(\max\{0,{\bm{\rho}}^{\rm T}{\bm{1}}-1\})^{q}.$ (59)
Then, the objective function can be formulated as
$L_{\rm s}({\bm{\rho}})={\bm{\rho}}^{\rm T}{\bm{1}}+\mu\alpha_{\rm s}({\bm{\rho}}).$ (60)
The complete P-NCG-ILS algorithm is summarized in Algorithm 3.
Algorithm 3 The P-NCG-ILS algorithm
Input: Positive scalar termination factors $\epsilon_{\mu}$ and $\epsilon_{\rm th}$, small positive constants $\sigma_{\upsilon}$, $\epsilon_{L}$ and $\epsilon_{\upsilon}$, initial step size $\upsilon_{0}$, initial penalty factor $\mu_{1}$ and its scaling factor $\varphi>1$; initial PA parameter ${\bm{\rho}}_{1}$ satisfying ${\bm{1}}^{\rm T}{\bm{\rho}}_{1}=1$, ${\bm{y}}_{1}={\bm{\rho}}_{1}$, and ${\bm{d}}_{1}=-\nabla L_{\rm s}({\bm{\rho}}_{1})/\|\nabla L_{\rm s}({\bm{\rho}}_{1})\|_{2}$; a counter $k=0$ and $i=j=1$.
Output: The PA vector ${\bm{y}}_{j}$.
1:  Step size: $\upsilon_{i}=\min\{\epsilon_{\upsilon}\upsilon_{i}^{\rm b},\upsilon_{0}\}$; while $L_{\rm s}({\bm{\rho}}_{i}+\upsilon_{i}{\bm{d}}_{i})-L_{\rm s}({\bm{\rho}}_{i})>\epsilon_{L}\upsilon_{i}{\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i})$, let $\upsilon_{i}=\sigma_{\upsilon}\upsilon_{i}$.
2:  Parameter update: ${\bm{\rho}}_{i+1}={\bm{\rho}}_{i}+\upsilon_{i}{\bm{d}}_{i}$.
3:  Convergence condition: If $|L_{\rm s}({\bm{\rho}}_{i+1})-L_{\rm s}({\bm{\rho}}_{i})|/|L_{\rm s}({\bm{\rho}}_{i+1})|\leq\epsilon_{\rm th}$, ${\bm{y}}_{j}={\bm{\rho}}_{i+1}$, and go to step 5; Otherwise, go to next step.
4:  Direction update: According to (45), if $\varsigma_{i}=\varsigma_{i}^{\rm HS}$, $k=k+1$, or $k=0$ otherwise; when $k=N$, let $\varsigma_{i}=0$ and $k=0$; update ${\bm{d}}_{i+1}=\tilde{\bm{d}}_{i+1}/(\|\tilde{\bm{d}}_{i+1}\|_{2}+\epsilon_{d})$ with $\tilde{\bm{d}}_{i+1}=-\nabla L_{\rm s}({\bm{\rho}}_{i+1})+\varsigma_{i}{\bm{d}}_{i}$; let $i=i+1$, and go back to step 1.
5:  Penalty update: If $\mu_{j}\alpha({\bm{y}}_{j})<\epsilon_{\mu}$, stop; Otherwise, let $\mu_{j+1}=\varphi\mu_{j}$, and ${\bm{d}}_{i+1}=-\nabla L_{\rm s}({\bm{\rho}}_{i+1})$; let $j=j+1$ and $i=i+1$, and go to step 1.
VII Algorithm Analysis
The proposed algorithms (e.g., PP-MCG-ILS, PP-MSD-ILS and P-NCG-ILS) can be extended and applied to a broad class of nonlinear and non-convex optimization problems. These problems are typically characterized by an affine equality constraint, if present, satisfied through a projection operator, along with nonlinear or non-convex constraints incorporated via penalty functions. The ILS mechanism is used to enhance convergence. In the following, we analyze the convergence behavior and computational complexity of the proposed algorithms.
VII-A Convergence analysis
We now provide the convergence condition of the proposed PP-MCG-ILS and PP-MSD-ILS algorithms. Before delving to the details, we define the increment of the parameter update and the objective function respectively as
$\Delta{\bm{\rho}}_{i+1}$ $\triangleq{\bm{\rho}}_{i+1}-{\bm{\rho}}_{i}=v_{i}{\bm{d}}_{i}$ (61)
and
$\Delta L({\bm{\rho}}_{i+1})\triangleq L({\bm{\rho}}_{i+1})-L({\bm{\rho}}_{i}).$ (62)
Lemma 1. The sufficient and necessary condition for $\|\Delta{\bm{\rho}}_{i+1}\|\rightarrow 0$ and $\Delta L({\bm{\rho}}_{i+1})\rightarrow 0$ is that, there exists an integer $I$ such that, ${\bm{d}}_{i}\rightarrow{\bm{0}}$ or $v_{i}\rightarrow 0$ when $i>I$.
Proof. The proof is given in Appendix A. $\blacksquare$
Definition. A function $L({\bm{\rho}})$ is Lipschitz gradient continuous over a set $\Omega$, if there exists a constant $C>0$, $\|\nabla L({\bm{\rho}}_{i})-\nabla L({\bm{\rho}}_{j})\|\leq C\|{\bm{\rho}}_{i}-{\bm{\rho}}_{j}\|$ for any ${\bm{\rho}}_{i},{\bm{\rho}}_{j}\in\Omega$ [58, 59].
Theorem 1. Given a nonempty convex set $\Omega$ where the objective function $L({\bm{\rho}}_{i})$ is bounded over its interior points, if the objective function is Lipschitz gradient continuous over $\Omega$, the ILS-based descent direction iterative algorithm will converge to a stationary point with a zero descent direction ${\bm{d}}_{i}\rightarrow{\bm{0}}$. In contrast, if it is not Lipschitz gradient continuous, the sequence $\{{\bm{\rho}}_{i}\}$ and $\{L({\bm{\rho}}_{i})\}$ may converge to a Lipschitz gradient noncontinuous point where the norm of the Hessian matrix may approach infinity.
Proof. The proof is given in Appendix B. $\blacksquare$
Theorem 1 provides the convergence analysis for the proposed PP-MCG-ILS and PP-MSD-ILS algorithms.
We now prove the Lipschitz gradient continuity of the objectivity functions in (31) and (60).
The feasible domain $\Omega\subset\mathbb{R}^{N}$ is closed, bounded, and convex. All CRLB-involved matrices, e.g., ${\bm{P}}(\rho)$, ${\bm{Y}}(\rho)$, and their inverses are uniformly bounded and positive definite over the set $\Omega$.
Firstly, each term $f_{\min,n}({\bm{\rho}})$ and $f_{\max,n}({\bm{\rho}})$ is piecewise quadratic, and its gradient is Lipschitz continuous due to continuity and boundedness. So does ${f}_{\rm s}({\bm{\rho}})$. Then, since ${\bm{C}}_{L}({\bm{\rho}})$ is uniformly positive definite and smooth in ${\bm{\rho}}$, it follows that: $f_{l}({\bm{\rho}})$ is continuously differentiable, and so is
$\nabla\tilde{f}_{l}({\bm{\rho}})=\begin{cases}2f_{l}({\bm{\rho}})\nabla f_{l}({\bm{\rho}}),&\text{if }f_{l}({\bm{\rho}})>0\\ 0,&\text{otherwise}.\end{cases}$ (63)
Given that $\nabla f_{l}({\bm{\rho}})$ is smooth and bounded (due to the matrix inverse being smooth on positive definite matrices), $\nabla\tilde{f}_{l}({\bm{\rho}})$ is Lipschitz continuous. A similar argument applies to $\tilde{f}_{v}({\bm{\rho}})$. Hence, $\nabla_{\bm{\rho}}\alpha({\bm{\rho}})$ and $\nabla_{\bm{\rho}}\alpha_{\rm s}({\bm{\rho}})$ are Lipschitz continuous over $\Omega$, and so are $\nabla_{\bm{\rho}}L({\bm{\rho}})$ and $\nabla_{\bm{\rho}}L_{\rm s}({\bm{\rho}})$.
Additionally, since ${\bm{\Theta}}$ is a constant matrix with $\|{\bm{\Theta}}\|_{2}=1$, and $\nabla_{\bm{\rho}}f({\bm{\rho}})$ is Lipschitz continuous, the composition $\nabla\mathcal{L}({\bm{\varpi}})={\bm{\Theta}}\nabla L({\bm{\rho}})$ is also Lipschitz continuous. There exists a constant $C>0$ such that $\forall{\bm{\varpi}}_{1},{\bm{\varpi}}_{2}\in\Omega_{\bm{\varpi}}$
$\|\nabla_{{\bm{\varpi}}}\mathcal{L}({\bm{\varpi}}_{1})-\nabla_{{\bm{\varpi}}}\mathcal{L}({\bm{\varpi}}_{2})\|_{2}\leq C\|{\bm{\varpi}}_{1}-{\bm{\varpi}}_{2}\|_{2}.$
Therefore, $\mathcal{L}({\bm{\varpi}})$ is Lipschitz gradient continuous on $\Omega_{\bm{\varpi}}$. $\blacksquare$
VII-B Complexity analysis
We observe that the additional computational complexity of the PP-MCG-ILS algorithm, compared to the PP-MSD-ILS algorithm, primarily stems from the computation of $\varsigma_{i}^{\rm b}$ and $\varsigma_{i}^{\rm HS}$ at each iteration, which requires approximately $4N$ multiplications and 2 divisions. In both algorithms, the dominant computational burden lies in evaluating the objective function during the ILS and computing the gradients for direction updates, which are inherently tied to the CRLB expressions. Therefore, a more meaningful comparison of total computational cost focuses on the number of ILS steps per iteration. This number is influenced by the step size threshold, given by the right-hand side of (67), the initial step size $\upsilon_{0,i}$ in (52), and the scaling factor. At a given iterate ${\bm{\rho}}_{i}$, a larger descent direction norm $\|{\bm{d}}_{i}\|$ generally leads to a smaller $\upsilon_{0,i}$ and a reduced step size threshold. As a result, the number of ILS steps is expected to be comparable between the PP-MCG-ILS and PP-MSD-ILS algorithms. Given this, the primary metrics for evaluating the relative performance of these two algorithms become their convergence rate and the quality of their steady-state solutions.
We now qualitatively compare the computational complexity between algorithms employing ILS and those without ILS. As previously discussed, the number of ILS steps significantly impacts the number of evaluations of the objective function. Thus, the computational complexity of the benchmark algorithms (without ILS) can be considerably lower at each iteration during step 1 compared to the PP-MCG-ILS and PP-MSD-ILS algorithms. However, all four algorithms have similar complexity in step 4 due to the inherently intensive gradient calculations involved. Moreover, the overall computational complexity heavily depends on the convergence rates, or equivalently, the total number of iterations required for convergence. As will be demonstrated in the subsequent simulation results, the proposed PP-MCG-ILS algorithm exhibits a notably higher convergence rate compared to its counterparts, with or without ILS. This superior convergence performance significantly reduces the overall computational effort, particularly benefiting the demanding computations of gradients.
VIII Simulation Results
In this section, we conduct some simulations to evaluate the proposed PP-MCG-ILS and PP-MSD-ILS algorithms for PA optimization. All APs, including $N$ single antenna transmitters and $K$ single antenna receivers, are assumed to be stationary. We employ the orthogonal chirp division multiplexing (OCDM) [60] waveform with a single Gaussian pulse shaping. The carrier frequency is $3$GHz. The lowpass equivalent is formulated by,
$s_{n}(t)=(2/T^{2})^{1/4}e^{-\pi t^{2}/T^{2}}e^{j\pi M/T^{2}(t-(n-1)T/M)^{2}},$ (64)
where $M$ is the number of OCDM chirps, and $T$ is proportional to the effective pulse time width.
VIII-A CRLBs
We first verify the CRLB in comparison with the MLE results. To maintain consistency with existing works [38, 40, 41, 39, 36, 34, 35], we use the signal energy-to-noise ratio (SENR), defined as $\tilde{\delta}_{\rm cn}\triangleq P/\sigma_{\rm cn}^{2}$, instead of the SCNR $\delta_{\rm cn}$ in evaluating the CRLBs. Due to the high complexity of four-dimensional MLE, we consider a simple $4\times 3$ nonsymmetric multi-static radar configuration given in Fig. 2. Consider the OCDM waveform with $M=16$, $T=10^{-2}$ s and a sampling rate of $f_{s}=1$ kHz. The velocity of the target is $(20,30)$ m/s.
Fig. 3 compares the mean squared error (MSE) of the MLE with the derived CRLB. It is observed that the derived CRLBs closely match the MSEs for SENR values above $-20$ dB. The small discrepancies in the low-SENR regime are attributed to the suboptimal behavior of the likelihood function under noise-dominated conditions. Overall, these results validate the effectiveness and accuracy of the proposed CRLBs.
VIII-B Power Allocation for ISAC
The performance of the proposed PA algorithms is evaluated by comparison with the PP-NCG and PP-NSD benchmark algorithms. To ensure communication quality, a larger number of APs are designated for transmitting ISAC signals, while a smaller subset is used for receiving target-reflected signals. Without loss of generality, we consider a CF MIMO-ISAC system comprising 10 transmitters and 2 receivers, as illustrated in Fig. 2. To guarantee the CRLB effectiveness, we assume the minimum PA coefficient vector is ${\bm{\rho}}_{\min}=0.01\times{\bm{1}}_{N}$. The PA budget is ${\bm{\rho}}_{\max}=0.3\times{\bm{1}}_{N}$. Consider a target moving slowly with velocity $l_{v}=[4,5]m/s$. A typical modulus squared RCS matrix is given by $\begin{bmatrix}0.37&0.70&1.38&0.65&2.40&0.18&0.25&0.82&0.14&0.35\\ 0.40&2.05&1.14&0.42&0.03&0.11&2.02&1.65&1.24&3.65\end{bmatrix}^{\rm T}$. The channel gains between the user and the transmitters are assumed to be $[2.11,12.57,5.63,0.75,0.61,1.75,0.20,2.34,14.79,9.68]^{\rm T}$. In this ISAC setup, the CRLB thresholds for the position and velocity are given by $[250,0.13]$.
Figure 2: The AP location distribution of the ISAC system: (a) $4\times 3$ radar setting; (b) CF MIMO-ISAC setting.
Figure 3: The MSE versus the CRLB under different SENRs.
The PA vector is initialized as the normalized PA vector ${\bm{\rho}_{1}}={\bm{1}}_{N}/N$, unless noted otherwise. The initial penalty factor is set to be $\mu_{1}=10^{4}$. We hereafter consider the ILS controlling factor $\epsilon_{L}=1e-3$, the initial step size controlling factor $\epsilon_{\upsilon}=0.9$, the penalty factor scaling coefficient $\phi=10$, and the penalty threshold $\epsilon_{\mu}=10^{-3}$, unless noted otherwise. The other simulation conditions are given in Table I. Note that the presented SINRs, PA vectors, and CRLBs, are obtained by averaging 100 points after convergence.
Firstly, we evaluate the convergence and steady-state performance of the proposed algorithms. The SINR learning curves of four algorithms are illustrated in Fig. 4. As apparently seen in Fig. 4, the PP-MCG-ILS has the highest convergence rate, followed by the PP-MSD-ILS, while the counterparts without ILS presents comparatively slow convergence. Additionally, the benchmark algorithms without ILS exhibit numerical fluctuations, particularly under large fixed step sizes. This behavior suggests potential instability and divergence risks. Consequently, we refrain from increasing the step sizes further during simulations to avoid compromising convergence reliability.
The evolution of the PA coefficients and corresponding CRLB values with respect to the number of iterations is illustrated in Fig. 5 and Fig. 6, respectively. The results show that the PP-MCG-ILS algorithm achieves significantly faster convergence compared to the PP-NCG algorithm. Ultimately, all four algorithms converge to a common PA vector, given by ${\bm{\rho}}=[0.01,0.30,0.0638,0.01,\dots,0.01,0.30,0.2762]^{\rm T}$, which satisfies the sensing QoS constraints. Notably, more power is allocated to transmitters 2, 9, and 11 due to their higher communication channel gains. Furthermore, the steady-state CRLB values remain within the predefined threshold limits, confirming the validity and effectiveness of the proposed optimization framework.
TABLE I: Simulation parameters
  Algorithms PP-MCG-ILS PP-MSD-ILS PP-NCG PP-NSD
  Fig. 4 $\{\epsilon_{\rm th},\sigma_{\upsilon}\}=\{1e-11,0.5\}$ same as PP-MCG-ILS $\upsilon\in\{4e-5,2e-5\}$ same as PP-NCG
Figs. 5 and 6 same as Fig. 4 $\upsilon=2e-5$
Figs. 7 and 8 $\epsilon_{\rm th}=1e-12$, $\sigma_{\upsilon}=0.7$ same as PP-MCG-ILS $\upsilon=2e-5$ same as PP-NCG
Figs. 9 and 10 $\{\epsilon_{\rm th},\sigma_{\upsilon}\}=\{1e-11,0.7\}$
Figure 4: SINR with sequential penalty factor
Figure 5: PA coefficient: (a) PP-MCG-ILS; (b) PP-NCG with $\upsilon=2\times 10^{-5}$.
Figure 6: CRLB values: (a) PP-MCG-ILS; (b) PP-NCG with $\upsilon=2\times 10^{-5}$.
The proposed sequential penalty-based algorithms leverage the optimal solution from the previous iteration loop as the initial point for the current loop under an updated penalty factor. To further explore the impact of initialization, we compare the performance of different initial PA strategies under a fixed penalty factor. In addition to the commonly used uniform PA initialization, we introduce a heuristic PA scheme, where the initial allocation is proportional to the communication channel gain, given by ${\bm{\rho}}_{1}={\bm{g}}/\sum_{n}|g_{n}|^{2}$. The results are illustrated in Figs. 7 and 8. When comparing to sequential penalty-based counterparts in Fig. 4, the PP-MCG-ILS and PP-NCG algorithms initialized with uniform PA exhibit similar convergence behavior. However, the PP-MSD-ILS and PP-NSD algorithms show degraded convergence, highlighting the superior convergence properties of CG-based algorithms over SD-based methods under identical conditions. When employing the heuristic PA initialization, all algorithms demonstrate significantly improved convergence rates compared to the uniform PA case. Moreover, both PP-NCG and PP-NSD achieve noticeable performance gains relative to their sequential penalty-based counterparts, particularly PP-NCG. In contrast, the PP-MCG-ILS algorithm shows robust and superior convergence behavior, and PP-MSD-ILS exhibits a slight degradation.
Figure 7: SINR with initial uniform PA
Figure 8: SINR with initial heuristic PA
Figure 9: Steady-state SINRs and CRLBs versus the location CRLB thresholds: (a) SINR; (b) CRLB.
Figure 10: Steady-state SINRs and CRLBs versus the velocity CRLB thresholds: (a) SINR; (b) CRLB.
To summarize, the proposed algorithms exhibit superior convergence behavior under the sequential penalty framework, consistently outperforming the benchmark methods. While the benchmark algorithms can benefit from specific conditions, such as an empirically tuned penalty factor and a well-initialized starting point, their performance is less robust and generalizable. In contrast, the PP-MCG-ILS algorithm demonstrates consistently strong convergence and steady-state accuracy across a wide range of settings, irrespective of initialization or penalty factor selection.
We now investigate the impact of the CRLB thresholds on the performance of the proposed PP-MCG-ILS algorithm. As shown in Fig. 9, relaxing the location CRLB thresholds leads to a gradual increase in the steady-state SINR, which eventually stabilizes. This trend aligns with the behavior observed in Fig. 9, where the achieved location CRLB value closely approaches its threshold and then converge. Furthermore, it is noteworthy that the velocity CRLB remains largely unaffected by variations in the location CRLB thresholds. In contrast, Fig. 10 shows that the performance of the PP-MCG-ILS algorithm exhibits negligible variation with the increased velocity CRLB threshold. The observed CRLB saturation with respect to the corresponding CRLB threshold may be an inherent outcome of the introduced power budget constraints.
Figure 11: Learning results for pure sensing: (a) total power; (b) CRLBs.
VIII-C Power Allocation for pure sensing
We now simply verify the effectiveness of the proposed P-MCG-ILS algorithm for total power minimization in joint location and velocity estimation. The initial penalty factor is set to be $\mu_{1}=1$, $\sigma_{\upsilon}=0.2$ and the penalty threshold $\epsilon_{\mu}=10^{-6}$. The other simulation conditions are the same as the PP-MCG-ILS algorithm in Fig. 4.
The learning results for total power optimization and the corresponding CRLB performance are illustrated in Fig. 11. As seen in Fig. 11, the algorithm demonstrates rapid convergence. Under the same CRLB threshold settings, the total power consumption in the pure sensing scenario is significantly lower than that in the ISAC case. This is expected, as ISAC systems must allocate sufficient power to meet both sensing and communication requirements, particularly for maximizing communication performance. Moreover, both the location and velocity CRLBs in the pure sensing case closely align with their respective thresholds. In contrast, in the ISAC scenario (refer to Fig. 6), the velocity CRLB matches its threshold loosely. This difference arises because, in pure sensing, any enhancement in CRLB performance directly benefits from increased sensing power. However, in ISAC, power allocation is optimized primarily for communication performance, with CRLB constraints being satisfied, potentially loosely, within acceptable margins.
IX Conclusions
This paper presents, for the first time, a comprehensive investigation of the power allocation problem in a cell-free (CF) MIMO-ISAC framework, where joint location and velocity estimation is considered as the sensing task. We employ newly derived approximate Cramér-Rao lower bounds (CRLBs) as sensing performance metrics, alongside user receive SINR as the communication metric. To address the inherent nonlinearity and non-convexity of the resulting optimization problem, we propose a penalty function and projection-based modified conjugate gradient algorithm with inexact line search (PP-MCG-ILS). Additionally, we introduce a penalty function-based normalized conjugate gradient algorithm with inexact line search (P-NCG-ILS) to handle the power minimization problem in pure sensing scenarios. We analyze the convergence behavior and qualitatively compare the computational complexity of all proposed methods. Simulation results confirm the accuracy of the derived CRLBs and demonstrate the efficacy of the proposed power allocation strategies in enhancing pure sensing and overall ISAC performance. The findings offer valuable insights into resource-efficient design for next-generation CF MIMO-ISAC networks and lay the groundwork for future extensions involving more complex sensing tasks and dynamic resource management.
Appendix A Proof of Lemma 1
It is obvious that $\|\Delta{\bm{\rho}}_{i+1}\|\rightarrow 0$ when $i>I$ if and only if either ${\bm{d}}_{i}\rightarrow{\bm{0}}$ or $v_{i}\rightarrow 0$ when $i>I$. On the other hand, in light of (46), we have
$\Delta L({\bm{\rho}}_{i+1})\approx\upsilon_{i}{\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i})+\upsilon_{i}^{2}/2{\bm{d}}_{i}^{\rm T}\nabla^{2}L({\bm{\rho}}_{i}){\bm{d}}_{i}.$ (65)
Firstly, if ${\bm{d}}_{i}\rightarrow{\bm{0}}$ or $v_{i}\rightarrow 0$ for $i>I$, we readily have $\Delta L({\bm{\rho}}_{i+1})\rightarrow 0$. Assume there exists $i>I$, where $\Delta L({\bm{\rho}}_{i+1})\rightarrow 0$ with neither ${\bm{d}}_{i}\rightarrow{\bm{0}}$ nor $v_{i}\rightarrow 0$. We can solve the step size for $\Delta L({\bm{\rho}}_{i+1})=0$ as
$\upsilon_{i}=\dfrac{-2{\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i})}{{\bm{d}}_{i}^{\rm T}\nabla^{2}L({\bm{\rho}}_{i}){\bm{d}}_{i}},$ (66)
which corresponds to the step size bound in the ILS with $\epsilon_{L}=0$. However, we require $\epsilon_{L}>0$ to perform ILS effectively. Thus, ${\bm{d}}_{i}\rightarrow{\bm{0}}$ or $v_{i}\rightarrow 0$ are necessary for $\Delta L({\bm{\rho}}_{i+1})\rightarrow 0$. $\blacksquare$
Appendix B Proof of Theorem 1
In light of (46), (48) and (65), the ILS condition requires
$\upsilon_{i}{\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i})+\upsilon_{i}^{2}/2{\bm{d}}_{i}^{\rm T}\nabla^{2}L({\bm{\rho}}_{i}){\bm{d}}_{i}\leq\epsilon_{L}\upsilon_{i}{\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i})$
$\Longrightarrow\ \upsilon_{i}\leq\dfrac{2(\epsilon_{L}-1){\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i})}{{\bm{d}}_{i}^{\rm T}\nabla^{2}L({\bm{\rho}}_{i}){\bm{d}}_{i}}.$ (67)
Where positive ${\bm{d}}_{i}^{\rm T}\nabla^{2}L({\bm{\rho}}_{i}){\bm{d}}_{i}$ is considered. Assuming the Lipschitz gradient continuity of $L({\bm{\rho}})$, $\|\nabla^{2}L({\bm{\rho}}_{i})\|$ is upper bounded, indicating a positive and finite ${\bm{d}}_{i}^{\rm T}\nabla^{2}L({\bm{\rho}}_{i}){\bm{d}}_{i}$ and thus non-zero $\upsilon_{i}$. Furthermore, in light of (48), we have,
$\Delta^{2}L({\bm{\rho}}_{i+1})\leq\epsilon_{L}^{2}\upsilon_{i}^{2}({\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i}))^{2}.$ (68)
Since $L({\bm{\rho}})$ is a monotonously decreasing and bounded function regarding ${\bm{\rho}}_{i}$, we have $\Delta^{2}L({\bm{\rho}}_{i+1})\rightarrow 0$ when $i>I$ for a large constant $I$. The deflection factor computed by (45) or the descent direction vector in (53) can guarantee the strict descent ${\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i})<0$ at each iteration $i$ unless ${\bm{d}}_{i}={\bm{0}}$. Consequently, sequence $\{L({\bm{\rho}}_{i})\}$ will converge to a stationary point with a zero descent direction ${\bm{d}}_{i}\rightarrow{\bm{0}}$ when it has Lipschitz gradient continuity. According to Lemma 1, we have $\|\Delta{\bm{\rho}}_{i+1}\|\rightarrow 0$ when ${\bm{d}}_{i}\rightarrow{\bm{0}}$. On the other hand, we assume the objective function is not Lipschitz gradient continuous. In this case, if the sequence $\{{\bm{\rho}}_{i}\}$ pass a Lipschitz noncontinuous point ${\bm{\rho}}_{I}$ where ${\bm{d}}_{I}^{\rm T}\nabla^{2}L({\bm{\rho}}_{I}){\bm{d}}_{I}$ is extremely large, to satisfy (67), the step size should meet $\upsilon_{I}\rightarrow 0$. According to Lemma 1, we have $\|\Delta{\bm{\rho}}_{i+1}\|\rightarrow 0$ and $\Delta L({\bm{\rho}}_{i+1})\rightarrow 0$ when $\upsilon_{i}\rightarrow 0$ for $i\geq I$, i.e., the sequences $\{{\bm{\rho}}_{i}\}$ and $\{L({\bm{\rho}}_{i})\}$ converge to a Lipschitz gradient noncontinuous point. $\blacksquare$

### v2

@@
-Power allocation for cell-free MIMO integrated sensing and communication
+Discrete-Time CRLB-based Power Allocation for CF MIMO-ISAC with Joint Localization and Velocity Sensing
 Abstract
-In this paper, we investigate integrated sensing and communication (ISAC) in a cell-free (CF) multiple-input multiple-output (MIMO) network with single-antenna access points (APs), where each AP functions either as a transmitter for both sensing and communication or as a receiver for target-reflected signals.
-We derive closed-form Cramér–Rao lower bounds (CRLBs) for location and velocity estimation under arbitrary power allocation ratios, assuming the radar cross-section (RCS) is deterministic and unknown over the observation interval.
-A power allocation optimization problem is formulated to maximize the communication signal-to-interference-plus-noise ratio (SINR), subject to CRLB-based sensing constraints and per-transmitter power limits.
+In this paper, we investigate integrated sensing and communication (ISAC) in a cell-free (CF) multiple-input multiple-output (MIMO) network, where each access point functions either as an ISAC transmitter or as a sensing receiver.
+We devote into the ISAC sensing metric using the discrete-time signal-based Cramér–Rao lower bounds (CRLBs) for joint location and velocity estimation under arbitrary power allocation ratios under the deterministic radar cross section assumption (RCS).
+Then, we consider the power allocation optimization problem for the CF MIMO-ISAC as the maximization of the communication signal-to-interference-plus-noise ratio (SINR), subject to CRLB-based sensing constraints and per-transmitter power limits.
 To solve the resulting nonlinear and non-convex problem, we propose a penalty function and projection-based modified conjugate gradient algorithm with inexact line search (PP-MCG-ILS), and an alternative method based on a modified steepest descent approach (PP-MSD-ILS).
-Additionally, for power minimization in pure sensing scenarios, we introduce a penalty function-based normalized conjugate gradient algorithm (P-NCG-ILS).
-We analyze the convergence behavior and qualitatively compare the computational complexity of the proposed algorithms.
+We show that the proposed algorithms are scalable and can be extended to a broad class of optimization problems involving nonlinear inequality constraints and affine equality constraints.
+In addition, we extend the PP-MCG-ILS algorithm to the pure sensing scenario, where a penalty function-based normalized conjugate gradient algorithm (P-NCG-ILS) is developed for sensing power minimization.
+Finally, we analyze the convergence behavior and qualitatively compare the computational complexity of the proposed algorithms.
 Simulation results confirm the accuracy of the derived CRLBs and demonstrate the effectiveness of the proposed power allocation strategies in enhancing both sensing and overall ISAC performance.
@@
 I Introduction
-Integrated sensing and communication (ISAC) has emerged as a key enabling technology for sixth-generation (6G) applications, such as smart cities, autonomous driving, and remote healthcare [1, 2].
-Unlike traditional systems with separate designs for communication and sensing, ISAC unifies these functionalities within shared hardware platforms and spectral resources, thereby significantly enhancing spectral, energy, and hardware efficiency [3, 4].
-Depending on the geometric deployment of transmitters and receivers, various ISAC system architectures have been proposed.
-In mono-static ISAC systems—where the transmitter and receiver are co-located—advanced beamforming techniques are employed to mitigate self-interference in full-duplex operations [5, 6].
-In contrast, cell-free (CF) or multi-static ISAC systems involve widely distributed transmitters and receivers.
-These systems inherit advantages from both CF communication and distributed multi-static radar due to their architectural similarities [7].
-MIMO (multiple-input, multiple-output) technology plays a central role by employing orthogonal waveforms across transmitters to achieve multiplexing and diversity gains for both communication and sensing [8].
-Owing to these benefits, CF MIMO-empowered ISAC has been actively explored for unified communication-sensing operations [9, 10, 11, 12, 13, 14].
-ISAC systems address a range of sensing tasks, including target detection [15, 16, 17], and parameter estimation for radar cross-section (RCS) [18, 19], location [20, 21, 22], velocity [23, 24], and angle [25, 26].
-These capabilities support diverse applications from navigation to environmental monitoring.
-A fundamental challenge in ISAC system design is the formulation of robust performance metrics, which are critical for resource optimization and balancing communication-sensing trade-offs [3, 4].
-While unified metrics based on mutual information [27] and Kullback–Leibler divergence [28] have been proposed, tailoring such metrics to specific sensing objectives remains challenging.
-Consequently, task-specific metrics for sensing and communication continue to be widely adopted.
+Integrated sensing and communication (ISAC) has emerged as a key enabling technology for sixth-generation (6G) applications, including smart cities, autonomous driving, and remote healthcare [1].
+Unlike traditional systems that separate the design of communication and sensing functionalities, ISAC integrates both within a unified framework using shared hardware platforms and spectral resources.
+This integration offers substantial improvements in spectral efficiency, energy savings, and hardware utilization [2, 3].
+ISAC systems can be broadly categorized based on the spatial configuration of transmitters and receivers.
+Monostatic ISAC systems, where the transmitter and receiver are co-located, typically rely on advanced beamforming techniques to mitigate self-interference in full-duplex operations [4, 5].
+In contrast, multistatic ISAC systems deploy spatially separated transmitters and receivers, facilitating spatial diversity and sensing coverage [6, 7].
+Building upon the multistatic architecture, the concept of cell-free (CF) multiple-input and multiple-output (MIMO) ISAC has recently emerged [8, 9].
+In CF MIMO-ISAC systems, a large number of distributed access points (APs) jointly serve communication users and provide sensing services in a coordinated manner without relying on conventional cell boundaries.
+This architecture inherits the advantages of both distributed multistatic radar and CF MIMO communication, offering improved spectral efficiency, spatial resolution, and system scalability [10].
+Recent advancements in CF MIMO-ISAC have extended from basic trade-off analysis [11] to complex sensing tasks, including target detection [12], and parameter estimation for radar cross-section (RCS) [13], location [14], velocity [15], and angle [16].
+For example, [12] studies the fundamental performance tradeoff between the target detection probability and the user’s achievable rate in ISAC systems by deriving the probability of false alarm and the successful probability of detection for the target of interest.
+[13] provides a unified framework integrating sensing (RCS estimation), computing, and communication to optimize limited system resource.
+[14] proposes a two-stage scheme for target localization, with communication signals reused as sensing reference signals.
+Additionally, joint sensing and communication beamforming has been explored in [17].
+Despite these advances, a fundamental challenge in ISAC system design, including CF MIMO-ISAC, is the formulation of robust performance metrics, which are critical for guiding system design and balancing communication-sensing trade-offs [2, 3].
+While unified metrics based on mutual information [18] and Kullback–Leibler divergence [19] have been proposed for general analysis, adapting these metrics to task-specific sensing objectives remains a significant challenge.
+Consequently, task-specific metrics for sensing and communication are still widely adopted.
 For communication, Shannon capacity underpins metrics such as spectral efficiency [20], achievable rate [21], and signal-to-interference-plus-noise ratio (SINR).
-Other metrics like coverage probability and energy efficiency have also been addressed in scenarios with stringent requirements [31].
-On the sensing side, metrics such as sensing mutual information [32, 33], sensing SINR [9, 10], and the Cramér–Rao lower bound (CRLB) [34, 35, 36, 37, 38, 39, 40, 41] have been explored extensively, with CRLB-based designs proving particularly effective.
-CRLB-based approaches fall into two categories depending on the RCS model:
-Random RCS assumption: RCS values are treated as random variables, and likelihood functions are derived by averaging over their distributions [34, 35, 36, 37].
-These studies established CRLBs for joint location and velocity estimation, and later extended them to multi-target scenarios.
-Deterministic RCS assumption: RCS values are assumed to be deterministic but unknown and spatially diverse [38, 40, 41, 39].
-This model is particularly relevant in slowly varying environments or when prior RCS knowledge is unavailable.
-Under this setting, CRLBs for both single-target and multiple-target location estimation have been analyzed, with [39] showing convergence to single-target performance as target separations increase.
-However, most of the above CRLB formulations rely on continuous-time signal models, while practical radar systems process discrete-time signals, where the sampling rate significantly impacts sensing accuracy.
-Our prior work [42] introduces a discrete-time likelihood function over Gaussian channels as a more realistic model.
-Building on this, we derive CRLBs for joint location and velocity estimation under the deterministic RCS assumption in both single- and multiple-target scenarios.
-Another critical challenge lies in resource allocation, involving time [43], frequency [44, 45], space [46, 47], and power [9, 10, 12, 11, 13, 14].
-While recent works have addressed power allocation for maximizing sensing SINR or minimizing power under CRLB and communication constraints [9, 10, 12, 11, 13], they typically consider localization only, neglecting velocity or angular estimation, and are based on continuous-time models that overlook the impact of sampling rate.
-In contrast, this paper focuses on joint location and velocity estimation using discrete-time signal models.
-Even with simplified CRLBs, the resulting CRLB-constrained power allocation problems are highly nonlinear and non-convex, posing significant challenges to analytical optimization.
-To address this, we leverage advanced optimization techniques including penalty methods, projection operations, and gradient-based search, building algorithms suitable for such complex constrained problems.
-I-A Contribution
-This paper addresses the power allocation problem in a CF MIMO-ISAC system with widely distributed single-antenna access points (APs), where each AP operates as either an ISAC transmitter or sensing receiver.
-We derive CRLBs for joint location and velocity estimation using discrete received signals under the deterministic RCS assumption.
-These CRLBs are adopted as sensing metrics to guide power allocation under communication SINR maximization objectives.
-The main contributions are as follows.
-1) We derive the approximate CRLBs for joint location and velocity estimation of a single target under arbitrary power allocations.
-The CRLBs are validated against MLE results, confirming their effectiveness as performance metrics.
-With the sensing requirements constrained by the CRLBs of both location and velocity estimation, we propose a Penalty-function based Normalized Modified Conjugate Gradient algorithm with Inexact Line Search (P-NCG-ILS) to generate a power allocation scheme for the total power minimization.
-Simulation results verify its effectiveness.
-2) We formulate the CF MIMO-ISAC power allocation problem as a communication SINR maximization task, subject to constraints on both location and velocity CRLBs and total/individual power budgets.
-To tackle the non-convexity and nonlinearity, we propose a Penalty-function and Projection-based Modified Conjugate Gradient algorithm with Inexact Line Search (PP-MCG-ILS) together with PP-MSD-ILS algorithm based on the modified steepest descent algorithm (MSD).
-3) The proposed PA algorithms for sensing and CF MIMO-ISAC can be broadly applicable to constrained non-convex optimization problems involving affine equality constraints (handled via projection) and nonlinear inequality constraints (handled via penalties).
-We analyze their convergence and computational complexity.
-Simulation results demonstrate the robust performance, fast convergence, and effectiveness of the proposed algorithms (e.g., PP-MCG-ILS) in optimizing joint sensing and communication performance.
-The remainder of the following parts of this paper is organized as follows.
-Section II describes the received signal model for ISAC.
-Section IV formulates the power allocation problem for ISAC.
-Section V details the proposed ISAC power allocation algorithms.
-Section VII gives the convergence and computational complexity analysis.
-Section VIII illustrates the simulation results.
-Section IX concludes this paper.
+Additional metrics, including coverage probability and energy efficiency, have also been considered, particularly in scenarios with stringent quality-of-service (QoS) or energy constraints [22].
+On the sensing side, a variety of performance metrics have been investigated, such as sensing mutual information [23], sensing SINR [24, 25], and the Cramér–Rao lower bound (CRLB) [26, 27, 28, 29].
+Among these, CRLB-based formulations have proven particularly effective for parameter estimation and algorithm benchmarking.
+CRLB-based approaches generally fall into two categories depending on the RCS model: deterministic and stochastic.
+Under the stochastic RCS assumption, RCS values are modeled as random variables, and the corresponding likelihood functions are derived by averaging over their statistical distributions.
+In contrast, the deterministic RCS assumption treats RCS values as deterministic but unknown quantities that may vary spatially.
+This model is particularly applicable in slowly varying environments or when prior RCS knowledge is unavailable.
+CRLBs for both single-target and multiple-target location estimation have been analyzed under the stochastic RCS assumption [26, 27] and the deterministic RCS assumption [28, 29].
+However, most existing CRLB formulations are built upon continuous-time signal models, whereas practical radar systems operate in the discrete-time domain, where the sampling rate critically affects sensing accuracy.
+Our prior work [30] addresses this gap by introducing a discrete-time likelihood function over Gaussian channels, offering a more realistic modeling framework.
+Building on this foundation, this paper devotes into the CRLBs derivation for joint location and velocity estimation under the deterministic RCS assumption, covering both single- and multi-target cases in CF MIMO-ISAC systems.
+Beyond performance characterization, another critical challenge lies in resource allocation, involving time [31], frequency [32], space [33], and power [24, 25, 34, 35, 36].
+Recent works in CF MIMO-ISAC systems have addressed power allocation (PA) under sensing SINR constraints [24, 25] or sensing CRLB constraints [34, 35, 36].
+These involved sensing CRLBs typically consider localization only, neglecting velocity or angular estimation, and depend on continuous-time models that overlook the impact of sampling rate.
+These considerations motivate us to consider a more comprehensive and realistic setting: joint location and velocity estimation in CF MIMO-ISAC systems under a discrete-time signal model with a deterministic RCS assumption.
+The CRLBs are serve as effective sensing metrics to guide PA across distributed APs, with the goal of maximizing communication SINR.
+This leads to a class of non-convex and highly nonlinear optimization problems.
+To address this issue, we develop a tailored optimization framework that integrates penalty methods, projection operations, and gradient-based search, enabling efficient and reliable algorithmic solutions under complex practical constraints.
+The main contributions are as follows:
+  • We derive novel closed-form approximate CRLBs for joint location and velocity estimation under arbitrary PA strategies.
+These CRLBs are rigorously validated against maximum likelihood estimation (MLE) results, confirming their effectiveness as practical sensing performance metrics.
+Based on these metrics, we formulate the PA optimization problems for ISAC, aiming at communication SINR maximization while ensuring joint CRLB and power budget constraints.
+  • To tackle the non-convexity and nonlinearity in PA optimization for ISAC, we propose a penalty-function and projection-based modified conjugate gradient algorithm with inexact line search (PP-MCG-ILS) together with PP-MSD-ILS algorithm based on the modified steepest descent algorithm (MSD).
+We further propose a penalty-function-based normalized modified conjugate gradient algorithm with inexact line search (P-NCG-ILS), which can effectively solve the power minimization problem subject to sensing accuracy constraints.
+  • We provide a comprehensive analysis of their convergence behavior and computational complexity.
+The design ensures stable updates, rapid convergence and low computational overhead.
+Additionally, we show that the proposed PP-MCG-ILS and P-NCG-ILS algorithms can be broadly scalable to a class of constrained non-convex optimization problems involving affine equality constraints (handled via projection) and nonlinear inequality constraints (handled via penalties).
+  • Extensive simulations validate the performance of the developed algorithms in both ISAC and sensing-only scenarios.
+The PP-MCG-ILS algorithm consistently exhibits fast convergence, robust numerical stability, and optimal PA efficiency across various conditions, as compared to benchmark methods.
+Results also highlight important trade-offs between sensing and communication performance and confirm that the proposed methods can meet stringent sensing QoS while maximizing communication throughput or minimizing power usage.
+Section II introduces the received signal model for the CF MIMO-ISAC system.
+Section III-B formulates the power allocation problem for CF MIMO-ISAC.
+Section IV presents the proposed power allocation algorithms for both CF MIMO-ISAC and the sensing-only scenario.
+Section VI analyzes the convergence properties and computational complexity of the proposed algorithms.
+Section VII provides simulation results to validate the theoretical analysis and algorithmic performance.
+Finally, Section VIII concludes the paper.
 […]
@@
 1, we consider a CF MIMO-ISAC system with widely separated APs.
-Each AP either serves as an ISAC transmitter or a sensing receiver, similar to the setup in [9, 10].
-Assume there are $N$ single-antenna transmitters and $K$ single-antenna receivers, with the positions of the transmitter $n$ and receiver $k$ denoted as ${\bm{l}}_{n}=[x_{n},y_{n}]^{\rm T}$ and ${\bm{l}}_{k}=[x_{k},y_{k}]^{\rm T}$, respectively.
+Each AP either serves as an ISAC transmitter or a sensing receiver, similar to the setup in [24].
+Assume that there are $N$ single-antenna transmitters and $K$ single-antenna receivers, with the positions of the transmitter $n$ and receiver $k$ denoted as ${\bm{l}}_{n}=[x_{n},y_{n}]^{\rm T}$ and ${\bm{l}}_{k}=[x_{k},y_{k}]^{\rm T}$, respectively.
 All transmitters and receivers are scheduled and managed by the edge cloud and the core network, e.g., the fifth generation (5G) communication networks.
-Denote the location and velocity vectors of the $q$th sensing target as ${\bm{l}}_{q}=[x_{q},y_{q}]^{\rm T}$ and ${\bm{v}}_{q}=[v_{x,q},v_{y,q}]^{\rm T}$, respectively.
+Denote the location and velocity vectors of the sensing target as ${\bm{l}}=[x,y]^{\rm T}$ and ${\bm{v}}=[v_{x},v_{y}]^{\rm T}$, respectively.
 Figure 1: CF MIMO system for ISAC.
-The Doppler frequency intorduced by the relative radial velocity of target $q$ versus any transmitter $n$ and receiver $k$, is computed as
-$\small f_{n,k,q}=\dfrac{1}{\lambda}{\bm{v}}_{q}^{\rm T}\left(\frac{{\bm{l}}_{n}-{\bm{l}}_{q}}{\|{\bm{l}}_{n}-{\bm{l}}_{q}\|_{2}}+\frac{{\bm{l}}_{k}-{\bm{l}}_{q}}{\|{\bm{l}}_{k}-{\bm{l}}_{q}\|_{2}}\right),$ (1)
-where $\lambda$ denotes the wavelength. Denote $\tau_{n,k,q}$ as the propagation delay between the $n$th transmitter and the $k$th receiver, as reflected by the $q$th target. Then, we have
-$\small\tau_{n,k,q}$ $=\tau_{n,q}+\tau_{k,q},$ (2a)
-$\tau_{n,q}$ $={\|{\bm{l}}_{n}-{\bm{l}}_{q}\|_{2}}/{c},$ (2b)
-$\tau_{k,q}$ $={\|{\bm{l}}_{k}-{\bm{l}}_{q}\|_{2}}/{c},$ (2c)
-where $\tau_{n,q}$ and $\tau_{q,k}$ represent the propagation delays from the $n$th transmitter to the $q$th target, and from the target to the $k$th receiver, respectively, and $c$ denotes the speed of light.
-In this paper, we consider the single-target scenario, and thus the subscript $q$ can be omitted.
+The Doppler frequency induced by the relative radial velocity of the target with respect to transmitter $n$ and receiver $k$ is computed as
+$f_{n,k}=\dfrac{1}{\lambda}{\bm{v}}^{\rm T}\left(\frac{{\bm{l}}_{n}-{\bm{l}}}{\|{\bm{l}}_{n}-{\bm{l}}\|_{2}}+\frac{{\bm{l}}_{k}-{\bm{l}}}{\|{\bm{l}}_{k}-{\bm{l}}\|_{2}}\right),$ (1)
+where $\lambda$ denotes the wavelength. Denote $\tau_{n,k}$ as the propagation delay between the $n$th transmitter and the $k$th receiver, as reflected by the target. Then, we have
+$\tau_{n,k}$ $=\tau_{n}+\tau_{k},$ (2a)
+$\tau_{n}={\|{\bm{l}}_{n}-{\bm{l}}\|_{2}}/{c},$ $\quad\tau_{k}={\|{\bm{l}}_{k}-{\bm{l}}\|_{2}}/{c},$ (2b)
+where $\tau_{n}$ and $\tau_{k}$ represent the propagation delays from the $n$th transmitter to the $q$th target, and from the target to the $k$th receiver, respectively, and $c$ denotes the speed of light.
 […]
 ${\bm{x}}_{n}(t)=\sqrt{P\rho_{n}}{b}_{n}\varsigma s_{n}(t),$ (3)
-where $P$ denotes the total available energy, $\rho_{n}$ is the power allocation (PA) factor with $\sum_{n=1}^{N}\rho_{n}=1$, ${b}_{n}$ is the normalized transmit beamforming weight with $|{b}_{n}|^{2}=1$, and $\varsigma$ is the complex data symbol with $\mathbb{E}\{|\varsigma|^{2}\}=1$.
-Denote ${\bm{\rho}}=[\rho_{1},\rho_{2},\cdots,\rho_{N}]^{\rm T}$ by the vector of the PA coefficient of the $N$ transmitters.
-For target-reflected reception, we refer the term “RCS” as encompassing both channel fading and radar RCS.
-Then, we define the RCS as ${\alpha}_{n,k}$ over transmitter $n$ and receiver $k$.
-We consider the Swerling-I sensing model, i.e.
-the deterministic RCS assumption where RCSs are assumed to be deterministic but unknown during the observation time [48, 38, 40, 41, 49].
-The lowpass equivalent of the reflected signal from the $n$th transmitter by a target received at receiver $k$ is given by [^1] The transmitted signals over different transmitters are orthogonal and maintain approximate orthogonality over different delays and Doppler frequencies such that the signals contributed from different transmitters can be separated at each receiver [50, 8, 34, 35, 36].
+where $P$ denotes the total available energy, $\rho_{n}$ is the PA factor with $\sum_{n=1}^{N}\rho_{n}=1$, ${b}_{n}\in\mathbb{C}$ is the normalized transmit beamforming weight with $|{b}_{n}|^{2}=1$, and $\varsigma\in\mathbb{C}$ has unit variance $\mathbb{E}\{|\varsigma|^{2}\}=1$.
+Let ${\bm{\rho}}=[\rho_{1},\rho_{2},\cdots,\rho_{N}]^{\rm T}$ be the vector of the PA coefficient of the $N$ transmitters.
+Sensing model: For target-reflected reception, we refer the term “RCS” as encompassing both channel fading and radar RCS.
+Denote the RCS associated with transmitter $n$ and receiver $k$ as ${\alpha}_{n,k}$.
+We consider the Swerling-I sensing model [37], i.e.
+the deterministic RCS assumption where RCSs are assumed to be deterministic but unknown during the observation time [29, 28].
+Then, the lowpass equivalent of the reflected signal from the $n$th transmitter by a target received at receiver $k$ is given by [^1] The transmitted signals over different transmitters are orthogonal and maintain approximate orthogonality over different delays and Doppler frequencies such that the signals contributed from different transmitters can be separated at each receiver [29, 26].
 $r_{n,k}(t)=\alpha_{n,k}y_{n,k}(t)+z_{n,k}(t),$ (4)
-where $z_{n,k}(t)$ denotes the clutter-plus-noise and $y_{n,k}(t)$ is the loss-free and noise-free received signal,
+where $z_{n,k}(t)$ denotes the clutter-plus-noise and $y_{n,k}(t)$ is the loss-free and noise-free received signal, i.e.,
 ${y}_{n,k}(t)=\sqrt{P\rho_{n}}b_{n}s_{n}(t-\tau_{n,k})e^{j2\pi f_{n,k}t}.$ (5)
-Note that clutters can initially be mitigated by Doppler processing or space-time adaptive processing [51] and whitened through a whitening filter [34].
-It is reasonable to assume the clutter-plus-noise to be additive white Gaussian noise with zero mean and variance $\sigma_{\rm cn}^{2}$ [8, 52, 53, 34, 35, 36].
-Define $\delta_{\rm cn}=\frac{P}{T_{\rm eff}\sigma_{\rm cn}^{2}}$ as the signal-to-clutter-plus-noise ratio (SCNR).
-The received signal of the communication user is given by,
-${r}_{u}(t)$ $=\sqrt{P}\sum_{n=1}^{N}g_{n}{b}_{n}\sqrt{\rho_{n}}s_{n}(t)+{z}(t),$ (6)
-where $g_{n}$ is the channel fading between the transmitter $n$ and the user, and the interference-plus-noise ${z}(t)$ is assumed to be the additive circular Gaussian white with zero mean and variance $\sigma_{z}^{2}$.
+Define the signal-to-clutter-plus-noise ratio (SCNR) as $\delta_{\rm cn}=\frac{P}{T_{\rm eff}\sigma_{\rm cn}^{2}}$ with $T_{\rm eff}$ denoting the effective waveform width.
+Communication model: The received signal of the communication user is given by
+$\small{r}_{u}(t)$ $=\sqrt{P}\sum\nolimits_{n=1}^{N}g_{n}{b}_{n}\sqrt{\rho_{n}}s_{n}(t)+{z}(t),$ (6)
+where $g_{n}$ is the channel fading between the transmitter $n$ and the user, and $z(t)$ denotes the interference-plus-noise with zero mean and variance $\sigma_{z}^{2}$.
 Then, the SINR at the user terminal can be given by
-$\gamma$ $=\sum_{n=1}^{N}|g_{n}{b}_{n}|^{2}\rho_{n}\delta\overset{(\text{i})}{=}\delta{\bm{\rho}}^{\rm T}{\bm{g}}$ (7)
-where $\delta=\frac{P}{T_{\rm eff}\sigma_{z}^{2}}$ is the total SINR and ${\bm{g}}$ denotes the channel gain [^2] The communication channels are assumed to be estimated a priori using pilot signals [9, 10, 11, 12, 13].
+$\small\gamma$ $=\sum\nolimits_{n=1}^{N}|g_{n}{b}_{n}|^{2}\rho_{n}\delta\overset{(\text{i})}{=}\delta{\bm{\rho}}^{\rm T}{\bm{g}}$ (7)
+where $\delta=\frac{P}{T_{\rm eff}\sigma_{z}^{2}}$ is the total SINR and ${\bm{g}}$ is the channel gain [^2] The communication channels are assumed to be estimated a priori using pilot signals [24, 35, 34, 36].
 with its $n$th element being $|g_{n}|^{2}$.
-Note that $(\mathrm{i})$ holds when using a normalized conjugate beamforming weight with ${b}_{n}={g}^{*}_{n}/|g_{n}|$.
-III Sensing Metric
+Note that $(\mathrm{i})$ holds when using a normalized conjugate beamforming with ${b}_{n}={g}^{*}_{n}/|g_{n}|$.
+III The Proposed Sensing Metric and Problem Formulation
 In this section, we present an approximate CRLB for CF MIMO radar under arbitrary power allocations.
-First define the squared effective bandwidth (SEBW) and squared effective pulse time width (SETW) as
+To begin with, we introduce the following definitions. Define the squared effective bandwidth (SEBW) and squared effective pulse time width (SETW) as
 […]
+[…]
+[…]
+[…]
+[…]
+respectively.
+III-A Sensing Metric
+Unlike the sensing performance metrics employed in [24, 25, 34, 35, 36], which only consider target detection or location estimation, we consider the joint location and velocity CRLB as the sensing metric, developed based on the discrete-time reception of the reflected signal under the deterministic RCS assumption.
+In particular, we have derived the approximate CRLBs for both location and velocity in [30] for simplified sensing metrics.
+They are respectively given by
+$\small{\bm{C}}_{\rm L}$ $=({\bm{P}}-{\bm{V}}{\bm{Y}}^{-1}{\bm{V}}^{\rm T})^{-1},$ (11)
 and
-[…]
-respectively. The cross-term between time and frequency is further defined as
-[…]
-We further define the average frequency as
-[…]
-and average time as
-[…]
-With a relatively large bandwidth and pulse time width, such that $\bar{f}_{n}^{2}\ll\bar{f_{n}^{2}}$ and $(\bar{t}_{n}+\tau_{n,k})^{2}\ll\bar{t_{n}^{2}}$, we have derived the approximate CRLBs for both location and velocity [42], respectively given by
-${\bm{C}}_{\rm L}$ $=({\bm{P}}-{\bm{V}}{\bm{Y}}^{-1}{\bm{V}}^{\rm T})^{-1},$ (13)
-and
-${\bm{C}}_{\rm V}$ $=({\bm{Y}}-{\bm{V}}^{\rm T}{\bm{P}}^{-1}{\bm{V}})^{-1}.$ (14)
-The corresponding $2\times 2$ sub-matrix ${\bm{P}}$ is given in (15), ${\bm{V}}$ given in (16),
-${\bm{P}}=\sum_{n=1}^{N}\sum_{k=1}^{K}\begin{bmatrix}\beta_{n,k}^{2}a_{n,k}+2\beta_{n,k}\eta_{n,k}b_{n,k}+\eta_{n,k}^{2}d_{n,k}&\beta_{n,k}\zeta_{n,k}a_{n,k}+(\beta_{n,k}\kappa_{n,k}+\zeta_{n,k}\eta_{n,k})b_{n,k}+\eta_{n,k}\kappa_{n,k}d_{n,k}\\ \zeta_{n,k}\beta_{n,k}a_{n,k}+(\zeta_{n,k}\eta_{n,k}+\beta_{n,k}\kappa_{n,k})b_{n,k}+\kappa_{n,k}\eta_{n,k}d_{n,k}&\zeta_{n,k}^{2}a_{n,k}+2\zeta_{n,k}\kappa_{n,k}b_{n,k}+\kappa_{n,k}^{2}d_{n,k}\end{bmatrix},$ (15)
-${\bm{V}}=\sum_{n=1}^{N}\sum_{k=1}^{K}\begin{bmatrix}\beta_{n,k}\xi_{n,k}b_{n,k}+\eta_{n,k}\xi_{n,k}d_{n,k}&\beta_{n,k}\varrho_{n,k}b_{n,k}+\eta_{n,k}\varrho_{n,k}d_{n,k}\\ \zeta_{n,k}\xi_{n,k}b_{n,k}+\kappa_{n,k}\xi_{n,k}d_{n,k}&\zeta_{n,k}\varrho_{n,k}b_{n,k}+\kappa_{n,k}\varrho_{n,k}d_{n,k}\end{bmatrix},$ (16)
-and
-${\bm{Y}}=\sum_{n=1}^{N}\sum_{k=1}^{K}\begin{bmatrix}\xi_{n,k}^{2}d_{n,k}&\xi_{n,k}\varrho_{n,k}d_{n,k}\\ \varrho_{n,k}\xi_{n,k}d_{n,k}&\varrho_{n,k}^{2}d_{n,k}\end{bmatrix}.$ (17)
-The involved waveform-related parameters, e.g., $a_{n,k}$, $b_{n,k}$ and $d_{n,k}$, are respectively given by,
-$a_{n,k}$ $=\frac{8\pi^{2}|\alpha_{n,k}|^{2}f_{s}P\rho_{n}\bar{f_{n}^{2}}}{\sigma_{z}^{2}},$ (18a)
-$b_{n,k}$ $=\frac{4\pi|\alpha_{n,k}|^{2}f_{s}P\rho_{n}}{\sigma_{z}^{2}}\mathfrak{T}\{\sigma_{tf}\},$ (18b)
-$d_{n,k}$ $\approx\frac{8\pi^{2}|\alpha_{n,k}|^{2}f_{s}P\rho_{n}\bar{t_{n}^{2}}}{\sigma_{z}^{2}},$ (18c)
-where the approximation holds in (18c) when the effective time width meets $\tau_{n,k}\ll T_{\rm eff}$.
-The relevant geometric spread parameters, including $\beta_{n,k}$, $\zeta_{n,k}$, $\xi_{n,k}$, $\varrho_{n,k}$, $\eta_{n,k}$ and $\kappa_{n,k}$, are omitted here and can be referred to [34, 42].
-To visualize the relationship between the CRLBs and the PA coefficients, we now reformulate the CRLB expressions.
-Given the equivalent location FIM Let $p_{ij}$ denote the $\{i,j\}$th element of matrix ${\bm{P}}$, similar for $v_{ij}$ and $y_{ij}$.
-The elements $p_{ij}$, $v_{ij}$ and $y_{ij}$ can be expressed as the weighted sum of $\rho_{n}$, i.e.,
-$\begin{cases}p_{ij}=\sum_{n=1}^{N}w_{n}^{p_{ij}}\rho_{n}={\bm{\rho}}^{\rm T}{\bm{w}}_{p_{ij}},\\ v_{ij}=\sum_{n=1}^{N}w_{n}^{v_{ij}}\rho_{n}={\bm{\rho}}^{\rm T}{\bm{w}}_{v_{ij}},\\ y_{ij}=\sum_{n=1}^{N}w_{n}^{y_{ij}}\rho_{n}={\bm{\rho}}^{\rm T}{\bm{w}}_{y_{ij}},\end{cases}$ (19)
-where $w_{n}^{p_{ij}}$ is the $n$th element of the weight vector ${\bm{w}}_{p_{ij}}$, similar for $w_{n}^{v_{ij}}$ and $w_{n}^{y_{ij}}$.
-These weights can be readily obtained from (15)-(17) by extracting the PA vector ${\bm{\rho}}$.
-By substituting (19) into (13), the CRLB matrix for location estimation can be rewritten as,
-${\bm{C}}_{\rm L}=\frac{1}{|{\bm{U}}|}\begin{bmatrix}u_{22}&-u_{12}\\ -u_{21}&u_{11}\end{bmatrix},$ (20)
-where $|{\bm{U}}|=u_{11}u_{22}-u_{12}u_{21}$ and
-$u_{11}=\frac{1}{|{\bm{Y}}|}\big{[}{\bm{\rho}}^{\rm T}({\bm{w}}_{p_{11}}{\bm{\rho}}^{\rm T}{\bm{W}}_{y}-{\bm{w}}_{v_{11}}{\bm{\rho}}^{\rm T}{\bm{W}}_{12}^{y}-{\bm{w}}_{v_{12}}{\bm{\rho}}^{\rm T}\breve{{\bm{W}}}_{11}^{y}){\bm{\rho}}\big{]},$ (21a)
-$u_{12}=\frac{1}{|{\bm{Y}}|}\big{[}{\bm{\rho}}^{\rm T}({\bm{w}}_{p_{12}}{\bm{\rho}}^{\rm T}{\bm{W}}_{y}-{\bm{w}}_{v_{21}}{\bm{\rho}}^{\rm T}{\bm{W}}_{12}^{y}-{\bm{w}}_{v_{22}}{\bm{\rho}}^{\rm T}\breve{{\bm{W}}}_{11}^{y}){\bm{\rho}}\big{]},$ (21b)
-$u_{22}=\frac{1}{|{\bm{Y}}|}\big{[}{\bm{\rho}}^{\rm T}({\bm{w}}_{p_{22}}{\bm{\rho}}^{\rm T}{\bm{W}}_{y}-{\bm{w}}_{v_{21}}{\bm{\rho}}^{\rm T}{\bm{W}}_{22}^{y}-{\bm{w}}_{v_{22}}{\bm{\rho}}^{\rm T}\breve{{\bm{W}}}_{21}^{y}){\bm{\rho}}\big{]},$ (21c)
-where $|{\bm{Y}}|={\bm{\rho}}^{\rm T}{\bm{W}}_{y}{\bm{\rho}}$. Elements $u_{ij}$ are respectively given by,
-$\begin{cases}{\bm{W}}_{y}={\bm{w}}_{y_{11}}{\bm{w}}_{y_{22}}^{\rm T}-{\bm{w}}_{y_{12}}{\bm{w}}_{y_{21}}^{\rm T},\\ {\bm{W}}^{y}_{12}={\bm{w}}_{v_{11}}{\bm{w}}_{y_{22}}^{\rm T}-{\bm{w}}_{v_{12}}{\bm{w}}_{y_{21}}^{\rm T},\\ {\bm{W}}^{y}_{22}={\bm{w}}_{v_{21}}{\bm{w}}_{y_{22}}^{\rm T}-{\bm{w}}_{v_{22}}{\bm{w}}_{y_{21}}^{\rm T},\\ \breve{\bm{W}}^{y}_{11}={\bm{w}}_{v_{12}}{\bm{w}}_{y_{11}}^{\rm T}-{\bm{w}}_{v_{11}}{\bm{w}}_{y_{12}}^{\rm T},\\ \breve{\bm{W}}^{y}_{21}={\bm{w}}_{v_{22}}{\bm{w}}_{y_{11}}^{\rm T}-{\bm{w}}_{v_{21}}{\bm{w}}_{y_{12}}^{\rm T}.\end{cases}$ (22)
-Similarly, the velocity CRLB matrix can be rewritten as,
-${\bm{C}}_{\rm V}=\frac{1}{|{\bm{H}}|}\begin{bmatrix}h_{22}&-h_{12}\\ -h_{21}&h_{11}\end{bmatrix},$ (23)
-where $|{\bm{H}}|=h_{11}h_{22}-h_{12}h_{21}$. Elements $h_{ij}$ are respectively given by,
-$h_{11}=\frac{1}{|{\bm{P}}|}\big{[}{\bm{\rho}}^{\rm T}({\bm{w}}_{y_{11}}{\bm{\rho}}^{\rm T}{\bm{W}}_{p}-{\bm{w}}_{v_{11}}{\bm{\rho}}^{\rm T}{\bm{W}}_{12}^{p}-{\bm{w}}_{v_{21}}{\bm{\rho}}^{\rm T}\breve{{\bm{W}}}_{11}^{p}){\bm{\rho}}\big{]},$ (24a)
-$h_{12}=\frac{1}{|{\bm{P}}|}\big{[}{\bm{\rho}}^{\rm T}({\bm{w}}_{y_{12}}{\bm{\rho}}^{\rm T}{\bm{W}}_{p}-{\bm{w}}_{v_{12}}{\bm{\rho}}^{\rm T}{\bm{W}}_{12}^{p}-{\bm{w}}_{v_{22}}{\bm{\rho}}^{\rm T}\breve{{\bm{W}}}_{11}^{p}){\bm{\rho}}\big{]},$ (24b)
-$h_{22}=\frac{1}{|{\bm{P}}|}\big{[}{\bm{\rho}}^{\rm T}({\bm{w}}_{y_{22}}{\bm{\rho}}^{\rm T}{\bm{W}}_{p}-{\bm{w}}_{v_{12}}{\bm{\rho}}^{\rm T}{\bm{W}}_{22}^{p}-{\bm{w}}_{v_{22}}{\bm{\rho}}^{\rm T}\breve{{\bm{W}}}_{21}^{p}){\bm{\rho}}\big{]},$ (24c)
+$\small{\bm{C}}_{\rm V}$ $=({\bm{Y}}-{\bm{V}}^{\rm T}{\bm{P}}^{-1}{\bm{V}})^{-1}.$ (12)
+The detailed expression of ${\bm{P}}$ and ${\bm{V}}$ are given at the top of the page, respectively, and
+${\bm{P}}=\sum_{n=1}^{N}\sum_{k=1}^{K}\begin{bmatrix}\beta_{n,k}^{2}a_{n,k}+2\beta_{n,k}\eta_{n,k}b_{n,k}+\eta_{n,k}^{2}d_{n,k}&\beta_{n,k}\zeta_{n,k}a_{n,k}\!+\!(\beta_{n,k}\kappa_{n,k}\!+\!\zeta_{n,k}\eta_{n,k})b_{n,k}+\eta_{n,k}\kappa_{n,k}d_{n,k}\\ \zeta_{n,k}\beta_{n,k}a_{n,k}\!+\!(\zeta_{n,k}\eta_{n,k}+\beta_{n,k}\kappa_{n,k})b_{n,k}+\kappa_{n,k}\eta_{n,k}d_{n,k}&\zeta_{n,k}^{2}a_{n,k}+2\zeta_{n,k}\kappa_{n,k}b_{n,k}\!+\!\kappa_{n,k}^{2}d_{n,k}\end{bmatrix},$
+$\small{\bm{V}}=\sum_{n=1}^{N}\sum_{k=1}^{K}\begin{bmatrix}\beta_{n,k}\xi_{n,k}b_{n,k}+\eta_{n,k}\xi_{n,k}d_{n,k}&\beta_{n,k}\varrho_{n,k}b_{n,k}+\eta_{n,k}\varrho_{n,k}d_{n,k}\\ \zeta_{n,k}\xi_{n,k}b_{n,k}+\kappa_{n,k}\xi_{n,k}d_{n,k}&\zeta_{n,k}\varrho_{n,k}b_{n,k}+\kappa_{n,k}\varrho_{n,k}d_{n,k}\end{bmatrix},$
+$\small{\bm{Y}}=\sum_{n=1}^{N}\sum_{k=1}^{K}\begin{bmatrix}\xi_{n,k}^{2}d_{n,k}&\xi_{n,k}\varrho_{n,k}d_{n,k}\\ \varrho_{n,k}\xi_{n,k}d_{n,k}&\varrho_{n,k}^{2}d_{n,k}\end{bmatrix}.$ (13)
+The relevant waveform-related parameters $a_{n,k}$, $b_{n,k}$ and $d_{n,k}$, and geometric-spreading-related factors $\beta_{n,k}$, $\zeta_{n,k}$, $\xi_{n,k}$, $\varrho_{n,k}$, $\eta_{n,k}$, and $\kappa_{n,k}$, are detailed in Appendix A.
+As observed from (61), the PA coefficients play a pivotal role in the formulation of the CRLBs.
+Meanwhile, these coefficients directly affect communication SINR, as described in (7).
+Unfortunately, the PA optimization relevant to the CRLB is intractable due to the implicit and complex forms of the expressions in (11) and (12).
+To facilitate analysis and optimization, we reformulate the CRLB expressions into a more explicit and tractable form.
+We denote the $(i,j)$th elements of matrices ${\bm{P}}$, ${\bm{V}}$ and ${\bm{Y}}$ as $p_{ij}$, $v_{ij}$ and $y_{ij}$, respectively. Then, one can simplify them as
+$\small\begin{cases}p_{ij}=\sum_{n=1}^{N}w_{n}^{p_{ij}}\rho_{n}={\bm{\rho}}^{\rm T}{\bm{w}}_{p_{ij}},\\ v_{ij}=\sum_{n=1}^{N}w_{n}^{v_{ij}}\rho_{n}={\bm{\rho}}^{\rm T}{\bm{w}}_{v_{ij}},\\ y_{ij}=\sum_{n=1}^{N}w_{n}^{y_{ij}}\rho_{n}={\bm{\rho}}^{\rm T}{\bm{w}}_{y_{ij}},\end{cases}$ (14)
+where $w_{n}^{p_{ij}}$, $w_{n}^{v_{ij}}$ and $w_{n}^{y_{ij}}$ are the $n$th element of the weight vector ${\bm{w}}_{p_{ij}}$, ${\bm{w}}_{v_{ij}}$ and ${\bm{w}}_{y_{ij}}$, respectively.
+These weights can be readily obtained from the definitions of ${\bm{P}}$, ${\bm{V}}$ and ${\bm{Y}}$, by extracting their PA coefficient $\rho_{n}$.
+For example, from (13), we have $w_{n}^{p_{11}}$ as
+$\small w_{n}^{p_{11}}=\sum\nolimits_{k=1}^{K}(\beta_{n,k}^{2}\bar{a}_{n,k}+2\beta_{n,k}\eta_{n,k}\bar{b}_{n,k}+\eta_{n,k}^{2}\bar{d}_{n,k}),$ (15)
+where the details of $\bar{a}_{n,k}$, $\bar{b}_{n,k}$ and $\bar{d}_{n,k}$ are referred to (61).
+By substituting (14) into (11), the CRLB matrix for location estimation can be rewritten as
+$\small{\bm{C}}_{\rm L}=\frac{1}{|{\bm{U}}|}\begin{bmatrix}u_{22}&-u_{12}\\ -u_{21}&u_{11}\end{bmatrix},$ (16)
+where $|{\bm{U}}|=u_{11}u_{22}-u_{12}u_{21}$. The relevant elements $u_{ij}$ are given respectively by
+$\small u_{11}=\frac{1}{|{\bm{Y}}|}\big{[}{\bm{\rho}}^{\rm T}({\bm{w}}_{p_{11}}{\bm{\rho}}^{\rm T}{\bm{W}}_{y}\!-\!{\bm{w}}_{v_{11}}{\bm{\rho}}^{\rm T}{\bm{W}}_{12}^{y}\!-\!{\bm{w}}_{v_{12}}{\bm{\rho}}^{\rm T}\breve{{\bm{W}}}_{11}^{y}){\bm{\rho}}\big{]},$ (17)
+$u_{12}=\frac{1}{|{\bm{Y}}|}\big{[}{\bm{\rho}}^{\rm T}({\bm{w}}_{p_{12}}{\bm{\rho}}^{\rm T}{\bm{W}}_{y}\!-\!{\bm{w}}_{v_{21}}{\bm{\rho}}^{\rm T}{\bm{W}}_{12}^{y}\!-\!{\bm{w}}_{v_{22}}{\bm{\rho}}^{\rm T}\breve{{\bm{W}}}_{11}^{y}){\bm{\rho}}\big{]},$
+$u_{22}=\frac{1}{|{\bm{Y}}|}\big{[}{\bm{\rho}}^{\rm T}({\bm{w}}_{p_{22}}{\bm{\rho}}^{\rm T}{\bm{W}}_{y}\!-\!{\bm{w}}_{v_{21}}{\bm{\rho}}^{\rm T}{\bm{W}}_{22}^{y}\!-\!{\bm{w}}_{v_{22}}{\bm{\rho}}^{\rm T}\breve{{\bm{W}}}_{21}^{y}){\bm{\rho}}\big{]},$
+where $|{\bm{Y}}|={\bm{\rho}}^{\rm T}{\bm{W}}_{y}{\bm{\rho}}$, and
+$\small\begin{cases}{\bm{W}}_{y}={\bm{w}}_{y_{11}}{\bm{w}}_{y_{22}}^{\rm T}-{\bm{w}}_{y_{12}}{\bm{w}}_{y_{21}}^{\rm T},\\ {\bm{W}}^{y}_{12}={\bm{w}}_{v_{11}}{\bm{w}}_{y_{22}}^{\rm T}-{\bm{w}}_{v_{12}}{\bm{w}}_{y_{21}}^{\rm T},\\ {\bm{W}}^{y}_{22}={\bm{w}}_{v_{21}}{\bm{w}}_{y_{22}}^{\rm T}-{\bm{w}}_{v_{22}}{\bm{w}}_{y_{21}}^{\rm T},\\ \breve{\bm{W}}^{y}_{11}={\bm{w}}_{v_{12}}{\bm{w}}_{y_{11}}^{\rm T}-{\bm{w}}_{v_{11}}{\bm{w}}_{y_{12}}^{\rm T},\\ \breve{\bm{W}}^{y}_{21}={\bm{w}}_{v_{22}}{\bm{w}}_{y_{11}}^{\rm T}-{\bm{w}}_{v_{21}}{\bm{w}}_{y_{12}}^{\rm T}.\end{cases}$ (18)
+Similarly, the velocity CRLB matrix can be rewritten as
+$\small{\bm{C}}_{\rm V}=\frac{1}{|{\bm{H}}|}\begin{bmatrix}h_{22}&-h_{12}\\ -h_{21}&h_{11}\end{bmatrix},$ (19)
+where $|{\bm{H}}|=h_{11}h_{22}-h_{12}h_{21}$. The relevant elements $h_{ij}$ are given respectively by
+$\small h_{11}=\frac{1}{|{\bm{P}}|}\big{[}{\bm{\rho}}^{\rm T}({\bm{w}}_{y_{11}}{\bm{\rho}}^{\rm T}{\bm{W}}_{p}\!-\!{\bm{w}}_{v_{11}}{\bm{\rho}}^{\rm T}{\bm{W}}_{12}^{p}\!-\!{\bm{w}}_{v_{21}}{\bm{\rho}}^{\rm T}\breve{{\bm{W}}}_{11}^{p}){\bm{\rho}}\big{]},$ (20)
+$h_{12}=\frac{1}{|{\bm{P}}|}\big{[}{\bm{\rho}}^{\rm T}({\bm{w}}_{y_{12}}{\bm{\rho}}^{\rm T}{\bm{W}}_{p}\!-\!{\bm{w}}_{v_{12}}{\bm{\rho}}^{\rm T}{\bm{W}}_{12}^{p}\!-\!{\bm{w}}_{v_{22}}{\bm{\rho}}^{\rm T}\breve{{\bm{W}}}_{11}^{p}){\bm{\rho}}\big{]},$
+$h_{22}=\frac{1}{|{\bm{P}}|}\big{[}{\bm{\rho}}^{\rm T}({\bm{w}}_{y_{22}}{\bm{\rho}}^{\rm T}{\bm{W}}_{p}\!-\!{\bm{w}}_{v_{12}}{\bm{\rho}}^{\rm T}{\bm{W}}_{22}^{p}\!-\!{\bm{w}}_{v_{22}}{\bm{\rho}}^{\rm T}\breve{{\bm{W}}}_{21}^{p}){\bm{\rho}}\big{]},$
 where $|{\bm{P}}|={\bm{\rho}}^{\rm T}{\bm{W}}_{p}{\bm{\rho}}$, and
-$\begin{cases}{\bm{W}}_{p}={\bm{w}}_{p_{11}}{\bm{w}}_{p_{22}}^{\rm T}-{\bm{w}}_{p_{12}}{\bm{w}}_{p_{21}}^{\rm T},\\ {\bm{W}}^{p}_{12}={\bm{w}}_{v_{11}}{\bm{w}}_{p_{22}}^{\rm T}-{\bm{w}}_{v_{21}}{\bm{w}}_{p_{21}}^{\rm T},\\ {\bm{W}}^{p}_{22}={\bm{w}}_{v_{12}}{\bm{w}}_{p_{22}}^{\rm T}-{\bm{w}}_{v_{22}}{\bm{w}}_{p_{21}}^{\rm T},\\ \breve{\bm{W}}^{p}_{11}={\bm{w}}_{v_{21}}{\bm{w}}_{p_{11}}^{\rm T}-{\bm{w}}_{v_{11}}{\bm{w}}_{p_{12}}^{\rm T},\\ \breve{\bm{W}}^{p}_{21}={\bm{w}}_{v_{22}}{\bm{w}}_{p_{11}}^{\rm T}-{\bm{w}}_{v_{12}}{\bm{w}}_{p_{12}}^{\rm T}.\end{cases}$ (25)
-IV Problem Formulation
-In general, a higher power budget leads to improved sensing and communication performance.
-This section investigates a novel PA scheme aimed at maximizing the overall performance of CF MIMO-ISAC systems.
-Depending on specific application requirements, the optimization objective can be either sensing or communication performance, while the quality of service (QoS) requirement of the other function is incorporated as a constraint in the optimization model.
-Let ${\bm{\rho}}_{\min}=[\rho_{\min,1},\rho_{\min,2},\cdots,\rho_{\min,N}]^{\rm T}$ denote the minimum PA coefficient to satisfy the so-called larger SNR for tight CRLBs, and ${\bm{\rho}}_{\max}=[\rho_{\max,1},\rho_{\max,2},\cdots,\rho_{\max,N}]^{\rm T}$ denote the maximum PA coefficient due to the power budgets of individual transmitters [40, 13].
-In this paper, we consider maximizing the communication data rate or SINR under the constraint of the sensing performance, i.e.,
-$\arg\min\limits_{{\bm{\rho}}}$ $\ -{\bm{\rho}}^{\rm T}{\bm{g}},$ (26a)
-${\rm s.t.}$ ${\bm{\rho}}^{\rm T}{\bm{1}}=1,$ (26b)
+$\small\begin{cases}{\bm{W}}_{p}={\bm{w}}_{p_{11}}{\bm{w}}_{p_{22}}^{\rm T}-{\bm{w}}_{p_{12}}{\bm{w}}_{p_{21}}^{\rm T},\\ {\bm{W}}^{p}_{12}={\bm{w}}_{v_{11}}{\bm{w}}_{p_{22}}^{\rm T}-{\bm{w}}_{v_{21}}{\bm{w}}_{p_{21}}^{\rm T},\\ {\bm{W}}^{p}_{22}={\bm{w}}_{v_{12}}{\bm{w}}_{p_{22}}^{\rm T}-{\bm{w}}_{v_{22}}{\bm{w}}_{p_{21}}^{\rm T},\\ \breve{\bm{W}}^{p}_{11}={\bm{w}}_{v_{21}}{\bm{w}}_{p_{11}}^{\rm T}-{\bm{w}}_{v_{11}}{\bm{w}}_{p_{12}}^{\rm T},\\ \breve{\bm{W}}^{p}_{21}={\bm{w}}_{v_{22}}{\bm{w}}_{p_{11}}^{\rm T}-{\bm{w}}_{v_{12}}{\bm{w}}_{p_{12}}^{\rm T}.\end{cases}$ (21)
+Note that the newly derived CRLB expressions in (16) and (19) represent explicit functions of the PA vector, thereby serving as effective and tractable sensing performance metrics.
+III-B Problem Formulation
+This section formulates the PA problem by maximizing the overall performance of CF MIMO-ISAC systems.
+Depending on specific application requirements, the optimization objective can be either sensing or communication performance, while the requirement of the other functionality is incorporated as a constraint in the optimization model.
+Let ${\bm{\rho}}_{\min}=[\rho_{\min,1},\rho_{\min,2},\cdots,\rho_{\min,N}]^{\rm T}$ denote the minimum PA coefficient to achieve the sufficiently large SNR for tight CRLBs [34, 36], and ${\bm{\rho}}_{\max}=[\rho_{\max,1},\rho_{\max,2},\cdots,\rho_{\max,N}]^{\rm T}$ denote the maximum PA coefficient due to the power budgets of individual transmitters [38, 36].
+In this paper, we consider the problem of maximizing the communication SINR, subject to constraints on the sensing performance, i.e.,
+$\small\arg\min\limits_{{\bm{\rho}}}$ $\ -{\bm{\rho}}^{\rm T}{\bm{g}},$ (22a)
+${\rm s.t.}$ ${\bm{1}}_{N}^{\rm T}{\bm{\rho}}=1,$ (22b)
 ${\bm{\rho}}_{\min}\leq{\bm{\rho}}\leq{\bm{\rho}}_{\max},$ (22c)
@@
 ${\rm tr}\{{\bm{C}}_{\rm V}\}\leq\delta_{v}^{2},$ (22e)
-where $\delta_{l}^{2}$ and $\delta_{v}^{2}$ denote the corresponding performance thresholds  [^3] The thresholds can be set by referring to the CRLBs under a uniform power allocation..
-Similar to the approaches in [40, 49, 11, 13], we assume that the target’s RCS, location, and velocity have been coarsely estimated by the ISAC system and are periodically updated over time [^4] The previously estimated target parameters can be reused for PA optimization at the current time step, particularly in consecutive estimation or target tracking scenarios..
-Therefore, the PA coefficient vector $\bm{\rho}$ becomes the sole optimization variable, while the other relevant parameters are implicitly embedded through their coupling in equations (22) and (25).
-V Algorithm Derivation
+where $\delta_{l}^{2}$ and $\delta_{v}^{2}$ denote the corresponding location and velocity CRLB thresholds.
+IV Power Allocation Algorithm for CF MIMO-ISAC
 In this section, we investigate the optimal PA strategy for maximizing the ISAC performance, as formulated in (22a)-(22e).
-However, the strong nonlinearity and non-convexity of the CRLB-based constraints in (26d) and (26e) render the problem analytically intractable.
-To overcome this challenge, we adopt a conjugate gradient (CG) iterative algorithm, enhanced with a penalty function and projection operator, to provide an efficient numerical solution.
-We first introduce a penalty (function) for violating the nonlinear inequality constraints in (26c)-(26e)
-$\alpha({\bm{\rho}})$ $=\tilde{f}_{l}({\bm{\rho}})+\tilde{f}_{v}({\bm{\rho}})+\sum_{n=1}^{N}\left(f_{\min,n}({\bm{\rho}})+f_{\max,n}({\bm{\rho}})\right),$ (27)
+Unfortunately, the strong nonlinearity and non-convexity of the CRLB-based constraints in (22d) and (22e) render the problem analytically intractable.
+To overcome this challenge, we adopt a CG iterative algorithm, enhanced with a penalty function and projection operator, to provide an efficient numerical solution.
+To derive a numerical solution, we start from converting the original constrained optimization problem into an unconstrained one.
+To this end, we first introduce a penalty (function) for violating the nonlinear inequality constraints in (22c)-(22e), defined as [39]
+$\alpha({\bm{\rho}})=\tilde{f}_{l}({\bm{\rho}})+\tilde{f}_{v}({\bm{\rho}})+\sum\nolimits_{n=1}^{N}\left(f_{\min,n}({\bm{\rho}})+f_{\max,n}({\bm{\rho}})\right),$
+where the involved penalty sub-functions are given respectively by
+$\small\begin{cases}\tilde{f}_{l}({\bm{\rho}})=(\max\{0,{f}_{l}({\bm{\rho}})\})^{q},\\ \tilde{f}_{v}({\bm{\rho}})=(\max\{0,{f}_{v}({\bm{\rho}})\})^{q},\\ f_{\min,n}({\bm{\rho}})=(\max\{0,\rho_{\min,n}-\rho_{n}\})^{q},\\ f_{\max,n}({\bm{\rho}})=(\max\{0,\rho_{n}-\rho_{\max,n}\})^{q},\end{cases}$ (23)
+with ${f}_{l}({\bm{\rho}})\triangleq{\rm tr}\{{\bm{C}}_{\rm L}\}-\delta_{l}^{2}$ and ${f}_{v}({\bm{\rho}})\triangleq{\rm tr}\{{\bm{C}}_{\rm V}\}-\delta_{v}^{2}$.
+$q$ is the power of the penalty function.
+We consider $q=2$ to guarantee their continuous differentiability.
+Then, the objective function for the penalty problem can be written as
+$\small L({\bm{\rho}})=-{\bm{\rho}}^{\rm T}{\bm{g}}+\mu\alpha({\bm{\rho}}),$ (24)
+where $\mu$ denotes the penalty factor.
+Additionally, to enforce this equality constraint (22b) and reduce both computational complexity and the nonlinearity of the optimization problem, we adopt a projection operator instead of incorporating another penalty sub-function, i.e.,
+$\small{\bm{\rho}}({\bm{\varpi}})={\bm{\Theta}}{\bm{\varpi}}+{\bm{\rho}}_{0},$ (25)
+where ${\bm{\varpi}}\in\mathbb{R}^{N\times 1}$ is an arbitrary vector and ${\bm{\Theta}}\triangleq{\bm{I}}_{N\times N}-\frac{{\bm{1}}_{N}{\bm{1}}_{N}^{\rm T}}{N}$ is a projection matrix onto the zero space of ${\bm{1}}_{N}$.
+We have ${\bm{1}}_{N}^{\rm T}{\bm{\Theta}}={\bm{0}}_{N}^{\rm T}$.
+Thus, the equality constraint ${\bm{1}}_{N}^{\rm T}{\bm{\rho}}({\bm{\varpi}})=1$ in (22b) always holds by initializing ${\bm{\rho}}_{0}$ such that ${\bm{1}}_{N}^{\rm T}{\bm{\rho}}_{0}=1$.
+Consequently, ${\bm{\rho}}({\bm{\varpi}})$ is the sum of a constant PA vector ${\bm{\rho}}_{0}$ and a projection of ${\bm{\varpi}}$ onto the zero space of ${\bm{1}}_{N}$.
+By substituting (25) into (24), we can convert the original constrained optimization to an unconstrained optimization problem as follows:
+$\small\arg\min\limits_{{\bm{\rho}}}$ $\ \mathcal{L}({\bm{\varpi}})=-{\bm{\rho}}^{\rm T}({\bm{\varpi}}){\bm{g}}+\mu\alpha({\bm{\rho}}({\bm{\varpi}})).$ (26)
+Then, unconstrained optimization methods, e.g., CG and SD methods, can be employed to search for the optimal solution.
+IV-A Modified Conjugate Gradient Algorithm
+To numerically solve (26), we propose a projection-based MCG algorithm. The update at the $i$th iteration is given by
+$\small{\bm{\varpi}}_{i+1}={\bm{\varpi}}_{i}+\upsilon_{i}{\bm{d}}_{i},$ (27)
+where ${\bm{d}}_{i}$ is the search direction of $\mathcal{L}({\bm{\varpi}}_{i})$ and $\upsilon_{i}$ is the step size. The search direction using the MCG is updated by
+${\bm{d}}_{i+1}$ $=-\nabla\mathcal{L}({\bm{\varpi}}_{i+1})+\varsigma_{i}{\bm{d}}_{i},$ (28)
+where $\varsigma_{i}$ denotes the deflection factor of a CG algorithm. Taking the derivative $\mathcal{L}({\bm{\varpi}})$ regarding ${\bm{\varpi}}$ yields
+(ϖ)=bold-italic-$\small\nabla\mathcal{L}({\bm{\varpi}})=$ ${\bm{\Theta}}\nabla L({\bm{\rho}})={\bm{\Theta}}\big{[}{\bm{g}}+\nabla\alpha({\bm{\rho}})\big{]},$ (29)
 where
-$\begin{cases}\tilde{f}_{l}({\bm{\rho}})=(\max\{0,{f}_{l}({\bm{\rho}})\})^{q},\\ \tilde{f}_{v}({\bm{\rho}})=(\max\{0,{f}_{v}({\bm{\rho}})\})^{q},\\ f_{\min,n}({\bm{\rho}})=(\max\{0,\rho_{\min,n}-\rho_{n}\})^{q},\\ f_{\max,n}({\bm{\rho}})=(\max\{0,\rho_{n}-\rho_{\max,n}\})^{q},\end{cases}$ (28)
-with ${f}_{l}({\bm{\rho}})\triangleq{\rm tr}\{{\bm{C}}_{\rm L}\}-\delta_{l}^{2}$ and ${f}_{v}({\bm{\rho}})\triangleq{\rm tr}\{{\bm{C}}_{\rm V}\}-\delta_{v}^{2}$.
-We consider $q=2$ to guarantee the continuous differentiable penalty functions.
-Then, the objective function for the penalty problem can be written as
-$L({\bm{\rho}})=-{\bm{\rho}}^{\rm T}{\bm{g}}+\mu\alpha({\bm{\rho}}).$ (29)
-As mentioned in [54], as the penalty factor $\mu$ is made large, the points generated for minimizing (29) approach an optimal solution of the original objective function (57a) from outside the feasible region and the iteration process terminate with $\mu\alpha({\bm{\rho}})\leq\epsilon_{\mu}$ for a small threshold $\epsilon_{\mu}>0$.
-This issue can be solved by setting $\epsilon_{\mu}$ small enough or employing much tighter thresholds than the required bounds to some extent to mitigate the loss due to the non-zero threshold $\epsilon_{\mu}$.
-On the other hand, intuitively, when orthogonal waveforms are employed, increasing the transmission power—or equivalently, the signal-to-noise ratio (SNR)—at any individual transmitter enhances both sensing and communication performance.
-Therefore, it is desirable to fully utilize the total power budget, as specified in (26b).
-To enforce this equality constraint and reduce both computational complexity and the nonlinearity of the optimization problem, we adopt a projection operator instead of incorporating another penalty sub-function
-${\bm{\rho}}({\bm{\varpi}})={\bm{\Theta}}{\bm{\varpi}}+{\bm{\rho}}_{0},$ (30)
-where ${\bm{\Theta}}\triangleq{\bm{I}}_{N\times N}-\frac{{\bm{1}}_{N}{\bm{1}}_{N}^{\rm T}}{N}$ is a projection matrix onto the zero space of ${\bm{1}}_{N}$.
-It is evident that ${\bm{1}}_{N}^{\rm T}{\bm{\rho}}({\bm{\varpi}})=1$ holds with ${\bm{1}}_{N}^{\rm T}{\bm{\rho}}_{0}=1$.
-Thus, ${\bm{\rho}}({\bm{\varpi}})$ is the sum of a constant PA vector ${\bm{\rho}}_{0}$ and a projection of any $N\times 1$ vector ${\bm{\varpi}}$ onto the zero space of ${\bm{1}}_{N}$.
-According to (30), the objective function can be further written as,
-$\mathcal{L}({\bm{\varpi}})=-{\bm{\rho}}^{\rm T}({\bm{\varpi}}){\bm{g}}+\mu\alpha({\bm{\rho}}({\bm{\varpi}})).$ (31)
-This objective function makes all the constraints satisfied and thus convert the original constrained optimization regarding ${\bm{\rho}}$ into an unconstrained one regarding ${\bm{\varpi}}$.
-V-A Modified Conjugate Gradient Algorithm
-We now propose a projection based modified CG (MCG) algorithm to minimize (31). Taking the derivative $\mathcal{L}({\bm{\varpi}})$ regarding ${\bm{\varpi}}$ yields
-(ϖ)=bold-italic-$\nabla\mathcal{L}({\bm{\varpi}})=$ ${\bm{\Theta}}\nabla L({\bm{\rho}})={\bm{\Theta}}\big{[}{\bm{g}}+\nabla\alpha({\bm{\rho}})\big{]}.$ (32)
-where
-$\nabla\alpha({\bm{\rho}})=$ $\mu\big{(}\nabla\tilde{f}_{l}({\bm{\rho}})+\nabla\tilde{f}_{v}({\bm{\rho}})$
-$+\sum_{n=1}^{N}(\nabla f_{\min,n}({\bm{\rho}})+\nabla f_{\max,n}({\bm{\rho}}))\big{)},$ (33)
-with components
-$\begin{cases}\nabla\tilde{f}_{l}({\bm{\rho}})=2\max\{f_{l}({\bm{\rho}}),0\}\nabla f_{l}({\bm{\rho}})\\ \nabla\tilde{f}_{v}({\bm{\rho}})=2\max\{f_{v}({\bm{\rho}}),0\}\nabla f_{v}({\bm{\rho}})\\ \nabla f_{\min,n}({\bm{\rho}})=2\min\{\rho_{n}-\rho_{\min,n},0\}{\bm{e}}_{n}\\ \nabla f_{\max,n}({\bm{\rho}})=2\max\{\rho_{n}-\rho_{\max,n},0\}{\bm{e}}_{n}.\end{cases}$ (34)
-Note that ${\bm{e}}_{n}$ is an index vector with its element at index $n$ being $1$ and the others being zeros.
-Solving (31) using an iterative method yields the update
-${\bm{\varpi}}_{i+1}={\bm{\varpi}}_{i}+\upsilon_{i}{\bm{d}}_{i},$ (35)
-where ${\bm{d}}_{i}$ is the descent direction of $\mathcal{L}({\bm{\varpi}}_{i})$ and $\upsilon_{i}$ is the step size. The descent direction using the MCG is updated by
-${\bm{d}}_{i+1}$ $=-{\bm{\Theta}}\nabla L({\bm{\rho}}_{i+1})+\varsigma_{i}{\bm{d}}_{i},$ (36)
-where $\varsigma_{i}$ denotes the deflection factor that characterizes a particular CG algorithm.
-Note that $-{\bm{\Theta}}\nabla L({\bm{\rho}}_{i})$ is still the descent direction of $L({\bm{\rho}}_{i})$ with $-\nabla L({\bm{\rho}}_{i})^{\rm T}{\bm{\Theta}}\nabla L({\bm{\rho}}_{i})\leq 0$ since the projection operator ${\bm{\Theta}}$ is a symmetric positive definite matrix.
+$\small\nabla\alpha({\bm{\rho}})=$ $\mu\big{(}\nabla\tilde{f}_{l}({\bm{\rho}})+\nabla\tilde{f}_{v}({\bm{\rho}})$
+$+\sum\nolimits_{n=1}^{N}(\nabla f_{\min,n}({\bm{\rho}})+\nabla f_{\max,n}({\bm{\rho}}))\big{)},$ (30)
+and
+$\small\begin{cases}\nabla\tilde{f}_{l}({\bm{\rho}})=2\max\{f_{l}({\bm{\rho}}),0\}\nabla f_{l}({\bm{\rho}})\\ \nabla\tilde{f}_{v}({\bm{\rho}})=2\max\{f_{v}({\bm{\rho}}),0\}\nabla f_{v}({\bm{\rho}})\\ \nabla f_{\min,n}({\bm{\rho}})=2\min\{\rho_{n}-\rho_{\min,n},0\}{\bm{e}}_{n}\\ \nabla f_{\max,n}({\bm{\rho}})=2\max\{\rho_{n}-\rho_{\max,n},0\}{\bm{e}}_{n}.\end{cases}$ (31)
+Note that ${\bm{e}}_{n}\in\mathbb{R}^{N\times 1}$ is an index vector with its element at index $n$ being $1$ and the others being zeros.
 Given the initial direction ${\bm{d}}_{1}=-{\bm{\Theta}}\nabla L({\bm{\rho}}_{1})$, and using the idempotent property of the projection matrix, i.e., ${\bm{\Theta}}^{2}={\bm{\Theta}}$, we have ${\bm{\Theta}}{\bm{d}}_{i}={\bm{d}}_{i}$, $\forall{i>0}$.
-Then, taking the projection of (35) leads to
-${\bm{\rho}}_{i+1}={\bm{\rho}}_{i}+\upsilon_{i}{\bm{d}}_{i}.$ (37)
-which has the same form as (35).
-Therefore, we hereafter directly adopt the update in (37).
-Moreover, note that the equality constraint ${\bm{1}}_{N}^{\rm T}{\bm{\rho}}_{i}=1$ holds for all $i>1$ if and only if ${\bm{1}}_{N}^{\rm T}{\bm{\rho}}_{1}=1$.
-This property follows from the iteration rules in (36) and (37).
-To enable the convergence, it is necessary to develop a suitable step size $\upsilon_{i}$ and search direction ${\bm{d}}_{i}$ in each iteration to guarantee the non-increasing objective function i.e.,
-$L({\bm{\rho}}_{i}+\upsilon_{i}{\bm{d}}_{i})-L({\bm{\rho}}_{i})\leq 0.$ (38)
-Firstly, for the non-increasing search direction for each iteration $i$, it requires
-${\bm{d}}_{i+1}^{\rm T}\nabla L({\bm{\rho}}_{i+1})\leq 0,$ (39)
-where the equality holds if and only if $\nabla L({\bm{\rho}}_{i+1})={\bm{0}}$. Using (36), we have,
-$\begin{cases}\varsigma_{i}\leq\varsigma_{i}^{\rm b},\ \varsigma_{i}^{\rm b}>0\\ \varsigma_{i}\geq\varsigma_{i}^{\rm b},\ \varsigma_{i}^{\rm b}<0\end{cases}$ (40)
-where $\varsigma_{i}^{\rm b}$ is the bound for $\varsigma_{i}$
-$\varsigma_{i}^{\rm b}=\dfrac{\nabla L({\bm{\rho}}_{i+1})^{\rm T}{\bm{\Theta}}\nabla L({\bm{\rho}}_{i+1})}{{\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i+1})}.$ (41)
-Note that $\nabla L({\bm{\rho}}_{i+1})^{\rm T}{\bm{\Theta}}\nabla L({\bm{\rho}}_{i+1})>0$ for any non-zero gradient $\nabla L({\bm{\rho}}_{i+1})$ due to the positive definite ${\bm{\Theta}}$.
-When considering inexact line searches, Hestenes and Stiefel’s [55, 56] choice for $\varsigma_{i}$ can be preferable. We have the modified HS deflection,
-$\varsigma_{i}^{\rm HS}=\dfrac{\nabla L({\bm{\rho}}_{i+1})^{\rm T}{\bm{q}}_{i}}{{\bm{d}}_{i}^{\rm T}{\bm{q}}_{i}},$ (42)
+Thus, taking the projection of (27) leads to
+$\small{\bm{\rho}}_{i+1}={\bm{\rho}}_{i}+\upsilon_{i}{\bm{d}}_{i},$ (32)
+where the search direction is updated by
+$\small{\bm{d}}_{i+1}$ $=-{\bm{\Theta}}\nabla\mathcal{L}({\bm{\rho}}_{i+1})+\varsigma_{i}{\bm{d}}_{i}.$ (33)
+We have ${\bm{1}}_{N}^{\rm T}{\bm{d}}_{i}=0$ for $\forall{i>1}$ due to ${\bm{1}}_{N}^{\rm T}{\bm{d}}_{1}=0$ and ${\bm{1}}_{N}^{\rm T}{\bm{\Theta}}={\bm{0}}_{N}^{\rm T}$.
+Therefore, the equality constraint in (22b), that is, ${\bm{1}}_{N}^{\rm T}{\bm{\rho}}_{i}=1$, always holds at $i$th iteration for $\forall{i>1}$ if and only if the initial PA vector ${\bm{\rho}}_{1}$ satisfies ${\bm{1}}_{N}^{\rm T}{\bm{\rho}}_{1}=1$.
+We now investigate the selection of the deflection factor $\varsigma_{i}$ and step size $\upsilon_{i}$ at each iteration, which is significant in guaranteeing the convergence of the MCG algorithm.
+One necessary condition for the convergence is the non-increasing objective function at each iteration, i.e.,
+$\small L({\bm{\rho}}_{i}+\upsilon_{i}{\bm{d}}_{i})-L({\bm{\rho}}_{i})\leq 0.$ (34)
+To this end, the search direction ${\bm{d}}_{i}$ should be the descent direction and the step size $\upsilon_{i}$ should be correspondingly upper bounded by a threshold at each iteration.
+IV-A1 Descent direction
+To ensure a descent direction ${\bm{d}}_{i+1}$ at the $(i+1)$th iteration, it requires
+$\small{\bm{d}}_{i+1}^{\rm T}\nabla L({\bm{\rho}}_{i+1})\leq 0,$ (35)
+where the equality holds if and only if $\nabla L({\bm{\rho}}_{i+1})={\bm{0}}$.
+The initial direction ${\bm{d}}_{1}=-{\bm{\Theta}}\nabla L({\bm{\rho}}_{1})$ is the descent direction of $L({\bm{\rho}}_{1})$ with $-\nabla L({\bm{\rho}}_{1})^{\rm T}{\bm{\Theta}}\nabla L({\bm{\rho}}_{1})\leq 0$ since the projection operator ${\bm{\Theta}}$ is a symmetric positive definite matrix.
+Substituting (33) into (35), we have
+$\small\begin{cases}\varsigma_{i}\leq\varsigma_{i}^{\rm b},\ \varsigma_{i}^{\rm b}>0\\ \varsigma_{i}\geq\varsigma_{i}^{\rm b},\ \varsigma_{i}^{\rm b}<0\end{cases}$ (36)
+where $\varsigma_{i}^{\rm b}$ is the bound for $\varsigma_{i}$ with its expression given by
+$\small\varsigma_{i}^{\rm b}=\dfrac{\nabla L({\bm{\rho}}_{i+1})^{\rm T}{\bm{\Theta}}\nabla L({\bm{\rho}}_{i+1})}{{\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i+1})}.$ (37)
+The numerator satisfies $\nabla L({\bm{\rho}}_{i+1})^{\rm T}{\bm{\Theta}}\nabla L({\bm{\rho}}_{i+1})>0$ for any non-zero gradient $\nabla L({\bm{\rho}}_{i+1})$ due to the positive definite ${\bm{\Theta}}$.
+When employing inexact line searches, the deflection selection proposed by Hestenes and Stiefel [40], referred to as HS method, for $\varsigma_{i}$ is preferable.
+Accordingly, we adopt the modified HS deflection, given by
+$\small\varsigma_{i}^{\rm HS}=\dfrac{\nabla L({\bm{\rho}}_{i+1})^{\rm T}{\bm{q}}_{i}}{{\bm{d}}_{i}^{\rm T}{\bm{q}}_{i}},$ (38)
 which is obtained from ${\bm{d}}_{i+1}^{\rm T}{\bm{q}}_{i}=0$ with ${\bm{q}}_{i}\triangleq{\bm{\Theta}}[\nabla L({\bm{\rho}}_{i+1})-\nabla L({\bm{\rho}}_{i})]$.
-In fact, when using the MCG algorithm, the orthogonality is desired between the (conjugate) gradients in adjacent iterations, i.e.,
+In fact, when using the MCG algorithm, it is desirable to maintain orthogonality between the conjugate gradients of adjacent iterations, i.e.,
 $|{\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i+1})|\ll|{\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i})|,$ (39)
 $|\nabla L({\bm{\rho}}_{i+1})^{\rm T}{\bm{\Theta}}\nabla L({\bm{\rho}}_{i})|\ll|\nabla L({\bm{\rho}}_{i+1})^{\rm T}{\bm{\Theta}}\nabla L({\bm{\rho}}_{i+1})|.$ (40)
-Equations (43) and (44) ensure the positivity of both ${\bm{d}}_{i}^{\rm T}{\bm{q}}i$ and $\nabla L({\bm{\rho}}{i+1})^{\rm T}{\bm{q}}_{i}$, respectively.
-This allow for a positive conjugate gradient scaling factor $\varsigma_{i}^{\rm HS}$.
+(39) and (40) ensure the positivity of ${\bm{d}}_{i}^{\rm T}{\bm{q}}i$ and $\nabla L({\bm{\rho}}{i+1})^{\rm T}{\bm{q}}_{i}$, respectively.
+This allows for a positive conjugate gradient scaling factor $\varsigma_{i}^{\rm HS}$.
 Furthermore, (39) and (40) also imply that $|{\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i+1})|\ll|{\bm{d}}_{i}^{\rm T}{\bm{q}}_{i}|$ and $|\nabla L({\bm{\rho}}_{i+1})^{\rm T}{\bm{q}}_{i}|\approx|\nabla L({\bm{\rho}}_{i+1})^{\rm T}{\bm{\Theta}}\nabla L({\bm{\rho}}_{i+1})|$, respectively.
@@
 ${\varsigma_{i}=}\begin{cases}\varsigma_{i}^{\rm HS},\ \min\{\nabla L({\bm{\rho}}_{i+1})^{\rm T}{\bm{q}}_{i},{\bm{d}}_{i}^{\rm T}{\bm{q}}_{i},|\varsigma_{i}^{\rm b}|-\varsigma_{i}^{\rm HS}\}>0,\\ 0,\ {\rm else}.\end{cases}$ (41)
-Note that, after every consecutive $N$ MCG steps with $\varsigma_{i}=\varsigma_{i}^{\rm HS}$, we will also let $\varsigma_{i}=0$ to restart the new loop of MCG iteration.
-A proper reboot is potential in mitigating the accumulation of numerical errors, restoring the orthogonality and conjugacy of direction vectors, and avoiding local minima in nonlinear and non-convex optimization.
-To optimize the step size selection, we define $f_{i}(\upsilon_{i})\triangleq L({\bm{\rho}}_{i}+\upsilon_{i}{\bm{d}}_{i})$.
-Then, we have $f_{i}(0)=L({\bm{\rho}}_{i})$.
+Note that, after every consecutive $N$ MCG steps with $\varsigma_{i}=\varsigma_{i}^{\rm HS}$, we reset $\varsigma_{i}=0$ to initiate a new cycle of MCG iterations.
+Such a reboot strategy can help mitigate the accumulation of numerical errors, restoring the orthogonality and conjugacy of direction vectors, and avoiding local minima in nonlinear and non-convex optimization.
+IV-A2 Step size optimization
+To optimize the step size selection, we define $f_{i}(\upsilon_{i})\triangleq L({\bm{\rho}}_{i}+\upsilon_{i}{\bm{d}}_{i})$ with $f_{i}(0)=L({\bm{\rho}}_{i})$.
 The second-order Taylor approximation of $f_{i}(\upsilon_{i})$ at $\upsilon_{i}=0$ is given by
-$f_{i}(\upsilon_{i})$ $\thickapprox f_{i}(0)+\upsilon_{i}\nabla f_{i}(0)+\upsilon_{i}^{2}\nabla^{2}f_{i}(0)/2$
+$\small f_{i}(\upsilon_{i})$ $\thickapprox f_{i}(0)+\upsilon_{i}\nabla f_{i}(0)+\upsilon_{i}^{2}\nabla^{2}f_{i}(0)/2$
 $=f_{i}(0)+\upsilon_{i}{\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i})+\upsilon_{i}^{2}/2{\bm{d}}_{i}^{\rm T}\nabla^{2}L({\bm{\rho}}_{i}){\bm{d}}_{i},$ (42)
@@
 According to (35), we have ${\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i})\leq 0$.
-If $\|\nabla^{2}L({\bm{\rho}}_{i})\|$ is locally bounded, there exists a non-negative step size bound $\upsilon_{b}$ for given $\epsilon_{L}\in[0,1)$ such that, when $\upsilon_{i}\leq\upsilon_{b}$, we have
-$f_{i}(\upsilon_{i})-f_{i}(0)\leq\epsilon_{L}\upsilon_{i}{\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i}).$ (47)
-When ${\bm{d}}_{i}^{\rm T}\nabla^{2}L({\bm{\rho}}_{i}){\bm{d}}_{i}\leq 0$, (47) always holds.
+If $\|\nabla^{2}L({\bm{\rho}}_{i})\|$ is locally bounded, there exists a non-negative step size bound $\upsilon_{b}$ for given $\epsilon_{L}\in[0,1)$ such that
+$\small f_{i}(\upsilon_{i})-f_{i}(0)\leq\epsilon_{L}\upsilon_{i}{\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i}),\quad\text{for}\quad\upsilon_{i}\leq\upsilon_{b}.$ (43)
+(43) always holds when ${\bm{d}}_{i}^{\rm T}\nabla^{2}L({\bm{\rho}}_{i}){\bm{d}}_{i}\leq 0$.
 In contrast, we have $\upsilon_{b}=\dfrac{2(\epsilon_{L}-1){\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i})}{{\bm{d}}_{i}^{\rm T}\nabla^{2}L({\bm{\rho}}_{i}){\bm{d}}_{i}}$ when ${\bm{d}}_{i}^{\rm T}\nabla^{2}L({\bm{\rho}}_{i}){\bm{d}}_{i}>0$.
 Due to the inherent high computational complexity of the Hessian matrix $\nabla^{2}L({\bm{\rho}}_{i})$ in calaulating $\upsilon_{b}$, we can apply an inexact line search (ILS) for the step size to enable
-$L({\bm{\rho}}_{i}+\upsilon_{i}{\bm{d}}_{i})-L({\bm{\rho}}_{i})\leq\epsilon_{L}\upsilon_{i}{\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i}).$ (48)
+$\small L({\bm{\rho}}_{i}+\upsilon_{i}{\bm{d}}_{i})-L({\bm{\rho}}_{i})\leq\epsilon_{L}\upsilon_{i}{\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i}).$ (44)
 According to (35), the term on the right-hand side of (44) is always non-positive.
-Thus, (48) contributes to (38).
-The ILS controlling factor $\epsilon_{L}$ is used to control the decrement of the function $L({\bm{\rho}})$ by moving a step $\upsilon_{i}$.
-When (48) is unmet for given $\upsilon_{i}$, it is scaled by,
-$\upsilon_{i}=\sigma_{\upsilon}\upsilon_{i}$ (49)
-where the scaling factor $\sigma_{\upsilon}\in(0,1)$ and an initial step size is denoted as $\upsilon_{i}=\upsilon_{0,i}$.
-On the other hand, the penalty function does not inherently guarantee that the individual power budget constraints at each transmitter are satisfied during the optimization process.
-As a result, negative PA values may emerge during intermediate iterations.
-These infeasible PA values can lead to misleading CRLB evaluations, such as yielding negative CRLBs instead of infinite ones, thereby artificially satisfying the CRLB constraints.
-To prevent unnecessary iterations within the infeasible region and to ensure physically meaningful results, it is essential to enforce non-negative PAs at each iteration step.
-To this end, the step size $\upsilon_{i}$ needs to satisfy,
-$0<\upsilon_{i}\leq\upsilon_{i}^{\rm b}$ (50)
-where $\upsilon_{i}^{\rm b}$ is a step size upper bound given by
+Thus, (44) can enable a descent function in (34).
+The ILS controlling factor $\epsilon_{L}$ is used to regulate the decrement of the function $L({\bm{\rho}})$ by moving ${\bm{\rho}}_{i}$ towards the direction ${\bm{d}}_{i}$ weighted by the step size $\upsilon_{i}$.
+When (44) is unmet for given $\upsilon_{i}$, the step size is scaled by
+$\small\upsilon_{i}=\sigma_{\upsilon}\upsilon_{i},$ (45)
+where $\sigma_{\upsilon}\in(0,1)$ denotes the scaling factor and an initial step size is denoted as $\upsilon_{i}=\upsilon_{0,i}$.
+Additionally, the penalty function does not inherently ensure that the individual power budget constraints at each transmitter are respected during the optimization process.
+As a result, negative PA values may arise in intermediate iterations.
+These infeasible PA values can lead to misleading CRLB evaluations, for instance, producing negative CRLBs instead of infinite ones, thereby falsely satisfying the CRLB constraints.
+To avoid unnecessary iterations within the infeasible region and to ensure physically meaningful results, it is crucial to enforce non-negative PAs at each iteration step.
+To this end, the step size $\upsilon_{i}$ needs to satisfy
+$\small 0<\upsilon_{i}\leq\upsilon_{i}^{\rm b},$ (46)
+where $\upsilon_{i}^{\rm b}$ denotes an upper bound on the step size, given by
 $\upsilon_{i}^{\rm b}=\min_{d_{n,i}<0}\{\rho_{n,i}/|{d}_{n,i}|\},$ (47)
-with $\rho_{n,i}$ denoting the $n$th element of ${\bm{\rho}}_{i}$ and similar for ${d}_{n,i}$. Thus, the initial step size for ILS can be selected as,
-$\upsilon_{0,i}=\epsilon_{\upsilon}\upsilon_{i}^{\rm b}$ (52)
+with $\rho_{n,i}$ denoting the $n$th element of ${\bm{\rho}}_{i}$ and similar for ${d}_{n,i}$. Thus, the initial step size for ILS can be selected as
+$\small\upsilon_{0,i}=\epsilon_{\upsilon}\upsilon_{i}^{\rm b}$ (48)
 with $\epsilon_{\upsilon}\in(0,1]$.
-Note that there must exist ${d}_{n,i}<0$ at each iteration $i$ unless ${\bm{d}}_{i}={\bm{0}}$ due to $\sum_{n=1}^{N}{d}_{n,i}=0$.
-The ILS method characterized in (48)-(52) can be seen as the extension and application of the Armijo’s Rule [57, 56] in our work.
-A termination condition is significant for the iterative algorithm.
-A natural choice, based on the ILS condition in (48), is that the absolute difference in objective function values between consecutive iterations satisfies $|L({\bm{\rho}}_{i+1})-L({\bm{\rho}}_{i})|\leq\epsilon$ where $\epsilon$ is a small positive threshold.
-To avoid numerical instability arising from the magnitude of the objective function itself, we apply a normalized termination condition, i.e., $|L({\bm{\rho}}_{i+1})-L({\bm{\rho}}_{i})|/|L({\bm{\rho}}_{i+1})|\leq\epsilon_{\rm th}$, as employed in both the MCG-based and MSD-based algorithms.
-We now present the Penalty-function and Projection-based Modified Conjugate Gradient algorithm with inexact line search (PP-MCG-ILS).
-A large penalty factor may lead to computational issues, such as an ill-conditioned Hessian matrix, which can adversely affect the convergence of the MCG algorithm.
-To mitigate this, we adopt a strategy of progressively increasing the penalty factor across iterations.
-For given penalty factor $\mu_{k}$, the MCG algorithm is executed until its termination condition is satisfied.
-Then, the penalty factor is updated as $\mu_{k+1}=\varphi\mu_{k}$ for the next loop of penalty-based iteration, where $\varphi>1$ is a scaling constant.
-This process continues until the convergence criterion $\mu_{k}\alpha(\bm{\rho})\leq\epsilon_{\mu}$ is met, with $\epsilon_{\mu}$ being a small, user-defined threshold [54].
-This approach ensures numerical stability and gradual enforcement of the constraint.
-The complete PP-MCG-ILS algorithm is summarized in Algorithm 1.
+Note that there must exist at least one component ${d}_{n,i}<0$ in each iteration $i$, unless ${\bm{d}}_{i}={\bm{0}}$, due to the constraint $\sum_{n=1}^{N}{d}_{n,i}=0$.
+The ILS method characterized in (44)-(48) can be seen as the extension and application of the Armijo’s Rule [41] in our work.
+IV-A3 The algorithm implementation
+A well-defined termination condition is essential for the proposed iterative algorithm.
+A natural choice, inspired by the ILS condition in (44), is to monitor the absolute change in the objective function between consecutive iterations, i.e., $|L({\bm{\rho}}_{i+1})-L({\bm{\rho}}_{i})|\leq\epsilon$, where $\epsilon$ is a predefined threshold with small value.
+To avoid numerical instability arising from the magnitude of the objective function itself, we apply a normalized termination condition, i.e., $|L({\bm{\rho}}_{i+1})-L({\bm{\rho}}_{i})|/|L({\bm{\rho}}_{i+1})|\leq\epsilon_{\rm th}$.
+This condition is employed in both the MCG-based and MSD-based algorithms.
+We now present the complete PP-MCG-ILS.
+Generally, a large penalty factor may lead to numerical issues, such as an ill-conditioned Hessian matrix, which can degrade the convergence behavior of the MCG algorithm.
+To mitigate this, we adopt a sequential penalty strategy that gradually increases the penalty factor across successive iteration loops.
+Specifically, for a given penalty factor $\mu_{k}$, the MCG algorithm is executed until its termination condition is satisfied.
+Then, the penalty factor is updated as $\mu_{k+1}=\varphi\mu_{k}$ where $\varphi>1$ is a scaling constant, for the next loop of penalty-based iteration.
+This iterative refinement continues until the convergence criterion $\mu_{k}\alpha(\bm{\rho})\leq\epsilon_{\mu}$ is satisfied, where $\epsilon_{\mu}$ is a small, predefined threshold [39].
+This sequential penalty strategy ensures both numerical stability and progressive enforcement of the constraint conditions.
+The detailed procedure of the proposed PP-MCG-ILS algorithm is summarized in Algorithm 1.
 Algorithm 1 The PP-MCG-ILS algorithm
@@
 Output: The PA vector ${\bm{y}}_{j}$.
-1:  Step size: Calculate $\upsilon_{i}^{\rm b}$ according to (51); set $\upsilon_{i}=\epsilon_{\upsilon}\upsilon_{i}^{\rm b}$ , while $L({\bm{\rho}}_{i}+\upsilon_{i}{\bm{d}}_{i})-L({\bm{\rho}}_{i})>\epsilon_{L}\upsilon_{i}{\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i})$, let $\upsilon_{i}=\sigma_{\upsilon}\upsilon_{i}$.
-2:  Parameter update: ${\bm{\rho}}_{i+1}={\bm{\rho}}_{i}+\upsilon_{i}{\bm{d}}_{i}$.
-3:  Convergence condition: If $|L({\bm{\rho}}_{i+1})-L({\bm{\rho}}_{i})|/|L({\bm{\rho}}_{i+1})|\leq\epsilon_{\rm th}$, ${\bm{y}}_{j}={\bm{\rho}}_{i+1}$, and go to step 5; Otherwise, go to next step.
-4:  Direction update: According to (45), if $\varsigma_{i}=\varsigma_{i}^{\rm HS}$, $k=k+1$, or $k=0$ otherwise; when $k=N$, let $\varsigma_{i}=0$ and $k=0$; update ${\bm{d}}_{i+1}=-{\bm{\Theta}}\nabla L({\bm{\rho}}_{i+1})+\varsigma_{i}{\bm{d}}_{i}$; let $i=i+1$, and go back to step 1.
-5:  Penalty update: If $\mu_{j}\alpha({\bm{y}}_{j})<\epsilon_{\mu}$, stop; Otherwise, let $\mu_{j+1}=\varphi\mu_{j}$, and ${\bm{d}}_{i+1}=-{\bm{\Theta}}\nabla L({\bm{\rho}}_{i+1})$; let $j=j+1$ and $i=i+1$, and go to step 1.
-V-B Modified steepest descent algorithm
-In fact, alternative unconstrained optimization methods can also be employed to solve the proposed projection-based penalty problem in (31).
-For comparison purposes, we also develop a Modified Steepest Descent (MSD) algorithm, adapted from the proposed MCG framework.
-The primary distinction between the MSD and MCG algorithms lies in the choice of the descent direction.
-Specifically, in the MSD method, the descent direction at each iteration is directly set as the negative gradient of the objective function
-${\bm{d}}_{i}$ $=-{\bm{\Theta}}\nabla L({\bm{\rho}}_{i}),$ (53)
-which can be directly obtained from MCG by setting the deflection factor to zero.
-We herein omit the detailed implementation steps of PP-MSD-ILS since they can be obtained from the PP-MCG-ILS by simply revising step 4 of Algorithm 1.
-V-C Benchmarks
-We now introduce two benchmark algorithms, derived from the PP-MCG-ILS and PP-MSD-ILS frameworks, respectively.
-These benchmark variants adopt fixed step sizes at each iteration instead of employing ILS.
-To mitigate potential instability arising from overly large or small gradient magnitudes under fixed step updates, we normalize the update directions using their $l_{2}$-norm.
-For the Penalty-function and Projection driven Normalized Steepest Descent (PP-NSD) algorithm, the descent direction is defined as
+1:  Step size bound: Calculate $\upsilon_{i}^{\rm b}$ according to (47);
+2:  Step size search: set $\upsilon_{i}=\epsilon_{\upsilon}\upsilon_{i}^{\rm b}$, while $L({\bm{\rho}}_{i}+\upsilon_{i}{\bm{d}}_{i})-L({\bm{\rho}}_{i})>\epsilon_{L}\upsilon_{i}{\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i})$, let $\upsilon_{i}=\sigma_{\upsilon}\upsilon_{i}$.
+3:  PA update: ${\bm{\rho}}_{i+1}={\bm{\rho}}_{i}+\upsilon_{i}{\bm{d}}_{i}$.
+4:  Convergence condition: If $|L({\bm{\rho}}_{i+1})-L({\bm{\rho}}_{i})|/|L({\bm{\rho}}_{i+1})|\leq\epsilon_{\rm th}$, ${\bm{y}}_{j}={\bm{\rho}}_{i+1}$, and go to step 7; Otherwise, go to next step.
+5:  Deflection choice: According to (41), if $\varsigma_{i}=\varsigma_{i}^{\rm HS}$, $k=k+1$, or $k=0$ otherwise; when $k=N$, let $\varsigma_{i}=0$ and $k=0$;
+6:  Direction update: ${\bm{d}}_{i+1}=-{\bm{\Theta}}\nabla L({\bm{\rho}}_{i+1})+\varsigma_{i}{\bm{d}}_{i}$; let $i=i+1$, and go back to step 1.
+7:  Penalty update: If $\mu_{j}\alpha({\bm{y}}_{j})<\epsilon_{\mu}$, stop; Otherwise, $\mu_{j+1}=\varphi\mu_{j}$, and ${\bm{d}}_{i+1}=-{\bm{\Theta}}\nabla L({\bm{\rho}}_{i+1})$; let $j=j+1$ and $i=i+1$, and go to step 1.
+IV-B Modified Steepest Descent Algorithm
+In fact, alternative unconstrained optimization methods can be applied to solve the proposed projection-based penalty problem in (26).
+In this paper, we also develop a MSD algorithm, adapted from the proposed MCG framework.
+The main difference between the MSD and MCG algorithms lies in how the search direction is determined.
+In the MSD approach, the search direction at each iteration is set to the negative gradient of the objective function, i.e.,
+$\small{\bm{d}}_{i}$ $=-{\bm{\Theta}}\nabla L({\bm{\rho}}_{i}),$ (49)
+which can be obtained from MCG by setting the deflection factor to zero.
+We herein omit the detailed implementation steps of PP-MSD-ILS since they can be obtained from the PP-MCG-ILS in Algorithm 1 by simply removing step 5 and performing the direction update in step 6 using (49).
+IV-C Benchmarking Schemes
+To facilitate performance comparison, we introduce two benchmark algorithms based on the PP-MCG-ILS and PP-MSD-ILS frameworks, called penalty function and projection-driven normalized SD (PP-NSD) and penalty function and projection-driven normalized CG (PP-NCG), respectively.
+Here, we employ the standard SD and CG methods proposed in [42], where fixed step sizes are adopted instead of ILS.
+To mitigate potential instability arising from overly large or small gradient magnitudes under fixed-step updates, we normalize the update directions using their $l_{2}$ norm.
+Specifically, for the PP-NSD algorithm, the search direction is defined as an $\epsilon_{d}$-regularized normalized gradient, given by
 ${\bm{d}}_{i}$ $=-{\bm{g}}_{i}/(\|{\bm{g}}_{i}\|_{2}+\epsilon_{d}),$ (50)
-where ${\bm{g}}_{i}={\bm{\Theta}}\nabla L({\bm{\rho}}_{i})$, and $\epsilon_{d}$ is a small constant to prevent division by zero.
-Similarly, for the penalty function and projection driven normalized CG (PP-NCG) algorithm, the descent direction is given by,
-${\bm{d}}_{i+1}$ $=-({\bm{g}}_{i+1}+\varsigma_{i}{\bm{d}}_{i})/(\|({\bm{g}}_{i+1}+\varsigma_{i}{\bm{d}}_{i})\|_{2}+\epsilon_{d}),$ (55)
+where ${\bm{g}}_{i}={\bm{\Theta}}\nabla L({\bm{\rho}}_{i})$, and $\epsilon_{d}$ is a small positive constant introduced to prevent division by zero.
+Similarly, for the PP-NCG algorithm, the search direction is given by an $\epsilon_{d}$-regularized normalized conjugate gradient, i.e.,
+$\small{\bm{d}}_{i+1}$ $=-({\bm{g}}_{i+1}+\varsigma_{i}{\bm{d}}_{i})/(\|({\bm{g}}_{i+1}+\varsigma_{i}{\bm{d}}_{i})\|_{2}+\epsilon_{d}),$ (51)
 where the deflection factor $\varsigma_{i}$ is computed by (41).
-The step size selection in both benchmark algorithms is
+The step size in both benchmark algorithms is given by
 $\upsilon_{i}=\min\{\epsilon_{\upsilon}\upsilon_{i}^{\rm b},\upsilon_{0}\},$ (52)
-with $\upsilon_{i}^{\rm b}$ provided by (51) and $\upsilon_{0}$ being a predetermined step size.
-Additionally, the descent function (38) may not hold when utilizing the fixed step sizes.
-Therefore, it is necessary to provide a sufficiently large integer $I$ as the upper bound of the number of iterations.
-In the same vain, due to the non-monotonic penalty function, the minimum may not occur at the final iteration but within a few steps before it.
-To address this, we introduce an alternative penalty stopping criterion given by $\mu_{j}\min\limits_{i+1-S\leq q\leq i+1}\{\alpha({\bm{\rho}}_{q})\}<\epsilon_{\mu}$ where $S$ represents the window size for searching the minimum penalty value over recent iterations.
-The detailed implementation steps of the PP-NSD algorithm are omitted here, as it can be readily constructed by modifying step 4 of the PP-NCG algorithm described in Algorithm 2.
+with $\upsilon_{i}^{\rm b}$ given in (47) and $\upsilon_{0}$ being a predetermined step size.
+The maximum number of iterations is set to $I$.
+Due to the non-monotonic penalty function, the minimum may not occur at the final iteration but within a few steps before it.
+To address this, we introduce an alternative penalty stopping criterion given by $\mu_{j}\min\limits_{i+1-S\leq q\leq i+1}\{\alpha({\bm{\rho}}_{q})\}<\epsilon_{\mu}$ where $S$ represents the window size for searching the minimum penalty value.
+The detailed implementation steps of the PP-NSD algorithm are omitted here, as it can be easily constructed from the PP-NCG described in Algorithm 2 by removing step 4 and performing the direction update in step 5 using (50).
 Algorithm 2 The PP-NCG algorithm
@@
 1:  Step size: $\upsilon_{i}=\min\{\epsilon_{\upsilon}\upsilon_{i}^{\rm b},\upsilon_{0}\}$.
-2:  Parameter update: ${\bm{\rho}}_{i+1}={\bm{\rho}}_{i}+\upsilon_{i}{\bm{d}}_{i}$.
-3:  Convergence condition: If $|L({\bm{\rho}}_{i+1})-L({\bm{\rho}}_{i})|/|L({\bm{\rho}}_{i+1})|\leq\epsilon_{\rm th}$ or $i+1=I$, ${\bm{y}}_{j}={\bm{\rho}}_{i+1}$, go to step 5; Otherwise, go to next step.
-4:  Direction update: According to (45), if $\varsigma_{i}=\varsigma_{i}^{\rm HS}$, $k=k+1$, or $k=0$ otherwise; when $k=N$, let $\varsigma_{i}=0$ and $k=0$; update ${\bm{d}}_{i+1}$ through (55); let $i=i+1$, and go back to step 1.
-5:  Penalty update: If $\mu_{j}\min\limits_{i+1-S\leq q\leq i+1}\{\alpha({\bm{\rho}}_{q})\}<\epsilon_{\mu}$, stop; Otherwise, let $\mu_{j+1}=\varphi\mu_{j}$, and ${\bm{d}}_{i+1}=-{\bm{\Theta}}\nabla L({\bm{\rho}}_{i+1})/(\|{\bm{\Theta}}\nabla L({\bm{\rho}}_{i+1})\|_{2}+\epsilon_{d})$; let $j=j+1$ and $i=i+1$, and go to step 1.
-VI Extension to pure sensing
-In this section, we propose a Penalty-function based Normalized Conjugate Gradient algorithm with inexact line search (P-NCG-ILS) for the PA in pure sensing scenario.
-We consider the power minimization under power budget constraints and CRLB constraints, i.e.,
-$\arg\min\limits_{{\bm{\rho}}}$ $\ {\bm{\rho}}^{\rm T}{\bm{1}},$ (57a)
+2:  PA update: ${\bm{\rho}}_{i+1}={\bm{\rho}}_{i}+\upsilon_{i}{\bm{d}}_{i}$.
+3:  Convergence condition: If $|L({\bm{\rho}}_{i+1})-L({\bm{\rho}}_{i})|/|L({\bm{\rho}}_{i+1})|\leq\epsilon_{\rm th}$ or $i+1=I$, ${\bm{y}}_{j}={\bm{\rho}}_{i+1}$, go to step 6; Otherwise, go to next step.
+4:  Deflection choice: According to (41), if $\varsigma_{i}=\varsigma_{i}^{\rm HS}$, $k=k+1$, or $k=0$ otherwise; when $k=N$, let $\varsigma_{i}=0$ and $k=0$;
+5:  Direction update: ${\bm{d}}_{i+1}$ through (51); let $i=i+1$, and go back to step 1.
+6:  Penalty update: If $\mu_{j}\min\limits_{i+1-S\leq q\leq i+1}\{\alpha({\bm{\rho}}_{q})\}<\epsilon_{\mu}$, stop; Otherwise, let $\mu_{j+1}=\varphi\mu_{j}$, and ${\bm{d}}_{i+1}=-{\bm{\Theta}}\nabla L({\bm{\rho}}_{i+1})/(\|{\bm{\Theta}}\nabla L({\bm{\rho}}_{i+1})\|_{2}+\epsilon_{d})$; let $j=j+1$ and $i=i+1$, and go to step 1.
+V Extension to Pure Sensing
+In this section, we consider a pure sensing scenario with joint location and velocity estimation.
+The P-NCG-ILS algorithm is proposed for PA in minimizing the total power under power budget constraints and CRLB constraints.
+The constrained power minimization problem is formulated as
+$\small\arg\min\limits_{{\bm{\rho}}}$ $\ {\bm{\rho}}^{\rm T}{\bm{1}},$ (53a)
 ${\rm s.t.}$ ${\bm{\rho}}^{\rm T}{\bm{1}}\leq 1,$ (53b)
 […]
-We present the penalty function as follows
-$\alpha_{\rm s}({\bm{\rho}})=$ ${f}_{\rm s}({\bm{\rho}})+\tilde{f}_{l}({\bm{\rho}})+\tilde{f}_{v}({\bm{\rho}})$
+Analogously to the derivations of PP-MCG-ILS, we set the penalty function for violating the inequality constraints as
+$\small\alpha_{\rm s}({\bm{\rho}})=$ ${f}_{\rm s}({\bm{\rho}})+\tilde{f}_{l}({\bm{\rho}})+\tilde{f}_{v}({\bm{\rho}})$
 $+\sum_{n=1}^{N}\left(f_{\min,n}({\bm{\rho}})+f_{\max,n}({\bm{\rho}})\right),$ (54)
-with the penalty function for the total power budget
-${f}_{\rm s}({\bm{\rho}})=(\max\{0,{\bm{\rho}}^{\rm T}{\bm{1}}-1\})^{q}.$ (59)
-Then, the objective function can be formulated as
-$L_{\rm s}({\bm{\rho}})={\bm{\rho}}^{\rm T}{\bm{1}}+\mu\alpha_{\rm s}({\bm{\rho}}).$ (60)
+where ${f}_{\rm s}({\bm{\rho}})$ denotes the penalty sub-function for the total power budget, i.e.,
+$\small{f}_{\rm s}({\bm{\rho}})=(\max\{0,{\bm{\rho}}^{\rm T}{\bm{1}}-1\})^{q}.$ (55)
+Then, the unconstrained optimization problem can be formulated as
+$\small\arg\min\limits_{{\bm{\rho}}}\ L_{\rm s}({\bm{\rho}})={\bm{\rho}}^{\rm T}{\bm{1}}+\mu\alpha_{\rm s}({\bm{\rho}}).$ (56)
+Similarly to the PP-MCG-ILS algorithm, we also use the CG-based method to solve (56).
+To avoid a potential numerical instability, we consider normalized CG search direction, given by ${\bm{d}}_{i+1}=\tilde{\bm{d}}_{i+1}/(\|\tilde{\bm{d}}_{i+1}\|_{2}+\epsilon_{d})$ with $\tilde{\bm{d}}_{i+1}=-\nabla L_{\rm s}({\bm{\rho}}_{i+1})+\varsigma_{i}{\bm{d}}_{i}$, where the deflection computation method is the same as that of the PP-MCG-ILS algorithm.
+Additionally, ILS is used to determine the step size at each iteration.
 The complete P-NCG-ILS algorithm is summarized in Algorithm 3.
@@
 Output: The PA vector ${\bm{y}}_{j}$.
-1:  Step size: $\upsilon_{i}=\min\{\epsilon_{\upsilon}\upsilon_{i}^{\rm b},\upsilon_{0}\}$; while $L_{\rm s}({\bm{\rho}}_{i}+\upsilon_{i}{\bm{d}}_{i})-L_{\rm s}({\bm{\rho}}_{i})>\epsilon_{L}\upsilon_{i}{\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i})$, let $\upsilon_{i}=\sigma_{\upsilon}\upsilon_{i}$.
-2:  Parameter update: ${\bm{\rho}}_{i+1}={\bm{\rho}}_{i}+\upsilon_{i}{\bm{d}}_{i}$.
-3:  Convergence condition: If $|L_{\rm s}({\bm{\rho}}_{i+1})-L_{\rm s}({\bm{\rho}}_{i})|/|L_{\rm s}({\bm{\rho}}_{i+1})|\leq\epsilon_{\rm th}$, ${\bm{y}}_{j}={\bm{\rho}}_{i+1}$, and go to step 5; Otherwise, go to next step.
-4:  Direction update: According to (45), if $\varsigma_{i}=\varsigma_{i}^{\rm HS}$, $k=k+1$, or $k=0$ otherwise; when $k=N$, let $\varsigma_{i}=0$ and $k=0$; update ${\bm{d}}_{i+1}=\tilde{\bm{d}}_{i+1}/(\|\tilde{\bm{d}}_{i+1}\|_{2}+\epsilon_{d})$ with $\tilde{\bm{d}}_{i+1}=-\nabla L_{\rm s}({\bm{\rho}}_{i+1})+\varsigma_{i}{\bm{d}}_{i}$; let $i=i+1$, and go back to step 1.
-5:  Penalty update: If $\mu_{j}\alpha({\bm{y}}_{j})<\epsilon_{\mu}$, stop; Otherwise, let $\mu_{j+1}=\varphi\mu_{j}$, and ${\bm{d}}_{i+1}=-\nabla L_{\rm s}({\bm{\rho}}_{i+1})$; let $j=j+1$ and $i=i+1$, and go to step 1.
-VII Algorithm Analysis
-The proposed algorithms (e.g., PP-MCG-ILS, PP-MSD-ILS and P-NCG-ILS) can be extended and applied to a broad class of nonlinear and non-convex optimization problems.
+1:  Step size bound: Calculate $\upsilon_{i}^{\rm b}$ according to (47);
+2:  Step size: $\upsilon_{i}=\min\{\epsilon_{\upsilon}\upsilon_{i}^{\rm b},\upsilon_{0}\}$; while $L_{\rm s}({\bm{\rho}}_{i}+\upsilon_{i}{\bm{d}}_{i})-L_{\rm s}({\bm{\rho}}_{i})>\epsilon_{L}\upsilon_{i}{\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i})$, let $\upsilon_{i}=\sigma_{\upsilon}\upsilon_{i}$.
+3:  PA update: ${\bm{\rho}}_{i+1}={\bm{\rho}}_{i}+\upsilon_{i}{\bm{d}}_{i}$.
+4:  Convergence condition: If $|L_{\rm s}({\bm{\rho}}_{i+1})-L_{\rm s}({\bm{\rho}}_{i})|/|L_{\rm s}({\bm{\rho}}_{i+1})|\leq\epsilon_{\rm th}$, ${\bm{y}}_{j}={\bm{\rho}}_{i+1}$, and go to step 5; Otherwise, go to next step.
+5:  Deflection choice: According to (41), if $\varsigma_{i}=\varsigma_{i}^{\rm HS}$, $k=k+1$, or $k=0$ otherwise; when $k=N$, let $\varsigma_{i}=0$ and $k=0$.
+6:  Direction update: ${\bm{d}}_{i+1}=\tilde{\bm{d}}_{i+1}/(\|\tilde{\bm{d}}_{i+1}\|_{2}+\epsilon_{d})$ with $\tilde{\bm{d}}_{i+1}=-\nabla L_{\rm s}({\bm{\rho}}_{i+1})+\varsigma_{i}{\bm{d}}_{i}$; let $i=i+1$, and go back to step 1.
+7:  Penalty update: If $\mu_{j}\alpha({\bm{y}}_{j})<\epsilon_{\mu}$, stop; Otherwise, $\mu_{j+1}=\varphi\mu_{j}$, and ${\bm{d}}_{i+1}=-\nabla L_{\rm s}({\bm{\rho}}_{i+1})$; let $j=j+1$ and $i=i+1$, and go to step 1.
+VI Algorithm Analysis
+The proposed algorithms (e.g., PP-MCG-ILS, PP-MSD-ILS) can be extended and applied to a broad class of nonlinear and non-convex optimization problems.
 These problems are typically characterized by an affine equality constraint, if present, satisfied through a projection operator, along with nonlinear or non-convex constraints incorporated via penalty functions.
 The ILS mechanism is used to enhance convergence.
+For example, the proposed P-NCG-ILS algorithm for sensing-only scenario can be seen as a simple extension of the PP-MCG-ILS algorithm.
 In the following, we analyze the convergence behavior and computational complexity of the proposed algorithms.
-VII-A Convergence analysis
-We now provide the convergence condition of the proposed PP-MCG-ILS and PP-MSD-ILS algorithms.
+VI-A Convergence Analysis
+We now provide the convergence analysis of the proposed PP-MCG-ILS and PP-MSD-ILS algorithms.
 Before delving to the details, we define the increment of the parameter update and the objective function respectively as
-$\Delta{\bm{\rho}}_{i+1}$ $\triangleq{\bm{\rho}}_{i+1}-{\bm{\rho}}_{i}=v_{i}{\bm{d}}_{i}$ (61)
+$\small\Delta{\bm{\rho}}_{i+1}$ $\triangleq{\bm{\rho}}_{i+1}-{\bm{\rho}}_{i}=v_{i}{\bm{d}}_{i},$ (57)
 and
@@
 The sufficient and necessary condition for $\|\Delta{\bm{\rho}}_{i+1}\|\rightarrow 0$ and $\Delta L({\bm{\rho}}_{i+1})\rightarrow 0$ is that, there exists an integer $I$ such that, ${\bm{d}}_{i}\rightarrow{\bm{0}}$ or $v_{i}\rightarrow 0$ when $i>I$.
-Proof. The proof is given in Appendix A. $\blacksquare$
+Proof. The proof is given in Appendix B. $\blacksquare$
 Definition.
-A function $L({\bm{\rho}})$ is Lipschitz gradient continuous over a set $\Omega$, if there exists a constant $C>0$, $\|\nabla L({\bm{\rho}}_{i})-\nabla L({\bm{\rho}}_{j})\|\leq C\|{\bm{\rho}}_{i}-{\bm{\rho}}_{j}\|$ for any ${\bm{\rho}}_{i},{\bm{\rho}}_{j}\in\Omega$ [58, 59].
+A function $L({\bm{\rho}})$ is Lipschitz gradient continuous over a set $\Omega$, if there exists a constant $C>0$, $\|\nabla L({\bm{\rho}}_{i})-\nabla L({\bm{\rho}}_{j})\|\leq C\|{\bm{\rho}}_{i}-{\bm{\rho}}_{j}\|$ for any ${\bm{\rho}}_{i},{\bm{\rho}}_{j}\in\Omega$ [43].
 Theorem 1.
-Given a nonempty convex set $\Omega$ where the objective function $L({\bm{\rho}}_{i})$ is bounded over its interior points, if the objective function is Lipschitz gradient continuous over $\Omega$, the ILS-based descent direction iterative algorithm will converge to a stationary point with a zero descent direction ${\bm{d}}_{i}\rightarrow{\bm{0}}$.
-In contrast, if it is not Lipschitz gradient continuous, the sequence $\{{\bm{\rho}}_{i}\}$ and $\{L({\bm{\rho}}_{i})\}$ may converge to a Lipschitz gradient noncontinuous point where the norm of the Hessian matrix may approach infinity.
-Proof. The proof is given in Appendix B. $\blacksquare$
-Theorem 1 provides the convergence analysis for the proposed PP-MCG-ILS and PP-MSD-ILS algorithms.
+Given a nonempty convex set $\Omega$ with the objective function $L({\bm{\rho}}_{i})$ bounded over its interior points, if the objective function is Lipschitz gradient continuous over $\Omega$, the ILS-based iterative algorithm will converge to a stationary point with a zero descent direction ${\bm{d}}_{i}\rightarrow{\bm{0}}$.
+In contrast, if it is Lipschitz gradient non-continuous, sequences $\{{\bm{\rho}}_{i}\}$ and $\{L({\bm{\rho}}_{i})\}$ may converge to a Lipschitz gradient noncontinuous point with the norm of the Hessian matrix approaching infinity.
+Proof. The proof is given in Appendix C. $\blacksquare$
+Theorem 1 provides the convergence analysis for the proposed iterative algorithms.
 We now prove the Lipschitz gradient continuity of the objectivity functions in (26) and (56).
-The feasible domain $\Omega\subset\mathbb{R}^{N}$ is closed, bounded, and convex.
-All CRLB-involved matrices, e.g., ${\bm{P}}(\rho)$, ${\bm{Y}}(\rho)$, and their inverses are uniformly bounded and positive definite over the set $\Omega$.
-Firstly, each term $f_{\min,n}({\bm{\rho}})$ and $f_{\max,n}({\bm{\rho}})$ is piecewise quadratic, and its gradient is Lipschitz continuous due to continuity and boundedness.
-So does ${f}_{\rm s}({\bm{\rho}})$.
-Then, since ${\bm{C}}_{L}({\bm{\rho}})$ is uniformly positive definite and smooth in ${\bm{\rho}}$, it follows that: $f_{l}({\bm{\rho}})$ is continuously differentiable, and so is
+Firstly, all CRLB-involved matrices, e.g., ${\bm{P}}(\rho)$, ${\bm{Y}}(\rho)$, and their inverses are smooth on ${\bm{\rho}}$ over the feasible domain $\Omega\subset\mathbb{R}^{N}$.
+Moreover, ${\bm{C}}_{L}({\bm{\rho}})$ is uniformly positive definite.
+It follows that ${\bm{C}}_{L}({\bm{\rho}})$ is continuously differentiable, and so is ${\rm tr}\{{\bm{C}}_{L}({\bm{\rho}})\}$.
+By the chain rule, we have
 $\nabla\tilde{f}_{l}({\bm{\rho}})=\begin{cases}2f_{l}({\bm{\rho}})\nabla f_{l}({\bm{\rho}}),&\text{if }f_{l}({\bm{\rho}})>0\\ 0,&\text{otherwise}.\end{cases}$ (59)
-Given that $\nabla f_{l}({\bm{\rho}})$ is smooth and bounded (due to the matrix inverse being smooth on positive definite matrices), $\nabla\tilde{f}_{l}({\bm{\rho}})$ is Lipschitz continuous.
-A similar argument applies to $\tilde{f}_{v}({\bm{\rho}})$.
-Hence, $\nabla_{\bm{\rho}}\alpha({\bm{\rho}})$ and $\nabla_{\bm{\rho}}\alpha_{\rm s}({\bm{\rho}})$ are Lipschitz continuous over $\Omega$, and so are $\nabla_{\bm{\rho}}L({\bm{\rho}})$ and $\nabla_{\bm{\rho}}L_{\rm s}({\bm{\rho}})$.
+where $\nabla f_{l}({\bm{\rho}})=-{\rm tr}\{{\bm{C}}_{L}({\bm{\rho}})\nabla({\bm{P}}-{\bm{V}}{\bm{Y}}^{-1}{\bm{V}}^{\rm T}){\bm{C}}_{L}({\bm{\rho}})\}$.
+Since the matrix inverse and trace are smooth on the set of positive definite matrices, $\nabla f_{l}({\bm{\rho}})$ is continuously differentiable.
+Moreover, all matrices involved (including ${\bm{C}}_{L}({\bm{\rho}})$) have bounded spectral norm, so the operator norm of the Hessian of $f_{l}({\bm{\rho}})$ is also bounded.
+Consequently, $\nabla\tilde{f}_{l}({\bm{\rho}})$ is Lipschitz continuous.
+A similar argument applies to $\nabla\tilde{f}_{v}({\bm{\rho}})$.
+On the other hand, $f_{\min,n}({\bm{\rho}})$, $f_{\max,n}({\bm{\rho}})$ and ${f}_{\rm s}({\bm{\rho}})$ are piecewise quadratic, and their gradients are Lipschitz continuous.
+Hence, $\nabla_{\bm{\rho}}\alpha({\bm{\rho}})$ and $\nabla_{\bm{\rho}}\alpha_{\rm s}({\bm{\rho}})$ are Lipschitz continuous over $\Omega$, and the same applies to $\nabla_{\bm{\rho}}L({\bm{\rho}})$ and $\nabla_{\bm{\rho}}L_{\rm s}({\bm{\rho}})$.
 Additionally, since ${\bm{\Theta}}$ is a constant matrix with $\|{\bm{\Theta}}\|_{2}=1$, and $\nabla_{\bm{\rho}}f({\bm{\rho}})$ is Lipschitz continuous, the composition $\nabla\mathcal{L}({\bm{\varpi}})={\bm{\Theta}}\nabla L({\bm{\rho}})$ is also Lipschitz continuous.
-There exists a constant $C>0$ such that $\forall{\bm{\varpi}}_{1},{\bm{\varpi}}_{2}\in\Omega_{\bm{\varpi}}$
-$\|\nabla_{{\bm{\varpi}}}\mathcal{L}({\bm{\varpi}}_{1})-\nabla_{{\bm{\varpi}}}\mathcal{L}({\bm{\varpi}}_{2})\|_{2}\leq C\|{\bm{\varpi}}_{1}-{\bm{\varpi}}_{2}\|_{2}.$
+In other words, there exists a constant $C>0$ such that
+$\small\|\nabla_{{\bm{\varpi}}}\mathcal{L}({\bm{\varpi}}_{1})-\nabla_{{\bm{\varpi}}}\mathcal{L}({\bm{\varpi}}_{2})\|_{2}\leq C\|{\bm{\varpi}}_{1}-{\bm{\varpi}}_{2}\|_{2},\forall{\bm{\varpi}}_{1},{\bm{\varpi}}_{2}\in\Omega_{\bm{\varpi}}$
 Therefore, $\mathcal{L}({\bm{\varpi}})$ is Lipschitz gradient continuous on $\Omega_{\bm{\varpi}}$. $\blacksquare$
-VII-B Complexity analysis
-We observe that the additional computational complexity of the PP-MCG-ILS algorithm, compared to the PP-MSD-ILS algorithm, primarily stems from the computation of $\varsigma_{i}^{\rm b}$ and $\varsigma_{i}^{\rm HS}$ at each iteration, which requires approximately $4N$ multiplications and 2 divisions.
+VI-B Complexity Analysis
+We now qualitatively analyze and compare the computational complexity of the proposed algorithms and their benchamrks.
+We observe that the increased computations of the PP-MCG-ILS algorithm, compared to the PP-MSD-ILS algorithm, primarily stems from the computation of $\varsigma_{i}^{\rm b}$ and $\varsigma_{i}^{\rm HS}$ at each iteration, which involves approximately $4N$ multiplications and 2 divisions.
 In both algorithms, the dominant computational burden lies in evaluating the objective function during the ILS and computing the gradients for direction updates, which are inherently tied to the CRLB expressions.
-Therefore, a more meaningful comparison of total computational cost focuses on the number of ILS steps per iteration.
-This number is influenced by the step size threshold, given by the right-hand side of (67), the initial step size $\upsilon_{0,i}$ in (52), and the scaling factor.
-At a given iterate ${\bm{\rho}}_{i}$, a larger descent direction norm $\|{\bm{d}}_{i}\|$ generally leads to a smaller $\upsilon_{0,i}$ and a reduced step size threshold.
-As a result, the number of ILS steps is expected to be comparable between the PP-MCG-ILS and PP-MSD-ILS algorithms.
-Given this, the primary metrics for evaluating the relative performance of these two algorithms become their convergence rate and the quality of their steady-state solutions.
+Therefore, a more meaningful comparison of the overall computational cost focuses on the number of ILS steps performed per iteration.
+This number is influenced by the step size threshold, given by the right-hand side of (C), the initial step size $\upsilon_{0,i}$ in (48), and the scaling factor.
+At a given iterate ${\bm{\rho}}_{i}$, a larger search direction norm $\|{\bm{d}}_{i}\|$ generally leads to a smaller $\upsilon_{0,i}$ and a reduced step size threshold.
+As a consequence, the number of ILS steps per iteration is expected to remain comparable between the PP-MCG-ILS and PP-MSD-ILS algorithms.
+Accordingly, the primary metrics for evaluating the relative performance of these two algorithms are their convergence rate and the quality of their steady-state solutions.
 We now qualitatively compare the computational complexity between algorithms employing ILS and those without ILS.
 As previously discussed, the number of ILS steps significantly impacts the number of evaluations of the objective function.
-Thus, the computational complexity of the benchmark algorithms (without ILS) can be considerably lower at each iteration during step 1 compared to the PP-MCG-ILS and PP-MSD-ILS algorithms.
-However, all four algorithms have similar complexity in step 4 due to the inherently intensive gradient calculations involved.
-Moreover, the overall computational complexity heavily depends on the convergence rates, or equivalently, the total number of iterations required for convergence.
+Thus, the computational complexity for step size selection of the benchmark algorithms can be considerably lower at each iteration compared to the PP-MCG-ILS and PP-MSD-ILS algorithms.
+However, all four algorithms have similar complexity in direction update due to the inherently intensive gradient calculations involved.
+Moreover, the overall computational complexity heavily depends on the convergence rates, or equivalently, on the total number of iterations required for convergence.
 As will be demonstrated in the subsequent simulation results, the proposed PP-MCG-ILS algorithm exhibits a notably higher convergence rate compared to its counterparts, with or without ILS.
-This superior convergence performance significantly reduces the overall computational effort, particularly benefiting the demanding computations of gradients.
-VIII Simulation Results
-In this section, we conduct some simulations to evaluate the proposed PP-MCG-ILS and PP-MSD-ILS algorithms for PA optimization.
-All APs, including $N$ single antenna transmitters and $K$ single antenna receivers, are assumed to be stationary.
+This superior convergence performance significantly reduces the overall computational effort, particularly in the sense that it benefits demanding gradient computations.
+VII Simulation Results
+In this section, we conduct numerical simulations to evaluate the proposed PP-MCG-ILS, PP-MSD-ILS and P-NCG-ILS algorithms for PA optimization.
+All APs are assumed to be stationary.
 We employ the orthogonal chirp division multiplexing (OCDM) [44] waveform with a single Gaussian pulse shaping.
 The carrier frequency is $3$GHz.
-The lowpass equivalent is formulated by,
+The lowpass equivalent is formulated by
 $s_{n}(t)=(2/T^{2})^{1/4}e^{-\pi t^{2}/T^{2}}e^{j\pi M/T^{2}(t-(n-1)T/M)^{2}},$ (60)
 where $M$ is the number of OCDM chirps, and $T$ is proportional to the effective pulse time width.
-VIII-A CRLBs
+VII-A CRLB Validation
 We first verify the CRLB in comparison with the MLE results.
-To maintain consistency with existing works [38, 40, 41, 39, 36, 34, 35], we use the signal energy-to-noise ratio (SENR), defined as $\tilde{\delta}_{\rm cn}\triangleq P/\sigma_{\rm cn}^{2}$, instead of the SCNR $\delta_{\rm cn}$ in evaluating the CRLBs.
+To maintain consistency with existing works [28, 38, 29], we use the signal energy-to-noise ratio (SENR), defined as $\tilde{\delta}_{\rm cn}\triangleq P/\sigma_{\rm cn}^{2}$, instead of the SCNR $\delta_{\rm cn}$ in evaluating the CRLBs.
 Due to the high complexity of four-dimensional MLE, we consider a simple $4\times 3$ nonsymmetric multi-static radar configuration given in Fig.
 2.
-Consider the OCDM waveform with $M=16$, $T=10^{-2}$ s and a sampling rate of $f_{s}=1$ kHz.
-The velocity of the target is $(20,30)$ m/s.
+Without loss of generality, we consider the OCDM waveform with $M=16$, $T=10^{-2}$ s and a sampling rate of $f_{s}=1$ kHz.
 Fig.
@@
 Overall, these results validate the effectiveness and accuracy of the proposed CRLBs.
-VIII-B Power Allocation for ISAC
+VII-B Power Allocation for ISAC
 The performance of the proposed PA algorithms is evaluated by comparison with the PP-NCG and PP-NSD benchmark algorithms.
@@
 The PA budget is ${\bm{\rho}}_{\max}=0.3\times{\bm{1}}_{N}$.
-Consider a target moving slowly with velocity $l_{v}=[4,5]m/s$.
-A typical modulus squared RCS matrix is given by $\begin{bmatrix}0.37&0.70&1.38&0.65&2.40&0.18&0.25&0.82&0.14&0.35\\ 0.40&2.05&1.14&0.42&0.03&0.11&2.02&1.65&1.24&3.65\end{bmatrix}^{\rm T}$.
-The channel gains between the user and the transmitters are assumed to be $[2.11,12.57,5.63,0.75,0.61,1.75,0.20,2.34,14.79,9.68]^{\rm T}$.
-In this ISAC setup, the CRLB thresholds for the position and velocity are given by $[250,0.13]$.
+The SENR is set to be $-10$ dB.
+Within the coherence time for RCS, a typical modulus-squared RCS matrix is given by $\begin{bmatrix}0.37&0.70&1.38&0.65&2.40&0.18&0.25&0.82&0.14&0.35\\ 0.40\!&\!2.05&1.14&0.42&0.03&0.11&2.02&1.65&1.24&3.65\end{bmatrix}^{\rm T}$.
+Similarly, the channel gains between the user and the transmitters are assumed to be ${\bm{g}}=[2.11,12.57,5.63,0.75,0.61,1.75,0.20,2.34,14.79,9.68]^{\rm T}$.
+We consider the OCDM waveform with $M=16$, $T=10^{-2}$ s and a sampling rate of $f_{s}=10$ kHz.
+Under the given ISAC system setup and waveform configuration, the CRLB thresholds for position and velocity are set to $[250,0.13]$.
 Figure 2: The AP location distribution of the ISAC system: (a) $4\times 3$ radar setting; (b) CF MIMO-ISAC setting.
 Figure 3: The MSE versus the CRLB under different SENRs.
-The PA vector is initialized as the normalized PA vector ${\bm{\rho}_{1}}={\bm{1}}_{N}/N$, unless noted otherwise.
+The PA vector is initialized as the uniform PA vector ${\bm{\rho}_{1}}={\bm{1}}_{N}/N$, unless noted otherwise.
 The initial penalty factor is set to be $\mu_{1}=10^{4}$.
@@
 The other simulation conditions are given in Table I.
-Note that the presented SINRs, PA vectors, and CRLBs, are obtained by averaging 100 points after convergence.
-Firstly, we evaluate the convergence and steady-state performance of the proposed algorithms.
-The SINR learning curves of four algorithms are illustrated in Fig.
+We begin by evaluating the convergence and steady-state performance of the proposed algorithms.
+The SINR results corresponding to the iterative PA vectors for all four algorithms are depicted in Fig.
 4.
-As apparently seen in Fig.
-4, the PP-MCG-ILS has the highest convergence rate, followed by the PP-MSD-ILS, while the counterparts without ILS presents comparatively slow convergence.
-Additionally, the benchmark algorithms without ILS exhibit numerical fluctuations, particularly under large fixed step sizes.
-This behavior suggests potential instability and divergence risks.
-Consequently, we refrain from increasing the step sizes further during simulations to avoid compromising convergence reliability.
-The evolution of the PA coefficients and corresponding CRLB values with respect to the number of iterations is illustrated in Fig.
-5 and Fig.
-6, respectively.
-The results show that the PP-MCG-ILS algorithm achieves significantly faster convergence compared to the PP-NCG algorithm.
-Ultimately, all four algorithms converge to a common PA vector, given by ${\bm{\rho}}=[0.01,0.30,0.0638,0.01,\dots,0.01,0.30,0.2762]^{\rm T}$, which satisfies the sensing QoS constraints.
-Notably, more power is allocated to transmitters 2, 9, and 11 due to their higher communication channel gains.
-Furthermore, the steady-state CRLB values remain within the predefined threshold limits, confirming the validity and effectiveness of the proposed optimization framework.
+As observed, the PP-MCG-ILS algorithm exhibits the fastest convergence, followed by the PP-MSD-ILS.
+In contrast, the benchmark algorithms converge significantly more slowly.
+Furthermore, the benchmarks without ILS display noticeable numerical fluctuations, particularly under large fixed step sizes, suggesting potential instability and the risk of divergence.
+To ensure convergence reliability, we refrain from further increasing the step sizes in these simulations.
+[…]
 TABLE I: Simulation parameters
@@
 Figs. 9 and 10 $\{\epsilon_{\rm th},\sigma_{\upsilon}\}=\{1e-11,0.7\}$
-Figure 4: SINR with sequential penalty factor
+Figure 4: SINR convergence behavior.
 Figure 5: PA coefficient: (a) PP-MCG-ILS; (b) PP-NCG with $\upsilon=2\times 10^{-5}$.
 Figure 6: CRLB values: (a) PP-MCG-ILS; (b) PP-NCG with $\upsilon=2\times 10^{-5}$.
-The proposed sequential penalty-based algorithms leverage the optimal solution from the previous iteration loop as the initial point for the current loop under an updated penalty factor.
-To further explore the impact of initialization, we compare the performance of different initial PA strategies under a fixed penalty factor.
+The sequential penalty-based algorithms above leverage the optimal solution from the previous iteration loop as the initial point for the current loop under an updated penalty factor.
+To further explore the impact of initialization, we consider fixing the penalty factor and comparing the ISAC performance under different PA initialization strategies.
 In addition to the commonly used uniform PA initialization, we introduce a heuristic PA scheme, where the initial allocation is proportional to the communication channel gain, given by ${\bm{\rho}}_{1}={\bm{g}}/\sum_{n}|g_{n}|^{2}$.
@@
 However, the PP-MSD-ILS and PP-NSD algorithms show degraded convergence, highlighting the superior convergence properties of CG-based algorithms over SD-based methods under identical conditions.
-When employing the heuristic PA initialization, all algorithms demonstrate significantly improved convergence rates compared to the uniform PA case.
+When employing the heuristic PA initialization, all algorithms demonstrate significantly better convergence rates compared to the uniform PA case.
 Moreover, both PP-NCG and PP-NSD achieve noticeable performance gains relative to their sequential penalty-based counterparts, particularly PP-NCG.
-In contrast, the PP-MCG-ILS algorithm shows robust and superior convergence behavior, and PP-MSD-ILS exhibits a slight degradation.
+In contrast, the PP-MCG-ILS algorithm shows robust and superior convergence behavior, and PP-MSD-ILS exhibits a slight degradation, relative to their sequential penalty-based counterparts.
 Figure 7: SINR with initial uniform PA
@@
 9, where the achieved location CRLB value closely approaches its threshold and then converge.
+These trends reveal the trade-off between the communication SINR and the location CRLBs.
 Furthermore, it is noteworthy that the velocity CRLB remains largely unaffected by variations in the location CRLB thresholds.
@@
 Figure 11: Learning results for pure sensing: (a) total power; (b) CRLBs.
-VIII-C Power Allocation for pure sensing
-We now simply verify the effectiveness of the proposed P-MCG-ILS algorithm for total power minimization in joint location and velocity estimation.
+VII-C Power Allocation for Pure Sensing
+We now verify the effectiveness of the proposed P-MCG-ILS algorithm for total power minimization in pure sensing scenario with joint location and velocity estimation.
 The initial penalty factor is set to be $\mu_{1}=1$, $\sigma_{\upsilon}=0.2$ and the penalty threshold $\epsilon_{\mu}=10^{-6}$.
@@
 4.
-The learning results for total power optimization and the corresponding CRLB performance are illustrated in Fig.
+The iterative results for total power optimization and the corresponding CRLB performance are illustrated in Fig.
 11.
@@
 This difference arises because, in pure sensing, any enhancement in CRLB performance directly benefits from increased sensing power.
-However, in ISAC, power allocation is optimized primarily for communication performance, with CRLB constraints being satisfied, potentially loosely, within acceptable margins.
-IX Conclusions
-This paper presents, for the first time, a comprehensive investigation of the power allocation problem in a cell-free (CF) MIMO-ISAC framework, where joint location and velocity estimation is considered as the sensing task.
-We employ newly derived approximate Cramér-Rao lower bounds (CRLBs) as sensing performance metrics, alongside user receive SINR as the communication metric.
-To address the inherent nonlinearity and non-convexity of the resulting optimization problem, we propose a penalty function and projection-based modified conjugate gradient algorithm with inexact line search (PP-MCG-ILS).
-Additionally, we introduce a penalty function-based normalized conjugate gradient algorithm with inexact line search (P-NCG-ILS) to handle the power minimization problem in pure sensing scenarios.
-We analyze the convergence behavior and qualitatively compare the computational complexity of all proposed methods.
-Simulation results confirm the accuracy of the derived CRLBs and demonstrate the efficacy of the proposed power allocation strategies in enhancing pure sensing and overall ISAC performance.
-The findings offer valuable insights into resource-efficient design for next-generation CF MIMO-ISAC networks and lay the groundwork for future extensions involving more complex sensing tasks and dynamic resource management.
-Appendix A Proof of Lemma 1
-It is obvious that $\|\Delta{\bm{\rho}}_{i+1}\|\rightarrow 0$ when $i>I$ if and only if either ${\bm{d}}_{i}\rightarrow{\bm{0}}$ or $v_{i}\rightarrow 0$ when $i>I$.
-On the other hand, in light of (46), we have
-$\Delta L({\bm{\rho}}_{i+1})\approx\upsilon_{i}{\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i})+\upsilon_{i}^{2}/2{\bm{d}}_{i}^{\rm T}\nabla^{2}L({\bm{\rho}}_{i}){\bm{d}}_{i}.$ (65)
-Firstly, if ${\bm{d}}_{i}\rightarrow{\bm{0}}$ or $v_{i}\rightarrow 0$ for $i>I$, we readily have $\Delta L({\bm{\rho}}_{i+1})\rightarrow 0$.
+However, in ISAC, PA is optimized primarily for communication performance, with CRLB constraints being satisfied, potentially loosely, within acceptable margins.
+VIII Conclusions
+This paper presents, for the first time, a comprehensive investigation of the power allocation problem in a CF MIMO-ISAC framework, where joint location and velocity estimation were considered as the sensing task.
+We employed newly derived approximate CRLBs as sensing performance metrics, alongside user receive SINR as the communication metric.
+To address the inherent nonlinearity and non-convexity of the resulting optimization problem, we propose the PP-MCG-ILS algorithm.
+Additionally, the P-NCG-ILS algorithm was further proposed to handle the power minimization problem in pure sensing scenarios.
+We devoted into the detailed analysis of the convergence behavior and qualitatively compare the computational complexity of all proposed methods.
+Simulation results confirmed the accuracy of the derived CRLBs and demonstrate the efficacy of the proposed power allocation strategies in enhancing pure sensing and overall ISAC performance.
+Our findings offer valuable insights into resource-efficient design for next-generation CF MIMO-ISAC networks and lay the groundwork for future extensions involving more complex sensing tasks and dynamic resource management.
+Appendix A Relevant parameters of the derived CRLBs
+The waveform-related parameters $a_{n,k}$, $b_{n,k}$ and $d_{n,k}$, are given by
+$\small[a_{n,k},\ b_{n,k},\ d_{n,k}]=\rho_{n}\big{[}\underbrace{8\pi^{2}S\bar{f_{n}^{2}}\delta_{\rm cn}}_{\bar{a}_{n,k}},\ \underbrace{4\pi S\mathfrak{T}\{\sigma_{tf}\}\delta_{\rm cn}}_{\bar{b}_{n,k}},\ \underbrace{8\pi^{2}S\bar{t_{n}^{2}}\delta_{\rm cn}}_{\bar{d}_{n,k}}\big{]}$ (61)
+where $S$ denotes the number of samples.
+In addition, the remaining parameters, including $\beta_{n,k}$, $\zeta_{n,k}$, $\xi_{n,k}$, $\varrho_{n,k}$, $\eta_{n,k}$, and $\kappa_{n,k}$, represent geometric spreading factors, given respectively by
+$[\beta_{n,k},\zeta_{n,k}]^{\rm T}=$ $\dfrac{|\alpha_{n,k}|}{c}\left(\dfrac{{\bm{l}}-{\bm{l}}_{n}}{\|{\bm{l}}_{n}-{\bm{l}}\|_{2}}+\dfrac{{\bm{l}}-{\bm{l}}_{k}}{\|{\bm{l}}_{k}-{\bm{l}}\|_{2}}\right),$ (62a)
+$[\xi_{n,k},\varrho_{n,k}]^{\rm T}=$ $\dfrac{f_{c}|\alpha_{n,k}|}{c}\left(\dfrac{{\bm{l}}_{n}-{\bm{l}}}{\|{\bm{l}}_{n}-{\bm{l}}\|_{2}}+\dfrac{{\bm{l}}_{k}-{\bm{l}}}{\|{\bm{l}}_{k}-{\bm{l}}\|_{2}}\right),$ (62b)
+$[\eta_{n,k},\kappa_{n,k}]^{\rm T}=$ $\dfrac{f_{c}|\alpha_{n,k}|}{c}\bigg{(}({\bm{l}}_{n}-{\bm{l}})\dfrac{{\bm{v}}^{\rm T}({\bm{l}}_{n}-{\bm{l}})}{\|{\bm{l}}_{n}-{\bm{l}}\|_{2}^{3}}-\dfrac{{\bm{v}}}{\|{\bm{l}}_{n}-{\bm{l}}\|_{2}}$
+$+({\bm{l}}_{k}-{\bm{l}})\dfrac{{\bm{v}}^{\rm T}({\bm{l}}_{k}-{\bm{l}})}{\|{\bm{l}}_{k}-{\bm{l}}\|_{2}^{3}}-\dfrac{{\bm{v}}}{\|{\bm{l}}_{k}-{\bm{l}}\|_{2}}\bigg{)}.$ (62c)
+Appendix B Proof of Lemma 1
+It is obvious that $\|\Delta{\bm{\rho}}_{i+1}\|\rightarrow 0$ if and only if either ${\bm{d}}_{i}\rightarrow{\bm{0}}$ or $v_{i}\rightarrow 0$ when $i>I$. On the other hand, in light of (42), we have
+$\small\Delta L({\bm{\rho}}_{i+1})\approx\upsilon_{i}{\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i})+\upsilon_{i}^{2}/2{\bm{d}}_{i}^{\rm T}\nabla^{2}L({\bm{\rho}}_{i}){\bm{d}}_{i}.$ (63)
+If ${\bm{d}}_{i}\rightarrow{\bm{0}}$ or $v_{i}\rightarrow 0$ for $i>I$, we readily have $\Delta L({\bm{\rho}}_{i+1})\rightarrow 0$.
 Assume there exists $i>I$, where $\Delta L({\bm{\rho}}_{i+1})\rightarrow 0$ with neither ${\bm{d}}_{i}\rightarrow{\bm{0}}$ nor $v_{i}\rightarrow 0$.
 We can solve the step size for $\Delta L({\bm{\rho}}_{i+1})=0$ as
-$\upsilon_{i}=\dfrac{-2{\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i})}{{\bm{d}}_{i}^{\rm T}\nabla^{2}L({\bm{\rho}}_{i}){\bm{d}}_{i}},$ (66)
+$\small\upsilon_{i}=\dfrac{-2{\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i})}{{\bm{d}}_{i}^{\rm T}\nabla^{2}L({\bm{\rho}}_{i}){\bm{d}}_{i}},$ (64)
 which corresponds to the step size bound in the ILS with $\epsilon_{L}=0$.
@@
 $\blacksquare$
-Appendix B Proof of Theorem 1
+Appendix C Proof of Theorem 1
 In light of (42), (44) and (63), the ILS condition requires
 $\upsilon_{i}{\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i})+\upsilon_{i}^{2}/2{\bm{d}}_{i}^{\rm T}\nabla^{2}L({\bm{\rho}}_{i}){\bm{d}}_{i}\leq\epsilon_{L}\upsilon_{i}{\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i})$
-$\Longrightarrow\ \upsilon_{i}\leq\dfrac{2(\epsilon_{L}-1){\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i})}{{\bm{d}}_{i}^{\rm T}\nabla^{2}L({\bm{\rho}}_{i}){\bm{d}}_{i}}.$ (67)
+$\Longrightarrow\ \upsilon_{i}\leq\dfrac{2(\epsilon_{L}-1){\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i})}{{\bm{d}}_{i}^{\rm T}\nabla^{2}L({\bm{\rho}}_{i}){\bm{d}}_{i}}.$
 where positive ${\bm{d}}_{i}^{\rm T}\nabla^{2}L({\bm{\rho}}_{i}){\bm{d}}_{i}$ is considered.
-Assuming the Lipschitz gradient continuity of $L({\bm{\rho}})$, $\|\nabla^{2}L({\bm{\rho}}_{i})\|$ is upper bounded, indicating a positive and finite ${\bm{d}}_{i}^{\rm T}\nabla^{2}L({\bm{\rho}}_{i}){\bm{d}}_{i}$ and thus non-zero $\upsilon_{i}$.
-Furthermore, in light of (48), we have,
-$\Delta^{2}L({\bm{\rho}}_{i+1})\leq\epsilon_{L}^{2}\upsilon_{i}^{2}({\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i}))^{2}.$ (68)
-Since $L({\bm{\rho}})$ is a monotonously decreasing and bounded function regarding ${\bm{\rho}}_{i}$, we have $\Delta^{2}L({\bm{\rho}}_{i+1})\rightarrow 0$ when $i>I$ for a large constant $I$.
-The deflection factor computed by (45) or the descent direction vector in (53) can guarantee the strict descent ${\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i})<0$ at each iteration $i$ unless ${\bm{d}}_{i}={\bm{0}}$.
+Assuming the Lipschitz gradient continuity of $L({\bm{\rho}})$, $\|\nabla^{2}L({\bm{\rho}}_{i})\|$ is upper bounded.
+This indicates a positive and finite ${\bm{d}}_{i}^{\rm T}\nabla^{2}L({\bm{\rho}}_{i}){\bm{d}}_{i}$ and thus non-zero $\upsilon_{i}$.
+Furthermore, in light of (44), we have
+$\small\Delta^{2}L({\bm{\rho}}_{i+1})\leq\epsilon_{L}^{2}\upsilon_{i}^{2}({\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i}))^{2}.$ (65)
+Since $L({\bm{\rho}})$ is a monotonously decreasing and bounded function regarding ${\bm{\rho}}_{i}$, the (conjugate) gradient-based descent algorithm converges to $\Delta^{2}L({\bm{\rho}}_{i+1})\rightarrow 0$ when $i>I$ for a large constant $I$.
+However, the deflection factor computed by (41) or the descent direction vector in (49) can guarantee a strict descent ${\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i})<0$ at each iteration $i$ unless ${\bm{d}}_{i}={\bm{0}}$.
 Consequently, sequence $\{L({\bm{\rho}}_{i})\}$ will converge to a stationary point with a zero descent direction ${\bm{d}}_{i}\rightarrow{\bm{0}}$ when it has Lipschitz gradient continuity.
@@
 On the other hand, we assume the objective function is not Lipschitz gradient continuous.
-In this case, if the sequence $\{{\bm{\rho}}_{i}\}$ pass a Lipschitz noncontinuous point ${\bm{\rho}}_{I}$ where ${\bm{d}}_{I}^{\rm T}\nabla^{2}L({\bm{\rho}}_{I}){\bm{d}}_{I}$ is extremely large, to satisfy (67), the step size should meet $\upsilon_{I}\rightarrow 0$.
+In this case, if the sequence $\{{\bm{\rho}}_{i}\}$ pass a Lipschitz noncontinuous point ${\bm{\rho}}_{I}$ where ${\bm{d}}_{I}^{\rm T}\nabla^{2}L({\bm{\rho}}_{I}){\bm{d}}_{I}$ is extremely large, to satisfy (C), the step size should meet $\upsilon_{I}\rightarrow 0$.
 According to Lemma 1, we have $\|\Delta{\bm{\rho}}_{i+1}\|\rightarrow 0$ and $\Delta L({\bm{\rho}}_{i+1})\rightarrow 0$ when $\upsilon_{i}\rightarrow 0$ for $i\geq I$, i.e., the sequences $\{{\bm{\rho}}_{i}\}$ and $\{L({\bm{\rho}}_{i})\}$ converge to a Lipschitz gradient noncontinuous point.

### v3

@@
 Abstract
-In this paper, we investigate integrated sensing and communication (ISAC) in a cell-free (CF) multiple-input multiple-output (MIMO) network, where each access point functions either as an ISAC transmitter or as a sensing receiver.
-We devote into the ISAC sensing metric using the discrete-time signal-based Cramér–Rao lower bounds (CRLBs) for joint location and velocity estimation under arbitrary power allocation ratios under the deterministic radar cross section assumption (RCS).
-Then, we consider the power allocation optimization problem for the CF MIMO-ISAC as the maximization of the communication signal-to-interference-plus-noise ratio (SINR), subject to CRLB-based sensing constraints and per-transmitter power limits.
-To solve the resulting nonlinear and non-convex problem, we propose a penalty function and projection-based modified conjugate gradient algorithm with inexact line search (PP-MCG-ILS), and an alternative method based on a modified steepest descent approach (PP-MSD-ILS).
-We show that the proposed algorithms are scalable and can be extended to a broad class of optimization problems involving nonlinear inequality constraints and affine equality constraints.
-In addition, we extend the PP-MCG-ILS algorithm to the pure sensing scenario, where a penalty function-based normalized conjugate gradient algorithm (P-NCG-ILS) is developed for sensing power minimization.
-Finally, we analyze the convergence behavior and qualitatively compare the computational complexity of the proposed algorithms.
-Simulation results confirm the accuracy of the derived CRLBs and demonstrate the effectiveness of the proposed power allocation strategies in enhancing both sensing and overall ISAC performance.
+This paper investigates integrated sensing and communication (ISAC) in a cell-free (CF) multiple-input multiple-output (MIMO) network, in which each access point operates either as an ISAC transmitter or as a sensing receiver.
+Discrete-time signal-based Cramér–Rao lower bounds (CRLBs) are employed as sensing performance metrics for joint location and velocity estimation under arbitrary power allocation ratios, assuming a deterministic radar cross section (RCS) model.
+Based on these metrics, a power allocation problem for CF MIMO-ISAC is formulated to maximize the communication signal-to-interference-plus-noise ratio (SINR), subject to CRLB-based sensing constraints and per-transmitter power limits.
+To address the resulting nonlinear and nonconvex optimization problem, a penalty-function and projection-based modified conjugate gradient algorithm with inexact line search (PP-MCG-ILS) is developed, together with an alternative method based on modified steepest descent (PP-MSD-ILS).
+The proposed algorithms are applicable to a class of constrained optimization problems involving nonlinear inequality constraints and affine equality constraints.
+In addition, the PP-MCG-ILS algorithm is extended to a sensing-only scenario, where a penalty-function-based normalized conjugate gradient algorithm (P-NCG-ILS) is employed for sensing power minimization.
+The convergence behavior of the proposed algorithms is analyzed, and their computational complexity is qualitatively evaluated.
+Simulation results validate the accuracy of the derived CRLBs and demonstrate the effectiveness of the proposed power allocation strategies in enhancing sensing performance and overall ISAC performance.
 Index Terms:
@@
 Our prior work [40] addresses this gap by introducing a discrete-time likelihood function over Gaussian channels, offering a more realistic modeling framework.
-Building on this foundation, this paper devotes into the CRLBs derivation for joint location and velocity estimation under the deterministic RCS assumption, covering both single- and multi-target cases in CF MIMO-ISAC systems.
+Building on this foundation, this paper devotes into the CRLBs derivation for joint location and velocity estimation under the deterministic RCS assumption in CF MIMO-ISAC systems.
 Beyond performance characterization, another critical challenge lies in resource allocation, involving time [43], frequency [12], space [23], and power [9, 13, 1, 22, 2].
@@
 These considerations motivate us to consider a more comprehensive and realistic setting: joint location and velocity estimation in CF MIMO-ISAC systems under a discrete-time signal model with a deterministic RCS assumption.
-The CRLBs are serve as effective sensing metrics to guide PA across distributed APs, with the goal of maximizing communication SINR.
+The CRLBs are served as effective sensing metrics to guide PA across distributed APs, with the goal of maximizing communication SINR.
 This leads to a class of non-convex and highly nonlinear optimization problems.
 To address this issue, we develop a tailored optimization framework that integrates penalty methods, projection operations, and gradient-based search, enabling efficient and reliable algorithmic solutions under complex practical constraints.
-The main contributions are as follows:
-  • We derive novel closed-form approximate CRLBs for joint location and velocity estimation under arbitrary PA strategies.
-These CRLBs are rigorously validated against maximum likelihood estimation (MLE) results, confirming their effectiveness as practical sensing performance metrics.
-Based on these metrics, we formulate the PA optimization problems for ISAC, aiming at communication SINR maximization while ensuring joint CRLB and power budget constraints.
-  • To tackle the non-convexity and nonlinearity in PA optimization for ISAC, we propose a penalty-function and projection-based modified conjugate gradient algorithm with inexact line search (PP-MCG-ILS) together with PP-MSD-ILS algorithm based on the modified steepest descent algorithm (MSD).
-We further propose a penalty-function-based normalized modified conjugate gradient algorithm with inexact line search (P-NCG-ILS), which can effectively solve the power minimization problem subject to sensing accuracy constraints.
-  • We provide a comprehensive analysis of their convergence behavior and computational complexity.
-The design ensures stable updates, rapid convergence and low computational overhead.
-Additionally, we show that the proposed PP-MCG-ILS and P-NCG-ILS algorithms can be broadly scalable to a class of constrained non-convex optimization problems involving affine equality constraints (handled via projection) and nonlinear inequality constraints (handled via penalties).
-  • Extensive simulations validate the performance of the developed algorithms in both ISAC and sensing-only scenarios.
-The PP-MCG-ILS algorithm consistently exhibits fast convergence, robust numerical stability, and optimal PA efficiency across various conditions, as compared to benchmark methods.
-Results also highlight important trade-offs between sensing and communication performance and confirm that the proposed methods can meet stringent sensing QoS while maximizing communication throughput or minimizing power usage.
+The main contributions of this work are summarized as follows:
+  • We develop tractable closed-form approximations of the CRLB for joint location and velocity estimation under arbitrary power allocation (PA) strategies.
+The accuracy of the proposed approximations is validated through comparisons with maximum-likelihood estimation (MLE) results, indicating their suitability as sensing performance metrics for ISAC-oriented resource allocation.
+Based on these metrics, PA optimization problems are formulated to maximize communication SINR subject to sensing accuracy and power budget constraints.
+  • To address the nonconvex and nonlinear nature of the resulting PA optimization problems, we introduce a penalty-function and projection-based modified conjugate gradient algorithm with inexact line search (PP-MCG-ILS), along with a PP-MSD-ILS algorithm based on modified steepest descent.
+In addition, a penalty-function-based normalized conjugate gradient algorithm with inexact line search (P-NCG-ILS) is developed for power minimization in sensing-only scenarios under prescribed sensing accuracy constraints.
+  • We provide a detailed analysis of the convergence behavior and computational complexity of the proposed algorithms.
+Under standard smoothness assumptions, the analysis establishes well-defined step size selection, stable iterative updates, and efficient numerical behavior.
+The algorithmic framework also admits straightforward extension to related constrained nonconvex optimization problems.
+  • Simulation results are provided to evaluate the proposed algorithms in both ISAC and sensing-only scenarios.
+The results demonstrate that the proposed methods achieve reliable convergence and competitive PA performance relative to representative benchmark schemes, while revealing the inherent trade-offs between sensing accuracy and communication performance.
 Section II introduces the received signal model for the CF MIMO-ISAC system.
-Section III-B formulates the power allocation problem for CF MIMO-ISAC.
-Section IV presents the proposed power allocation algorithms for both CF MIMO-ISAC and the sensing-only scenario.
+Section III formulates the PA problem for CF MIMO-ISAC based on the proposed metrics.
+Section IV presents the proposed PA algorithms for both CF MIMO-ISAC and the sensing-only scenario.
 Section VI analyzes the convergence properties and computational complexity of the proposed algorithms.
@@
 Finally, Section VIII concludes the paper.
-[…]
+Notation: $\mathbb{R}$ and $\mathbb{C}$ denote the field of real and complex numbers, respectively.
+Scalars are denoted by lower-case letters, vectors and matrices respectively by lower- and upper-case boldface letters.
+The conjugate and transpose are denoted by $(\cdot)^{*}$ and $(\cdot)^{\rm T}$, respectively.
+$\mathbb{E}\{\cdot\}$ and $|\cdot|$ denote the mathematical expectation and element-wise modulus, respectively.
+$\|\cdot\|$ denotes the $l_{2}$ norm of a matrix.
+The notations $\min\{\cdot\}$ and $\max\{\cdot\}$ denote the minimum and maximum element of the enclosed set $\{\cdot\}$, respectively.
 II Signal Model
@@
 Figure 1: CF MIMO system for ISAC.
-The Doppler frequency induced by the relative radial velocity of the target with respect to transmitter $n$ and receiver $k$ is computed as
-$f_{n,k}=\dfrac{1}{\lambda}{\bm{v}}^{\rm T}\left(\frac{{\bm{l}}_{n}-{\bm{l}}}{\|{\bm{l}}_{n}-{\bm{l}}\|_{2}}+\frac{{\bm{l}}_{k}-{\bm{l}}}{\|{\bm{l}}_{k}-{\bm{l}}\|_{2}}\right),$ (1)
-where $\lambda$ denotes the wavelength. Denote $\tau_{n,k}$ as the propagation delay between the $n$th transmitter and the $k$th receiver, as reflected by the target. Then, we have
+Denote $\tau_{n,k}$ as the propagation delay between the $n$th transmitter and the $k$th receiver, as reflected by the target. Then, we have
 $\tau_{n,k}$ $=\tau_{n}+\tau_{k},$ (1a)
@@
 where $\tau_{n}$ and $\tau_{k}$ represent the propagation delays from the $n$th transmitter to the $q$th target, and from the target to the $k$th receiver, respectively, and $c$ denotes the speed of light.
-[…]
+The Doppler frequency induced by the relative radial velocity of the target with respect to transmitter $n$ and receiver $k$ is
+$f_{n,k}=\dfrac{1}{\lambda}{\bm{v}}^{\rm T}\left(\frac{{\bm{l}}_{n}-{\bm{l}}}{\|{\bm{l}}_{n}-{\bm{l}}\|_{2}}+\frac{{\bm{l}}_{k}-{\bm{l}}}{\|{\bm{l}}_{k}-{\bm{l}}\|_{2}}\right),$ (2)
+where $\lambda$ denotes the wavelength.
+A waveform $s_{n}(t)$, with unit energy, i.e., $\varepsilon_{s}=\int|{s}_{n}(t)|^{2}\,dt=1$, is considered for both sensing and communication at the $n$th transmitter.
+Then, the baseband ISAC transmitted signal at transmitter $n$ is denoted as
 ${\bm{x}}_{n}(t)=\sqrt{P\rho_{n}}{b}_{n}\varsigma s_{n}(t),$ (3)
-where $P$ denotes the total available energy, $\rho_{n}$ is the PA factor with $\sum_{n=1}^{N}\rho_{n}=1$, ${b}_{n}\in\mathbb{C}$ is the normalized transmit beamforming weight with $|{b}_{n}|^{2}=1$, and $\varsigma\in\mathbb{C}$ has unit variance $\mathbb{E}\{|\varsigma|^{2}\}=1$.
-Let ${\bm{\rho}}=[\rho_{1},\rho_{2},\cdots,\rho_{N}]^{\rm T}$ be the vector of the PA coefficient of the $N$ transmitters.
-Sensing model: For target-reflected reception, we refer the term “RCS” as encompassing both channel fading and radar RCS.
+where $P$ denotes the total available energy, $\rho_{n}$ is the PA factor with $\sum_{n=1}^{N}\rho_{n}=1$, ${b}_{n}\in\mathbb{C}$ is the normalized precoding weight with $|{b}_{n}|^{2}=1$, and $\varsigma\in\mathbb{C}$ has unit variance $\mathbb{E}\{|\varsigma|^{2}\}=1$.
+Let ${\bm{\rho}}=[\rho_{1},\rho_{2},\cdots,\rho_{N}]^{\rm T}$ be the PA coefficient vector over the $N$ transmitters.
+Sensing model: For target-reflected reception, we hereafter refer to the term “RCS” as encompassing both channel fading and radar RCS.
 Denote the RCS associated with transmitter $n$ and receiver $k$ as ${\alpha}_{n,k}$.
-We consider the Swerling-I sensing model [37], i.e.
-the deterministic RCS assumption where RCSs are assumed to be deterministic but unknown during the observation time [29, 28].
+The sensing model under the deterministic RCS assumption is considered where RCSs are assumed to be deterministic but unknown during the observation time [3, 16, 38].
 Then, the lowpass equivalent of the reflected signal from the $n$th transmitter by a target received at receiver $k$ is given by [^1] The transmitted signals over different transmitters are orthogonal and maintain approximate orthogonality over different delays and Doppler frequencies such that the signals contributed from different transmitters can be separated at each receiver [3, 19].
@@
 $\small{r}_{u}(t)$ $=\sqrt{P}\sum\nolimits_{n=1}^{N}g_{n}{b}_{n}\sqrt{\rho_{n}}s_{n}(t)+{z}(t),$ (6)
-where $g_{n}$ is the channel fading between the transmitter $n$ and the user, and $z(t)$ denotes the interference-plus-noise with zero mean and variance $\sigma_{z}^{2}$.
+where $g_{n}$ is the channel fading [^2] The communication channels are assumed to be estimated a priori using pilot signals [9, 22, 1, 2].
+between the transmitter $n$ and the user, and $z(t)$ denotes the interference-plus-noise with zero mean and variance $\sigma_{z}^{2}$.
 Then, the SINR at the user terminal can be given by
 $\small\gamma$ $=\sum\nolimits_{n=1}^{N}|g_{n}{b}_{n}|^{2}\rho_{n}\delta\overset{(\text{i})}{=}\delta{\bm{\rho}}^{\rm T}{\bm{g}}$ (7)
-where $\delta=\frac{P}{T_{\rm eff}\sigma_{z}^{2}}$ is the total SINR and ${\bm{g}}$ is the channel gain [^2] The communication channels are assumed to be estimated a priori using pilot signals [24, 35, 34, 36].
-with its $n$th element being $|g_{n}|^{2}$.
-Note that $(\mathrm{i})$ holds when using a normalized conjugate beamforming with ${b}_{n}={g}^{*}_{n}/|g_{n}|$.
+where $\delta=\frac{P}{T_{\rm eff}\sigma_{z}^{2}}$ is the total loss-free SINR and $\bm{g}$} is the squared channel fading with its $n$th element being $|g_{n}|^{2}$.
+Note that $(\mathrm{i})$ holds when using a normalized conjugate precoding with ${b}_{n}={g}^{*}_{n}/|g_{n}|$.
+Because both sensing and communication channels are assumed quasi-static/deterministic within the coherence observation time, the deterministic RCS model and the estimated communication gains remain valid for each optimization cycle.
 III The Proposed Sensing Metric and Problem Formulation
-In this section, we present an approximate CRLB for CF MIMO radar under arbitrary power allocations.
+In this section, we propose a sensing metric of the joint location and velocity estimation based on our derived approximate CRLB for CF MIMO radar under arbitrary PAs.
 To begin with, we introduce the following definitions. Define the squared effective bandwidth (SEBW) and squared effective pulse time width (SETW) as
-[…]
-[…]
-[…]
-[…]
-[…]
+$\small\bar{f_{n}^{2}}=\frac{1}{4\pi^{2}}\int\bigg|\frac{d{s_{n}(t-\tau_{n,k})}}{d{\tau_{n,k}}}\bigg|^{2}\,dt=\int f^{2}|S_{n}(f)|^{2}\,df,$ (8)
+and $\bar{t_{n}^{2}}=\int t^{2}|s_{n}(t)|^{2}\,dt$, respectively. The cross-term between time and frequency is further defined as
+$\small\sigma_{tf}=\int{ts_{n}^{*}(t-\tau_{n,k})}\frac{d{s_{n}(t-\tau_{n,k})}}{d{\tau_{n,k}}}\,dt.$ (9)
+$|s_{n}(t)|^{2}$ and $|S_{n}(f)|^{2}$ can be interpreted as the probability density functions for consecutive random variables $t$ and $f$, respectively, since $\int|S_{n}(f)|^{2}\ df=\int|s_{n}(t)|^{2}\ dt=1$.
+Based on this interpretation, we define the average time and average frequency of the transmission waveform as
+$\small\bar{f}_{n}=\int f|S_{n}(f)|^{2}\ df\quad\text{and}\quad\bar{t}_{n}=\int t|s_{n}(t)|^{2}\ dt,$ (10)
 respectively.
@@
 Unlike the sensing performance metrics employed in [9, 13, 1, 22, 2], which only consider target detection or location estimation, we consider the joint location and velocity CRLB as the sensing metric, developed based on the discrete-time reception of the reflected signal under the deterministic RCS assumption.
-In particular, we have derived the approximate CRLBs for both location and velocity in [30] for simplified sensing metrics.
-They are respectively given by
+We derive the approximate CRLBs of the joint location and velocity estimation [^3] For completeness, the detailed derivations are referred to [40]., respectively, given by
 $\small{\bm{C}}_{\rm L}$ $=({\bm{P}}-{\bm{V}}{\bm{Y}}^{-1}{\bm{V}}^{\rm T})^{-1},$ (11)
@@
 $\small{\bm{C}}_{\rm V}$ $=({\bm{Y}}-{\bm{V}}^{\rm T}{\bm{P}}^{-1}{\bm{V}})^{-1}.$ (12)
-The detailed expression of ${\bm{P}}$ and ${\bm{V}}$ are given at the top of the page, respectively, and
+where $\bm{P}$} and $\bm{V}$} are presented at the top of the page, respectively, and
 ${\bm{P}}=\sum_{n=1}^{N}\sum_{k=1}^{K}\begin{bmatrix}\beta_{n,k}^{2}a_{n,k}+2\beta_{n,k}\eta_{n,k}b_{n,k}+\eta_{n,k}^{2}d_{n,k}&\beta_{n,k}\zeta_{n,k}a_{n,k}\!+\!(\beta_{n,k}\kappa_{n,k}\!+\!\zeta_{n,k}\eta_{n,k})b_{n,k}+\eta_{n,k}\kappa_{n,k}d_{n,k}\\ \zeta_{n,k}\beta_{n,k}a_{n,k}\!+\!(\zeta_{n,k}\eta_{n,k}+\beta_{n,k}\kappa_{n,k})b_{n,k}+\kappa_{n,k}\eta_{n,k}d_{n,k}&\zeta_{n,k}^{2}a_{n,k}+2\zeta_{n,k}\kappa_{n,k}b_{n,k}\!+\!\kappa_{n,k}^{2}d_{n,k}\end{bmatrix}$,
@@
 $\small{\bm{Y}}=\sum_{n=1}^{N}\sum_{k=1}^{K}\begin{bmatrix}\xi_{n,k}^{2}d_{n,k}&\xi_{n,k}\varrho_{n,k}d_{n,k}\\ \varrho_{n,k}\xi_{n,k}d_{n,k}&\varrho_{n,k}^{2}d_{n,k}\end{bmatrix}$. (13)
-The relevant waveform-related parameters $a_{n,k}$, $b_{n,k}$ and $d_{n,k}$, and geometric-spreading-related factors $\beta_{n,k}$, $\zeta_{n,k}$, $\xi_{n,k}$, $\varrho_{n,k}$, $\eta_{n,k}$, and $\kappa_{n,k}$, are detailed in Appendix A.
-As observed from (61), the PA coefficients play a pivotal role in the formulation of the CRLBs.
-Meanwhile, these coefficients directly affect communication SINR, as described in (7).
-Unfortunately, the PA optimization relevant to the CRLB is intractable due to the implicit and complex forms of the expressions in (11) and (12).
-To facilitate analysis and optimization, we reformulate the CRLB expressions into a more explicit and tractable form.
-We denote the $(i,j)$th elements of matrices ${\bm{P}}$, ${\bm{V}}$ and ${\bm{Y}}$ as $p_{ij}$, $v_{ij}$ and $y_{ij}$, respectively. Then, one can simplify them as
+The waveform-related parameters $a_{n,k}$, $b_{n,k}$ and $d_{n,k}$, and geometric-spreading-related factors $\beta_{n,k}$, $\zeta_{n,k}$, $\xi_{n,k}$, $\varrho_{n,k}$, $\eta_{n,k}$, and $\kappa_{n,k}$, are detailed in Appendix A.
+As described in (7), the PA coefficients directly affect communication SINR.
+However, the PA optimization relevant to the CRLB is intractable due to the implicit and complex forms of the expressions in (11) and (12).
+To facilitate analysis and optimization, we reformulate the CRLB expressions into a more explicit and tractable form regarding the PA coefficients.
+We denote the $(i,j)$th elements of matrices $\bm{P}$}, $\bm{V}$} and $\bm{Y}$} by $p_{ij}$, $v_{ij}$ and $y_{ij}$, respectively. They can be rewritten as
 $\small\begin{cases}p_{ij}=\sum_{n=1}^{N}w_{n}^{p_{ij}}\rho_{n}={\bm{\rho}}^{\rm T}{\bm{w}}_{p_{ij}},\\ v_{ij}=\sum_{n=1}^{N}w_{n}^{v_{ij}}\rho_{n}={\bm{\rho}}^{\rm T}{\bm{w}}_{v_{ij}},\\ y_{ij}=\sum_{n=1}^{N}w_{n}^{y_{ij}}\rho_{n}={\bm{\rho}}^{\rm T}{\bm{w}}_{y_{ij}},\end{cases}$ (14)
 where $w_{n}^{p_{ij}}$, $w_{n}^{v_{ij}}$ and $w_{n}^{y_{ij}}$ are the $n$th element of the weight vector ${\bm{w}}_{p_{ij}}$, ${\bm{w}}_{v_{ij}}$ and ${\bm{w}}_{y_{ij}}$, respectively.
-These weights can be readily obtained from the definitions of ${\bm{P}}$, ${\bm{V}}$ and ${\bm{Y}}$, by extracting their PA coefficient $\rho_{n}$.
-For example, from (13), we have $w_{n}^{p_{11}}$ as
+These weights can be readily obtained by extracting $\rho_{n}$ from the element-wise definitions of $\bm{P}$}, $\bm{V}$} and $\bm{Y}$}.
+For example, we have
 $\small w_{n}^{p_{11}}=\sum\nolimits_{k=1}^{K}(\beta_{n,k}^{2}\bar{a}_{n,k}+2\beta_{n,k}\eta_{n,k}\bar{b}_{n,k}+\eta_{n,k}^{2}\bar{d}_{n,k}),$ (15)
@@
 $\small{\bm{C}}_{\rm L}=\frac{1}{|{\bm{U}}|}\begin{bmatrix}u_{22}&-u_{12}\\ -u_{21}&u_{11}\end{bmatrix}$, (16)
-where $|{\bm{U}}|=u_{11}u_{22}-u_{12}u_{21}$. The relevant elements $u_{ij}$ are given respectively by
-$\small u_{11}=\frac{1}{|{\bm{Y}}|}\big{[}{\bm{\rho}}^{\rm T}({\bm{w}}_{p_{11}}{\bm{\rho}}^{\rm T}{\bm{W}}_{y}\!-\!{\bm{w}}_{v_{11}}{\bm{\rho}}^{\rm T}{\bm{W}}_{12}^{y}\!-\!{\bm{w}}_{v_{12}}{\bm{\rho}}^{\rm T}\breve{{\bm{W}}}_{11}^{y}){\bm{\rho}}\big{]},$ (17)
+where $|{\bm{U}}|=u_{11}u_{22}-u_{12}u_{21}$. The elements $u_{ij}$ are given by
+$u_{11}=\frac{1}{|{\bm{Y}}|}\big[{\bm{\rho}}^{\rm T}({\bm{w}}_{p_{11}}{\bm{\rho}}^{\rm T}{\bm{W}}_{y}\!-\!{\bm{w}}_{v_{11}}{\bm{\rho}}^{\rm T}{\bm{W}}_{12}^{y}\!-\!{\bm{w}}_{v_{12}}{\bm{\rho}}^{\rm T}\breve{{\bm{W}}}_{11}^{y}){\bm{\rho}}\big],$ (17)
 $u_{12}=\frac{1}{|{\bm{Y}}|}\big[{\bm{\rho}}^{\rm T}({\bm{w}}_{p_{12}}{\bm{\rho}}^{\rm T}{\bm{W}}_{y}\!-\!{\bm{w}}_{v_{21}}{\bm{\rho}}^{\rm T}{\bm{W}}_{12}^{y}\!-\!{\bm{w}}_{v_{22}}{\bm{\rho}}^{\rm T}\breve{{\bm{W}}}_{11}^{y}){\bm{\rho}}\big],$
@@
 $\small\begin{cases}{\bm{W}}_{p}={\bm{w}}_{p_{11}}{\bm{w}}_{p_{22}}^{\rm T}-{\bm{w}}_{p_{12}}{\bm{w}}_{p_{21}}^{\rm T},\\ {\bm{W}}^{p}_{12}={\bm{w}}_{v_{11}}{\bm{w}}_{p_{22}}^{\rm T}-{\bm{w}}_{v_{21}}{\bm{w}}_{p_{21}}^{\rm T},\\ {\bm{W}}^{p}_{22}={\bm{w}}_{v_{12}}{\bm{w}}_{p_{22}}^{\rm T}-{\bm{w}}_{v_{22}}{\bm{w}}_{p_{21}}^{\rm T},\\ \breve{\bm{W}}^{p}_{11}={\bm{w}}_{v_{21}}{\bm{w}}_{p_{11}}^{\rm T}-{\bm{w}}_{v_{11}}{\bm{w}}_{p_{12}}^{\rm T},\\ \breve{\bm{W}}^{p}_{21}={\bm{w}}_{v_{22}}{\bm{w}}_{p_{11}}^{\rm T}-{\bm{w}}_{v_{12}}{\bm{w}}_{p_{12}}^{\rm T}.\end{cases}$ (21)
-Note that the newly derived CRLB expressions in (16) and (19) represent explicit functions of the PA vector, thereby serving as effective and tractable sensing performance metrics.
+The newly derived CRLB expressions in (16) and (19) represent explicit functions of the PA coefficient vector $\bm{\rho}$}, thereby serving as effective and tractable sensing performance metrics.
 III-B Problem Formulation
@@
 Depending on specific application requirements, the optimization objective can be either sensing or communication performance, while the requirement of the other functionality is incorporated as a constraint in the optimization model.
-Let ${\bm{\rho}}_{\min}=[\rho_{\min,1},\rho_{\min,2},\cdots,\rho_{\min,N}]^{\rm T}$ denote the minimum PA coefficient to achieve the sufficiently large SNR for tight CRLBs [34, 36], and ${\bm{\rho}}_{\max}=[\rho_{\max,1},\rho_{\max,2},\cdots,\rho_{\max,N}]^{\rm T}$ denote the maximum PA coefficient due to the power budgets of individual transmitters [38, 36].
-In this paper, we consider the problem of maximizing the communication SINR, subject to constraints on the sensing performance, i.e.,
+This paper considers the problem of maximizing the communication SINR in (7), subject to constraints on the performance of joint location and velocity estimation, i.e.,
 $\small\arg\min\limits_{{\bm{\rho}}}$ $\ -{\bm{\rho}}^{\rm T}{\bm{g}},$ (22a)
-${\rm s.t.}$ ${\bm{1}}_{N}^{\rm T}{\bm{\rho}}=1,$ (22b)
+${\rm s.t.}\$${\bm{1}}_{N}^{\rm T}{\bm{\rho}}=1,$ (22b)
 ${\bm{\rho}}_{\min}\leq{\bm{\rho}}\leq{\bm{\rho}}_{\max},$ (22c)
@@
 ${\rm tr}\{{\bm{C}}_{\rm V}\}\leq\delta_{v}^{2},$ (22e)
-where $\delta_{l}^{2}$ and $\delta_{v}^{2}$ denote the corresponding location and velocity CRLB thresholds.
+where $\delta_{l}^{2}$ and $\delta_{v}^{2}$ denote the corresponding location and velocity CRLB thresholds  [^4] The thresholds can be specified by referring to the CRLB values under uniform PA..
+Note that ${\bm{\rho}}_{\min}=[\rho_{\min,1},\rho_{\min,2},\ldots,\rho_{\min,N}]^{\mathrm{T}}$ denote the minimum PA coefficient vector required to ensure a sufficiently high SNR for achieving tight CRLBs [1, 2], and ${\bm{\rho}}_{\max}=[\rho_{\max,1},\rho_{\max,2},\ldots,\rho_{\max,N}]^{\mathrm{T}}$ denote the maximum PA coefficient vector determined by the individual transmitters’ power budgets [17, 2].
+Specifically, (22b) and (22c) represent the total power and per-transmitter power constraints, respectively, while (22d) and (22e) correspond to the location and velocity CRLB constraints, respectively.
 IV Power Allocation Algorithm for CF MIMO-ISAC
-In this section, we investigate the optimal PA strategy for maximizing the ISAC performance, as formulated in (22a)-(22e).
-Unfortunately, the strong nonlinearity and non-convexity of the CRLB-based constraints in (22d) and (22e) render the problem analytically intractable.
-To overcome this challenge, we adopt a CG iterative algorithm, enhanced with a penalty function and projection operator, to provide an efficient numerical solution.
-To derive a numerical solution, we start from converting the original constrained optimization problem into an unconstrained one.
-To this end, we first introduce a penalty (function) for violating the nonlinear inequality constraints in (22c)-(22e), defined as [39]
+We now investigate the PA strategy by solving the optimization model formulated in (22a)–(22e).
+The strong nonlinearity and nonconvexity of the CRLB-based constraints in (22d) and (22e) make the problem analytically intractable.
+Although numerous numerical optimization methods developed in the literature in past several decades, these methods cannot be directly applied to the nonlinear and nonconvex constrained PA optimization problem considered in this work.
+To address this challenge, we develop a penalty function and projection operator enhanced conjugate gradient (CG)–based iterative algorithm, to yield an efficient numerical solution.
+IV-A Unconstrained Optimization Enabled by Penalty Function and Projection Operator
+To facilitate optimization, we start from converting the original constrained problem into an unconstrained one.
+Firstly, we consider incorporating the nonlinear inequality constraints in (22c)-(22e) onto the target function (22a), by introducing a penalty (function) for violating these constraints.
+It can be defined as [32]
 $\alpha({\bm{\rho}})=\tilde{f}_{l}({\bm{\rho}})+\tilde{f}_{v}({\bm{\rho}})+\sum\nolimits_{n=1}^{N}\left(f_{\min,n}({\bm{\rho}})+f_{\max,n}({\bm{\rho}})\right),$
-where the involved penalty sub-functions are given respectively by
+where the involved penalty sub-functions are
 $\small\begin{cases}\tilde{f}_{l}({\bm{\rho}})=(\max\{0,{f}_{l}({\bm{\rho}})\})^{q},\\ \tilde{f}_{v}({\bm{\rho}})=(\max\{0,{f}_{v}({\bm{\rho}})\})^{q},\\ f_{\min,n}({\bm{\rho}})=(\max\{0,\rho_{\min,n}-\rho_{n}\})^{q},\\ f_{\max,n}({\bm{\rho}})=(\max\{0,\rho_{n}-\rho_{\max,n}\})^{q},\end{cases}$ (23)
 with ${f}_{l}({\bm{\rho}})\triangleq{\rm tr}\{{\bm{C}}_{\rm L}\}-\delta_{l}^{2}$ and ${f}_{v}({\bm{\rho}})\triangleq{\rm tr}\{{\bm{C}}_{\rm V}\}-\delta_{v}^{2}$.
-$q$ is the power of the penalty function.
-We consider $q=2$ to guarantee their continuous differentiability.
-Then, the objective function for the penalty problem can be written as
+$q$ denotes the power to the penalty function, which is set to $q=2$ in this work to ensure continuous differentiability.
+Then, the penalty-driven objective function is given by
 $\small L({\bm{\rho}})=-{\bm{\rho}}^{\rm T}{\bm{g}}+\mu\alpha({\bm{\rho}}),$ (24)
 where $\mu$ denotes the penalty factor.
-Additionally, to enforce this equality constraint (22b) and reduce both computational complexity and the nonlinearity of the optimization problem, we adopt a projection operator instead of incorporating another penalty sub-function, i.e.,
+Additionally, to reduce both computational complexity and the nonlinearity of the optimization problem, we adopt a projection operator to enforce the equality constraint (22b) instead of incorporating another penalty sub-function, i.e.,
 $\small{\bm{\rho}}({\bm{\varpi}})={\bm{\Theta}}{\bm{\varpi}}+{\bm{\rho}}_{0},$ (25)
-where ${\bm{\varpi}}\in\mathbb{R}^{N\times 1}$ is an arbitrary vector and ${\bm{\Theta}}\triangleq{\bm{I}}_{N\times N}-\frac{{\bm{1}}_{N}{\bm{1}}_{N}^{\rm T}}{N}$ is a projection matrix onto the zero space of ${\bm{1}}_{N}$.
-We have ${\bm{1}}_{N}^{\rm T}{\bm{\Theta}}={\bm{0}}_{N}^{\rm T}$.
-Thus, the equality constraint ${\bm{1}}_{N}^{\rm T}{\bm{\rho}}({\bm{\varpi}})=1$ in (22b) always holds by initializing ${\bm{\rho}}_{0}$ such that ${\bm{1}}_{N}^{\rm T}{\bm{\rho}}_{0}=1$.
-Consequently, ${\bm{\rho}}({\bm{\varpi}})$ is the sum of a constant PA vector ${\bm{\rho}}_{0}$ and a projection of ${\bm{\varpi}}$ onto the zero space of ${\bm{1}}_{N}$.
-By substituting (25) into (24), we can convert the original constrained optimization to an unconstrained optimization problem as follows:
+where ${\bm{\varpi}}\in\mathbb{R}^{N\times 1}$ is an arbitrary vector and ${\bm{\Theta}}\triangleq{\bm{I}}_{N\times N}-\frac{{\bm{1}}_{N}{\bm{1}}_{N}^{\rm T}}{N}$ is a orthogonal projection matrix onto the null space of ${\bm{1}}_{N}^{\rm T}$, i.e., ${\bm{1}}_{N}^{\rm T}{\bm{\Theta}}={\bm{0}}_{N}^{\rm T}$.
+Thus, the equality constraint ${\bm{1}}_{N}^{\rm T}{\bm{\rho}}({\bm{\varpi}})=1$ in (22b) always holds provided that ${\bm{1}}_{N}^{\rm T}{\bm{\rho}}_{0}=1$.
+Consequently, ${\bm{\rho}}({\bm{\varpi}})$ is the sum of a constant PA vector ${\bm{\rho}}_{0}$ and an orthogonal projection of $\bm{\varpi}$} onto the null space of ${\bm{1}}_{N}^{\rm T}$.
+Substituting (25) into (24) can convert the original constrained optimization to an unconstrained problem as follows:
 $\small\arg\min\limits_{{\bm{\rho}}}$ $\ \mathcal{L}({\bm{\varpi}})=-{\bm{\rho}}^{\rm T}({\bm{\varpi}}){\bm{g}}+\mu\alpha({\bm{\rho}}({\bm{\varpi}})).$ (26)
-Then, unconstrained optimization methods, e.g., CG and SD methods, can be employed to search for the optimal solution.
-IV-A Modified Conjugate Gradient Algorithm
-To numerically solve (26), we propose a projection-based MCG algorithm. The update at the $i$th iteration is given by
+Building upon the formulated problem, the next section presents unconstrained optimization methods for efficiently finding the PA solution.
+IV-B Modified Conjugate Gradient Algorithm
+To numerically solve (26), we employ an iterative search algorithm. The $(i+1)$th update is given by
 $\small{\bm{\varpi}}_{i+1}={\bm{\varpi}}_{i}+\upsilon_{i}{\bm{d}}_{i},$ (27)
-where ${\bm{d}}_{i}$ is the search direction of $\mathcal{L}({\bm{\varpi}}_{i})$ and $\upsilon_{i}$ is the step size. The search direction using the MCG is updated by
+where ${\bm{d}}_{i}$ is the search direction of ${\bm{\varpi}}_{i}$ and $\upsilon_{i}$ is the step size.
+For second-order convergence, we consider using a CG algorithm, with the search direction updated by
 ${\bm{d}}_{i+1}$ $=-\nabla\mathcal{L}({\bm{\varpi}}_{i+1})+\varsigma_{i}{\bm{d}}_{i},$ (28)
-where $\varsigma_{i}$ denotes the deflection factor of a CG algorithm. Taking the derivative $\mathcal{L}({\bm{\varpi}})$ regarding ${\bm{\varpi}}$ yields
-(ϖ)=bold-italic-$\small\nabla\mathcal{L}({\bm{\varpi}})=$ ${\bm{\Theta}}\nabla L({\bm{\rho}})={\bm{\Theta}}\big{[}{\bm{g}}+\nabla\alpha({\bm{\rho}})\big{]},$ (29)
+where $\varsigma_{i}$ denotes the deflection factor of a CG algorithm. $\nabla\mathcal{L}({\bm{\varpi}}_{i})$ is the derivative $\mathcal{L}({\bm{\varpi}})$ regarding $\bm{\varpi}$}, given by
+$\small\nabla\mathcal{L}({\bm{\varpi}})=\$${\bm{\Theta}}\nabla L({\bm{\rho}})={\bm{\Theta}}\big[{\bm{g}}+\mu\nabla\alpha({\bm{\rho}})\big],$ (29)
 where
-$\small\nabla\alpha({\bm{\rho}})=$ $\mu\big{(}\nabla\tilde{f}_{l}({\bm{\rho}})+\nabla\tilde{f}_{v}({\bm{\rho}})$
+$\small\nabla\alpha({\bm{\rho}})=\$$\mu\big(\nabla\tilde{f}_{l}({\bm{\rho}})+\nabla\tilde{f}_{v}({\bm{\rho}})$
 $+\sum\nolimits_{n=1}^{N}(\nabla f_{\min,n}({\bm{\rho}})+\nabla f_{\max,n}({\bm{\rho}}))\big),$ (30)
@@
 Note that ${\bm{e}}_{n}\in\mathbb{R}^{N\times 1}$ is an index vector with its element at index $n$ being $1$ and the others being zeros.
-Given the initial direction ${\bm{d}}_{1}=-{\bm{\Theta}}\nabla L({\bm{\rho}}_{1})$, and using the idempotent property of the projection matrix, i.e., ${\bm{\Theta}}^{2}={\bm{\Theta}}$, we have ${\bm{\Theta}}{\bm{d}}_{i}={\bm{d}}_{i}$, $\forall{i>0}$.
-Thus, taking the projection of (27) leads to
+Different from classic CG algorithm, (29) is a projected gradient direction regarding $\bm{\rho}$}.
+In the following, we will directly solve PA vector $\bm{\rho}$} instead of updating $\bm{\varpi}$}.
+By using the idempotent property of the projection matrix, i.e., ${\bm{\Theta}}^{2}={\bm{\Theta}}$, we firstly have ${\bm{\Theta}}\nabla\mathcal{L}({\bm{\varpi}})=\nabla\mathcal{L}({\bm{\varpi}})$ for the gradient $\nabla\mathcal{L}({\bm{\varpi}})$ in (29).
+Then, by setting the initial direction ${\bm{d}}_{0}=-{\bm{\Theta}}\nabla L({\bm{\rho}}_{0})$ in (28), we further have ${\bm{\Theta}}{\bm{d}}_{i}={\bm{d}}_{i}$, $\forall{i>0}$.
+Finally, taking the projection of (27) leads to
 $\small{\bm{\rho}}_{i+1}={\bm{\rho}}_{i}+\upsilon_{i}{\bm{d}}_{i},$ (32)
@@
 $\small{\bm{d}}_{i+1}$ $=-{\bm{\Theta}}\nabla\mathcal{L}({\bm{\rho}}_{i+1})+\varsigma_{i}{\bm{d}}_{i}.$ (33)
-We have ${\bm{1}}_{N}^{\rm T}{\bm{d}}_{i}=0$ for $\forall{i>1}$ due to ${\bm{1}}_{N}^{\rm T}{\bm{d}}_{1}=0$ and ${\bm{1}}_{N}^{\rm T}{\bm{\Theta}}={\bm{0}}_{N}^{\rm T}$.
-Therefore, the equality constraint in (22b), that is, ${\bm{1}}_{N}^{\rm T}{\bm{\rho}}_{i}=1$, always holds at $i$th iteration for $\forall{i>1}$ if and only if the initial PA vector ${\bm{\rho}}_{1}$ satisfies ${\bm{1}}_{N}^{\rm T}{\bm{\rho}}_{1}=1$.
-We now investigate the selection of the deflection factor $\varsigma_{i}$ and step size $\upsilon_{i}$ at each iteration, which is significant in guaranteeing the convergence of the MCG algorithm.
-One necessary condition for the convergence is the non-increasing objective function at each iteration, i.e.,
+We refer to (32) and (33) as the penalty-function and projection-based modified CG method (PP-MCG).
+Different from (27), (32) yields the numerically iterative solution of the penalty-driven objective function (24).
+Additionally, with the search direction in (33), the equality constraint (22b), that is, ${\bm{1}}_{N}^{\rm T}{\bm{\rho}}_{i}=1$, always holds at the $i$th iteration for $\forall{i>0}$ if and only if the initial PA vector ${\bm{\rho}}_{0}$ satisfies ${\bm{1}}_{N}^{\rm T}{\bm{\rho}}_{0}=1$.
+This is because ${\bm{1}}_{N}^{\rm T}{\bm{d}}_{i}=0$ for $\forall{i>0}$ provided that ${\bm{1}}_{N}^{\rm T}{\bm{d}}_{0}=0$, due to ${\bm{1}}_{N}^{\rm T}{\bm{\Theta}}={\bm{0}}_{N}^{\rm T}$.
+IV-C Parameter Optimization of PP-MCG Algorithm
+A necessary condition for the convergence of the PP-MCG algorithm is that the objective function is non-increasing at each iteration, i.e.,
 $\small L({\bm{\rho}}_{i}+\upsilon_{i}{\bm{d}}_{i})-L({\bm{\rho}}_{i})\leq 0.$ (34)
-To this end, the search direction ${\bm{d}}_{i}$ should be the descent direction and the step size $\upsilon_{i}$ should be correspondingly upper bounded by a threshold at each iteration.
-IV-A1 Descent direction
-To ensure a descent direction ${\bm{d}}_{i+1}$ at the $(i+1)$th iteration, it requires
+To this end, ${\bm{d}}_{i}$ should be a descent direction, and the step size $\upsilon_{i}$ should be optimized at each iteration.
+To ensure the descent direction, the deflection factor $\varsigma_{i}$ is crucial.
+IV-C1 Descent direction
+To ensure that ${\bm{d}}_{i+1}$ is a descent direction of $L({\bm{\rho}}_{i+1})$, it is required that
 $\small{\bm{d}}_{i+1}^{\rm T}\nabla L({\bm{\rho}}_{i+1})\leq 0,$ (35)
 where the equality holds if and only if $\nabla L({\bm{\rho}}_{i+1})={\bm{0}}$.
-The initial direction ${\bm{d}}_{1}=-{\bm{\Theta}}\nabla L({\bm{\rho}}_{1})$ is the descent direction of $L({\bm{\rho}}_{1})$ with $-\nabla L({\bm{\rho}}_{1})^{\rm T}{\bm{\Theta}}\nabla L({\bm{\rho}}_{1})\leq 0$ since the projection operator ${\bm{\Theta}}$ is a symmetric positive definite matrix.
-Substituting (33) into (35), we have
+The initial direction ${\bm{d}}_{0}=-{\bm{\Theta}}\nabla L({\bm{\rho}}_{0})$ is descent with ${\bm{d}}_{0}^{\rm T}\nabla L({\bm{\rho}}_{0})\leq 0$ since the projection operator $\bm{\Theta}$} is a symmetric positive definite matrix.
+By substituting (33) into (35), the feasible range of the deflection factor can be obtained, i.e.,
 $\small\begin{cases}\varsigma_{i}\leq\varsigma_{i}^{\rm b},\ \varsigma_{i}^{\rm b}>0\\ \varsigma_{i}\geq\varsigma_{i}^{\rm b},\ \varsigma_{i}^{\rm b}<0\end{cases}$ (36)
-where $\varsigma_{i}^{\rm b}$ is the bound for $\varsigma_{i}$ with its expression given by
+where $\varsigma_{i}^{\rm b}$ denotes the bound in selecting $\varsigma_{i}$, given by
 $\small\varsigma_{i}^{\rm b}=\dfrac{\nabla L({\bm{\rho}}_{i+1})^{\rm T}{\bm{\Theta}}\nabla L({\bm{\rho}}_{i+1})}{{\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i+1})}.$ (37)
-The numerator satisfies $\nabla L({\bm{\rho}}_{i+1})^{\rm T}{\bm{\Theta}}\nabla L({\bm{\rho}}_{i+1})>0$ for any non-zero gradient $\nabla L({\bm{\rho}}_{i+1})$ due to the positive definite ${\bm{\Theta}}$.
-When employing inexact line searches, the deflection selection proposed by Hestenes and Stiefel [40], referred to as HS method, for $\varsigma_{i}$ is preferable.
-Accordingly, we adopt the modified HS deflection, given by
-$\small\varsigma_{i}^{\rm HS}=\dfrac{\nabla L({\bm{\rho}}_{i+1})^{\rm T}{\bm{q}}_{i}}{{\bm{d}}_{i}^{\rm T}{\bm{q}}_{i}},$ (38)
+Note that $\nabla L({\bm{\rho}}_{i+1})^{\rm T}{\bm{\Theta}}\nabla L({\bm{\rho}}_{i+1})>0$ for any nonzero gradient $\nabla L({\bm{\rho}}_{i+1})$, owing to the positive definiteness of $\bm{\Theta}$}.
+When employing inexact line searches, the deflection selection for $\varsigma_{i}$ proposed by Hestenes and Stiefel [21], referred to as HS method, is preferable.
+Herein, we modify the HS deflection for the projection-driven iterative method, i.e.,
+$\small\varsigma_{i}^{\rm HS}={\nabla L({\bm{\rho}}_{i+1})^{\rm T}{\bm{q}}_{i}}/{{\bm{d}}_{i}^{\rm T}{\bm{q}}_{i}},$ (38)
 which is obtained from ${\bm{d}}_{i+1}^{\rm T}{\bm{q}}_{i}=0$ with ${\bm{q}}_{i}\triangleq{\bm{\Theta}}[\nabla L({\bm{\rho}}_{i+1})-\nabla L({\bm{\rho}}_{i})]$.
@@
 $|\nabla L({\bm{\rho}}_{i+1})^{\rm T}{\bm{\Theta}}\nabla L({\bm{\rho}}_{i})|\ll|\nabla L({\bm{\rho}}_{i+1})^{\rm T}{\bm{\Theta}}\nabla L({\bm{\rho}}_{i+1})|.$ (40)
-(39) and (40) ensure the positivity of ${\bm{d}}_{i}^{\rm T}{\bm{q}}i$ and $\nabla L({\bm{\rho}}{i+1})^{\rm T}{\bm{q}}_{i}$, respectively.
-This allows for a positive conjugate gradient scaling factor $\varsigma_{i}^{\rm HS}$.
+Note that (39) and (40) ensure the positivity of ${\bm{d}}_{i}^{\rm T}{\bm{q}}_{i}$ and $\nabla L({\bm{\rho}}_{i+1})^{\rm T}{\bm{q}}_{i}$, respectively.
+This allows for a positive conjugate gradient deflection factor $\varsigma_{i}^{\rm HS}$.
 Furthermore, (39) and (40) also imply that $|{\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i+1})|\ll|{\bm{d}}_{i}^{\rm T}{\bm{q}}_{i}|$ and $|\nabla L({\bm{\rho}}_{i+1})^{\rm T}{\bm{q}}_{i}|\approx|\nabla L({\bm{\rho}}_{i+1})^{\rm T}{\bm{\Theta}}\nabla L({\bm{\rho}}_{i+1})|$, respectively.
 As a result, $\varsigma_{i}^{\rm HS}$ is expected to be smaller than $|\varsigma_{i}^{\rm b}|$, thereby satisfying the deflection constraint defined in (36).
-Rather than strictly relying on the potentially tight and unquantifiable conjugacy conditions in (39) and (40), we instead enforce the positivity of both the numerator and denominator in (38), along with the inequality $\varsigma_{i}^{\rm HS}<|\varsigma_{i}^{\rm b}|$.
+Rather than strictly relying on the potentially tight and unquantifiable conjugacy conditions in (39) and (40), we instead directly enforce the positivity of both the numerator and denominator in (38), along with the inequality $\varsigma_{i}^{\rm HS}<|\varsigma_{i}^{\rm b}|$.
 When these conditions hold, we set $\varsigma_{i}=\varsigma_{i}^{\rm HS}$.
@@
 To sum up, the deflection factor is updated by
-${\varsigma_{i}=}\begin{cases}\varsigma_{i}^{\rm HS},\ \min\{\nabla L({\bm{\rho}}_{i+1})^{\rm T}{\bm{q}}_{i},{\bm{d}}_{i}^{\rm T}{\bm{q}}_{i},|\varsigma_{i}^{\rm b}|-\varsigma_{i}^{\rm HS}\}>0,\\ 0,\ {\rm else}.\end{cases}$ (41)
-Note that, after every consecutive $N$ MCG steps with $\varsigma_{i}=\varsigma_{i}^{\rm HS}$, we reset $\varsigma_{i}=0$ to initiate a new cycle of MCG iterations.
+$\small{\varsigma_{i}=}\begin{cases}\varsigma_{i}^{\rm HS},\ \min\{\nabla L({\bm{\rho}}_{i+1})^{\rm T}{\bm{q}}_{i},{\bm{d}}_{i}^{\rm T}{\bm{q}}_{i},|\varsigma_{i}^{\rm b}|-\varsigma_{i}^{\rm HS}\}>0,\\ 0,\ {\rm else}.\end{cases}$ (41)
+Note that, after every consecutive $Q$ MCG steps with $\varsigma_{i}=\varsigma_{i}^{\rm HS}$, we reset $\varsigma_{i}=0$ to initiate a new cycle of MCG iterations.
 Such a reboot strategy can help mitigate the accumulation of numerical errors, restoring the orthogonality and conjugacy of direction vectors, and avoiding local minima in nonlinear and non-convex optimization.
-IV-A2 Step size optimization
-To optimize the step size selection, we define $f_{i}(\upsilon_{i})\triangleq L({\bm{\rho}}_{i}+\upsilon_{i}{\bm{d}}_{i})$ with $f_{i}(0)=L({\bm{\rho}}_{i})$.
+IV-C2 Step size optimization
+We have explored the selection of deflection factor above to ensure a descent direction in (35). The second step is to optimize the step size selection.
+Let us define $f_{i}(\upsilon_{i})\triangleq L({\bm{\rho}}_{i}+\upsilon_{i}{\bm{d}}_{i})$ with $f_{i}(0)=L({\bm{\rho}}_{i})$.
 The second-order Taylor approximation of $f_{i}(\upsilon_{i})$ at $\upsilon_{i}=0$ is given by
@@
 where $\nabla^{2}L({\bm{\rho}}_{i})$ is the Hessian matrix of $L({\bm{\rho}})$ at ${\bm{\rho}}_{i}$.
-According to (35), we have ${\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i})\leq 0$.
-If $\|\nabla^{2}L({\bm{\rho}}_{i})\|$ is locally bounded, there exists a non-negative step size bound $\upsilon_{b}$ for given $\epsilon_{L}\in[0,1)$ such that
-$\small f_{i}(\upsilon_{i})-f_{i}(0)\leq\epsilon_{L}\upsilon_{i}{\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i}),\quad\text{for}\quad\upsilon_{i}\leq\upsilon_{b}.$ (43)
-(43) always holds when ${\bm{d}}_{i}^{\rm T}\nabla^{2}L({\bm{\rho}}_{i}){\bm{d}}_{i}\leq 0$.
-In contrast, we have $\upsilon_{b}=\dfrac{2(\epsilon_{L}-1){\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i})}{{\bm{d}}_{i}^{\rm T}\nabla^{2}L({\bm{\rho}}_{i}){\bm{d}}_{i}}$ when ${\bm{d}}_{i}^{\rm T}\nabla^{2}L({\bm{\rho}}_{i}){\bm{d}}_{i}>0$.
-Due to the inherent high computational complexity of the Hessian matrix $\nabla^{2}L({\bm{\rho}}_{i})$ in calaulating $\upsilon_{b}$, we can apply an inexact line search (ILS) for the step size to enable
+Using the optimized deflection factor above, we have ${\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i})\leq 0$.
+If $\|\nabla^{2}L({\bm{\rho}}_{i})\|$ is locally bounded, there will exist a non-negative step size bound $\upsilon_{i,\rm b}$ for a given constant $\epsilon_{L}\in[0,1)$ such that
+$\small f_{i}(\upsilon_{i})-f_{i}(0)\leq\epsilon_{L}\upsilon_{i}{\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i}),\quad\text{for}\quad\upsilon_{i}\leq\upsilon_{i,\rm b}.$ (43)
+When ${\bm{d}}_{i}^{\rm T}\nabla^{2}L({\bm{\rho}}_{i}){\bm{d}}_{i}\leq 0$, (43) always holds.
+Otherwise, we have $\upsilon_{i,\rm b}=\dfrac{2(\epsilon_{L}-1){\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i})}{{\bm{d}}_{i}^{\rm T}\nabla^{2}L({\bm{\rho}}_{i}){\bm{d}}_{i}}$ when ${\bm{d}}_{i}^{\rm T}\nabla^{2}L({\bm{\rho}}_{i}){\bm{d}}_{i}>0$.
+One can choose a step size at each iteration as $\upsilon_{i}=\delta_{i}\upsilon_{i,\rm b}$ for the latter case, with a small constant $\delta_{i}\in(0,1)$.
+However, in this work, due to the high computational complexity of the Hessian matrix $\nabla^{2}L({\bm{\rho}}_{i})$ when computing $\upsilon_{i,\mathrm{b}}$, we adopt an inexact line search (ILS) strategy for step-size selection to satisfy (43), i.e.,
 $\small L({\bm{\rho}}_{i}+\upsilon_{i}{\bm{d}}_{i})-L({\bm{\rho}}_{i})\leq\epsilon_{L}\upsilon_{i}{\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i}).$ (44)
-According to (35), the term on the right-hand side of (44) is always non-positive.
-Thus, (44) can enable a descent function in (34).
+Note that (44) can ensure a descent function in (34) since the term on the right-hand side is always non-positive.
 The ILS controlling factor $\epsilon_{L}$ is used to regulate the decrement of the function $L({\bm{\rho}})$ by moving ${\bm{\rho}}_{i}$ towards the direction ${\bm{d}}_{i}$ weighted by the step size $\upsilon_{i}$.
-When (44) is unmet for given $\upsilon_{i}$, the step size is scaled by
+When (44) is unmet for a given $\upsilon_{i}$, the step size is scaled by
 $\small\upsilon_{i}=\sigma_{\upsilon}\upsilon_{i},$ (45)
-where $\sigma_{\upsilon}\in(0,1)$ denotes the scaling factor and an initial step size is denoted as $\upsilon_{i}=\upsilon_{0,i}$.
-Additionally, the penalty function does not inherently ensure that the individual power budget constraints at each transmitter are respected during the optimization process.
-As a result, negative PA values may arise in intermediate iterations.
-These infeasible PA values can lead to misleading CRLB evaluations, for instance, producing negative CRLBs instead of infinite ones, thereby falsely satisfying the CRLB constraints.
-To avoid unnecessary iterations within the infeasible region and to ensure physically meaningful results, it is crucial to enforce non-negative PAs at each iteration step.
+where $\sigma_{\upsilon}\in(0,1)$ is a scaling factor. An initial step size, denoted as $\upsilon_{i}=\upsilon_{i,0}$, can be chosen as described later.
+Additionally, negative PA values may occur in intermediate iterations.
+Such infeasible values can lead to misleading CRLB evaluations, e.g., producing negative rather than infinite CRLBs, and may thus falsely appear to satisfy the CRLB constraints.
+To prevent unnecessary iterations within the infeasible region and to ensure physically meaningful results, it is essential to enforce non-negative PAs at each iteration step.
 To this end, the step size $\upsilon_{i}$ needs to satisfy
@@
 $\upsilon_{i}^{\rm b}=\min_{d_{n,i}<0}\{\rho_{n,i}/|{d}_{n,i}|\},$ (47)
-with $\rho_{n,i}$ denoting the $n$th element of ${\bm{\rho}}_{i}$ and similar for ${d}_{n,i}$. Thus, the initial step size for ILS can be selected as
-$\small\upsilon_{0,i}=\epsilon_{\upsilon}\upsilon_{i}^{\rm b}$ (48)
+with $\rho_{n,i}$ denoting the $n$th element of ${\bm{\rho}}_{i}$ and similar for ${d}_{n,i}$.
+Note that there must exist at least one component ${d}_{n,i}<0$ in each iteration $i$, unless ${\bm{d}}_{i}={\bm{0}}$, due to the constraint ${\bm{1}}_{N}^{\rm T}{\bm{d}}_{i}=0$.
+Thus, the initial step size for ILS can be selected as
+$\small\upsilon_{i,0}=\epsilon_{\upsilon}\upsilon_{i}^{\rm b}$ (48)
 with $\epsilon_{\upsilon}\in(0,1]$.
-Note that there must exist at least one component ${d}_{n,i}<0$ in each iteration $i$, unless ${\bm{d}}_{i}={\bm{0}}$, due to the constraint $\sum_{n=1}^{N}{d}_{n,i}=0$.
-The ILS method characterized in (44)-(48) can be seen as the extension and application of the Armijo’s Rule [41] in our work.
-IV-A3 The algorithm implementation
+The ILS method, characterized by (44)–(48), can be regarded as an adaptation of Armijo’s rule [7], customized to handle the non-negative PA optimization under projection constraint in our proposed framework.
+IV-C3 The algorithm implementation
 A well-defined termination condition is essential for the proposed iterative algorithm.
-A natural choice, inspired by the ILS condition in (44), is to monitor the absolute change in the objective function between consecutive iterations, i.e., $|L({\bm{\rho}}_{i+1})-L({\bm{\rho}}_{i})|\leq\epsilon$, where $\epsilon$ is a predefined threshold with small value.
+A natural choice, inspired by the ILS condition in (44), is to monitor the absolute change in the objective function between consecutive iterations, i.e., $|L({\bm{\rho}}_{i+1})-L({\bm{\rho}}_{i})|\leq\epsilon$, where $\epsilon$ is a predefined small threshold.
 To avoid numerical instability arising from the magnitude of the objective function itself, we apply a normalized termination condition, i.e., $|L({\bm{\rho}}_{i+1})-L({\bm{\rho}}_{i})|/|L({\bm{\rho}}_{i+1})|\leq\epsilon_{\rm th}$.
-This condition is employed in both the MCG-based and MSD-based algorithms.
-We now present the complete PP-MCG-ILS.
+This is employed in both the MCG-based and the following MSD-based algorithms.
+Finally, the selection of penalty factor is also important in guaranteeing a good learning performance.
 Generally, a large penalty factor may lead to numerical issues, such as an ill-conditioned Hessian matrix, which can degrade the convergence behavior of the MCG algorithm.
@@
 This sequential penalty strategy ensures both numerical stability and progressive enforcement of the constraint conditions.
-The detailed procedure of the proposed PP-MCG-ILS algorithm is summarized in Algorithm 1.
+The detailed procedure of the PP-MCG-ILS algorithm is summarized in Algorithm 1.
 Algorithm 1 The PP-MCG-ILS algorithm
-Input: The positive scalar termination factor $\epsilon_{\mu}$ and $\epsilon_{\rm th}$, small positive constant $\sigma_{\upsilon}$, $\epsilon_{L}$ and $\epsilon_{\upsilon}$, initial penalty factor $\mu_{1}$, scaling factor $\varphi>1$; initial PA parameter ${\bm{\rho}}_{1}$ satisfying ${\bm{1}}^{\rm T}{\bm{\rho}}_{1}=1$, projection operator ${\bm{\Theta}}$, ${\bm{y}}_{1}={\bm{\rho}}_{1}$, and ${\bm{d}}_{1}=-{\bm{\Theta}}\nabla L({\bm{\rho}}_{1})$; a counter $k=0$ and $i=j=1$.
+Input: The positive scalar termination factor $\epsilon_{\mu}$ and $\epsilon_{\rm th}$, small positive constant $\sigma_{\upsilon}$, $\epsilon_{L}$ and $\epsilon_{\upsilon}$, initial penalty factor $\mu_{1}$, scaling factor $\varphi>1$; initial PA parameter ${\bm{\rho}}_{0}$ satisfying ${\bm{1}}^{\rm T}{\bm{\rho}}_{0}=1$, projection operator $\bm{\Theta}$}, ${\bm{y}}_{1}={\bm{\rho}}_{1}$, and ${\bm{d}}_{1}=-{\bm{\Theta}}\nabla L({\bm{\rho}}_{1})$; a counter $k=0$ and $i=j=1$.
 Output: The PA vector ${\bm{y}}_{j}$.
@@
 4:  Convergence condition: If $|L({\bm{\rho}}_{i+1})-L({\bm{\rho}}_{i})|/|L({\bm{\rho}}_{i+1})|\leq\epsilon_{\rm th}$, ${\bm{y}}_{j}={\bm{\rho}}_{i+1}$, and go to step 7; Otherwise, go to next step.
-5:  Deflection choice: According to (41), if $\varsigma_{i}=\varsigma_{i}^{\rm HS}$, $k=k+1$, or $k=0$ otherwise; when $k=N$, let $\varsigma_{i}=0$ and $k=0$;
+5:  Deflection choice: According to (41), if $\varsigma_{i}=\varsigma_{i}^{\rm HS}$, $k=k+1$, or $k=0$ otherwise; when $k=Q$, let $\varsigma_{i}=0$ and $k=0$;
 6:  Direction update: ${\bm{d}}_{i+1}=-{\bm{\Theta}}\nabla L({\bm{\rho}}_{i+1})+\varsigma_{i}{\bm{d}}_{i}$; let $i=i+1$, and go back to step 1.
 7:  Penalty update: If $\mu_{j}\alpha({\bm{y}}_{j})<\epsilon_{\mu}$, stop; Otherwise, $\mu_{j+1}=\varphi\mu_{j}$, and ${\bm{d}}_{i+1}=-{\bm{\Theta}}\nabla L({\bm{\rho}}_{i+1})$; let $j=j+1$ and $i=i+1$, and go to step 1.
-IV-B Modified Steepest Descent Algorithm
+IV-D Modified Steepest Descent Algorithm
 In fact, alternative unconstrained optimization methods can be applied to solve the proposed projection-based penalty problem in (26).
-In this paper, we also develop a MSD algorithm, adapted from the proposed MCG framework.
+In this paper, we also develop an MSD algorithm, adapted from the proposed MCG framework.
 The main difference between the MSD and MCG algorithms lies in how the search direction is determined.
-In the MSD approach, the search direction at each iteration is set to the negative gradient of the objective function, i.e.,
+In the MSD approach, the search direction is set to the projected negative gradient of the penalty-based objective function, i.e.,
 $\small{\bm{d}}_{i}$ $=-{\bm{\Theta}}\nabla L({\bm{\rho}}_{i}),$ (49)
@@
 We herein omit the detailed implementation steps of PP-MSD-ILS since they can be obtained from the PP-MCG-ILS in Algorithm 1 by simply removing step 5 and performing the direction update in step 6 using (49).
-IV-C Benchmarking Schemes
-To facilitate performance comparison, we introduce two benchmark algorithms based on the PP-MCG-ILS and PP-MSD-ILS frameworks, called penalty function and projection-driven normalized SD (PP-NSD) and penalty function and projection-driven normalized CG (PP-NCG), respectively.
-Here, we employ the standard SD and CG methods proposed in [42], where fixed step sizes are adopted instead of ILS.
-To mitigate potential instability arising from overly large or small gradient magnitudes under fixed-step updates, we normalize the update directions using their $l_{2}$ norm.
-Specifically, for the PP-NSD algorithm, the search direction is defined as an $\epsilon_{d}$-regularized normalized gradient, given by
+IV-E Benchmarking Schemes
+Here, we employ the standard SD and CG methods proposed in [37] as benchmarks, where fixed step sizes are adopted instead of ILS.
+To facilitate performance comparison, these two benchmark algorithms should be adapted to our problem below, called penalty function and projection-driven normalized SD (PP-NSD) and penalty function and projection-driven normalized CG (PP-NCG), respectively.
+To mitigate potential instability arising from overly large or small gradient magnitudes under fixed-step updates, we normalize the update directions of CG or SD algorithms using their $l_{2}$ norm.
+Specifically, for the PP-NSD algorithm, the search direction is expressed by an $\epsilon_{d}$-regularized normalized gradient
 ${\bm{d}}_{i}$ $=-{\bm{g}}_{i}/(\|{\bm{g}}_{i}\|_{2}+\epsilon_{d}),$ (50)
-where ${\bm{g}}_{i}={\bm{\Theta}}\nabla L({\bm{\rho}}_{i})$, and $\epsilon_{d}$ is a small positive constant introduced to prevent division by zero.
-Similarly, for the PP-NCG algorithm, the search direction is given by an $\epsilon_{d}$-regularized normalized conjugate gradient, i.e.,
+where ${\bm{g}}_{i}={\bm{\Theta}}\nabla L({\bm{\rho}}_{i})$, and $\epsilon_{d}$ is a small positive constant to prevent division by zero.
+Similarly, for the PP-NCG algorithm, the search direction is given by an $\epsilon_{d}$-regularized normalized CG, i.e.,
 $\small{\bm{d}}_{i+1}$ $=-({\bm{g}}_{i+1}+\varsigma_{i}{\bm{d}}_{i})/(\|({\bm{g}}_{i+1}+\varsigma_{i}{\bm{d}}_{i})\|_{2}+\epsilon_{d}),$ (51)
@@
 with $\upsilon_{i}^{\rm b}$ given in (47) and $\upsilon_{0}$ being a predetermined step size.
-The maximum number of iterations is set to $I$.
-Due to the non-monotonic penalty function, the minimum may not occur at the final iteration but within a few steps before it.
+Due to the uncertain convergence of the benchmarks, their maximum number of iterations is set to $I$.
+Moreover, due to the non-monotonic penalty function, the minimum may not occur at the final iteration but within a few steps before it.
 To address this, we introduce an alternative penalty stopping criterion given by $\mu_{j}\min\limits_{i+1-S\leq q\leq i+1}\{\alpha({\bm{\rho}}_{q})\}<\epsilon_{\mu}$ where $S$ represents the window size for searching the minimum penalty value.
 The detailed implementation steps of the PP-NSD algorithm are omitted here, as it can be easily constructed from the PP-NCG described in Algorithm 2 by removing step 4 and performing the direction update in step 5 using (50).
-Algorithm 2 The PP-NCG algorithm
-Input: Positive scalar termination factor $\epsilon_{\mu}$ and $\epsilon_{\rm th}$, a large integer $I$ and a small integer $S$, small positive constants $\sigma_{\upsilon}$, $\epsilon_{L}$, $\epsilon_{d}$ and $\epsilon_{\upsilon}$, initial step size $\upsilon_{0}$, initial penalty factor $\mu_{1}$ and its scaling factor $\varphi>1$; initial PA parameter ${\bm{\rho}}_{1}$ satisfying ${\bm{1}}^{\rm T}{\bm{\rho}}_{1}=1$, projection operator ${\bm{\Theta}}$, ${\bm{y}}_{1}={\bm{\rho}}_{1}$, and ${\bm{d}}_{1}=-{\bm{\Theta}}\nabla L({\bm{\rho}}_{1})/(\|{\bm{\Theta}}\nabla L({\bm{\rho}}_{1})\|_{2}+\epsilon_{d})$; a counter $k=0$ and $i=j=1$.
+Algorithm 2 The $P$-NCG algorithm
+Input: Positive scalar termination factor $\epsilon_{\mu}$ and $\epsilon_{\rm th}$, a large integer $I$ and a small integer $S$, small positive constants $\sigma_{\upsilon}$, $\epsilon_{L}$, $\epsilon_{d}$ and $\epsilon_{\upsilon}$, initial step size $\upsilon_{0}$, initial penalty factor $\mu_{1}$ and its scaling factor $\varphi>1$; initial PA parameter ${\bm{\rho}}_{0}$ satisfying ${\bm{1}}^{\rm T}{\bm{\rho}}_{0}=1$, projection operator $\bm{\Theta}$}, ${\bm{y}}_{1}={\bm{\rho}}_{1}$, and ${\bm{d}}_{1}=-{\bm{\Theta}}\nabla L({\bm{\rho}}_{1})/(\|{\bm{\Theta}}\nabla L({\bm{\rho}}_{1})\|_{2}+\epsilon_{d})$; a counter $k=0$ and $i=j=1$.
 Output: The PA vector ${\bm{y}}_{j}$.
@@
 3:  Convergence condition: If $|L({\bm{\rho}}_{i+1})-L({\bm{\rho}}_{i})|/|L({\bm{\rho}}_{i+1})|\leq\epsilon_{\rm th}$ or $i+1=I$, ${\bm{y}}_{j}={\bm{\rho}}_{i+1}$, go to step 6; Otherwise, go to next step.
-4:  Deflection choice: According to (41), if $\varsigma_{i}=\varsigma_{i}^{\rm HS}$, $k=k+1$, or $k=0$ otherwise; when $k=N$, let $\varsigma_{i}=0$ and $k=0$;
+4:  Deflection choice: According to (41), if $\varsigma_{i}=\varsigma_{i}^{\rm HS}$, $k=k+1$, or $k=0$ otherwise; when $k=Q$, let $\varsigma_{i}=0$ and $k=0$;
 5:  Direction update: ${\bm{d}}_{i+1}$ through (51); let $i=i+1$, and go back to step 1.
@@
 In this section, we consider a pure sensing scenario with joint location and velocity estimation.
-The P-NCG-ILS algorithm is proposed for PA in minimizing the total power under power budget constraints and CRLB constraints.
-The constrained power minimization problem is formulated as
+The P-NCG-ILS algorithm is proposed for PA optimization in minimizing the total power under power budget constraints and CRLB constraints.
+The constrained power minimization problem is given by
 $\small\arg\min\limits_{{\bm{\rho}}}$ $\ {\bm{\rho}}^{\rm T}{\bm{1}},$ (53a)
-${\rm s.t.}$ ${\bm{\rho}}^{\rm T}{\bm{1}}\leq 1,$ (53b)
+${\rm s.t.}\$${\bm{\rho}}^{\rm T}{\bm{1}}\leq 1,$ (53b)
 […]
 Analogously to the derivations of PP-MCG-ILS, we set the penalty function for violating the inequality constraints as
-$\small\alpha_{\rm s}({\bm{\rho}})=$ ${f}_{\rm s}({\bm{\rho}})+\tilde{f}_{l}({\bm{\rho}})+\tilde{f}_{v}({\bm{\rho}})$
-$+\sum_{n=1}^{N}\left(f_{\min,n}({\bm{\rho}})+f_{\max,n}({\bm{\rho}})\right),$ (54)
+$\small\alpha_{\rm s}({\bm{\rho}})=\$${f}_{\rm s}({\bm{\rho}})+\tilde{f}_{l}({\bm{\rho}})+\tilde{f}_{v}({\bm{\rho}})$
+$+\sum\nolimits_{n=1}^{N}\left(f_{\min,n}({\bm{\rho}})+f_{\max,n}({\bm{\rho}})\right),$ (54)
 where ${f}_{\rm s}({\bm{\rho}})$ denotes the penalty sub-function for the total power budget, i.e.,
@@
 4:  Convergence condition: If $|L_{\rm s}({\bm{\rho}}_{i+1})-L_{\rm s}({\bm{\rho}}_{i})|/|L_{\rm s}({\bm{\rho}}_{i+1})|\leq\epsilon_{\rm th}$, ${\bm{y}}_{j}={\bm{\rho}}_{i+1}$, and go to step 5; Otherwise, go to next step.
-5:  Deflection choice: According to (41), if $\varsigma_{i}=\varsigma_{i}^{\rm HS}$, $k=k+1$, or $k=0$ otherwise; when $k=N$, let $\varsigma_{i}=0$ and $k=0$.
+5:  Deflection choice: According to (41), if $\varsigma_{i}=\varsigma_{i}^{\rm HS}$, $k=k+1$, or $k=0$ otherwise; when $k=Q$, let $\varsigma_{i}=0$ and $k=0$.
 6:  Direction update: ${\bm{d}}_{i+1}=\tilde{\bm{d}}_{i+1}/(\|\tilde{\bm{d}}_{i+1}\|_{2}+\epsilon_{d})$ with $\tilde{\bm{d}}_{i+1}=-\nabla L_{\rm s}({\bm{\rho}}_{i+1})+\varsigma_{i}{\bm{d}}_{i}$; let $i=i+1$, and go back to step 1.
@@
 VI Algorithm Analysis
-The proposed algorithms (e.g., PP-MCG-ILS, PP-MSD-ILS) can be extended and applied to a broad class of nonlinear and non-convex optimization problems.
-These problems are typically characterized by an affine equality constraint, if present, satisfied through a projection operator, along with nonlinear or non-convex constraints incorporated via penalty functions.
-The ILS mechanism is used to enhance convergence.
-For example, the proposed P-NCG-ILS algorithm for sensing-only scenario can be seen as a simple extension of the PP-MCG-ILS algorithm.
+The proposed algorithms (e.g., PP-MCG-ILS and PP-MSD-ILS) can be readily extended to a broad class of nonlinear and nonconvex optimization problems.
+Such problems typically involve affine equality constraints, which, when present, are enforced via projection operators, while nonlinear or nonconvex constraints are handled through penalty functions.
+The ILS mechanism is further employed to improve convergence stability.
+As an example, the proposed P-NCG-ILS algorithm for the sensing-only scenario can be regarded as a direct extension of the PP-MCG-ILS algorithm.
 In the following, we analyze the convergence behavior and computational complexity of the proposed algorithms.
@@
 Theorem 1.
-Given a nonempty convex set $\Omega$ with the objective function $L({\bm{\rho}}_{i})$ bounded over its interior points, if the objective function is Lipschitz gradient continuous over $\Omega$, the ILS-based iterative algorithm will converge to a stationary point with a zero descent direction ${\bm{d}}_{i}\rightarrow{\bm{0}}$.
-In contrast, if it is Lipschitz gradient non-continuous, sequences $\{{\bm{\rho}}_{i}\}$ and $\{L({\bm{\rho}}_{i})\}$ may converge to a Lipschitz gradient noncontinuous point with the norm of the Hessian matrix approaching infinity.
+Consider the CRLB-based penalty objective $L(\bm{\rho})$ defined in (24), over a nonempty and bounded convex feasible set $\Omega\subset\mathbb{R}^{n}$.
+If $L$ is continuously differentiable and admits a Lipschitz continuous gradient on $\Omega$, the proposed ILS-based schemes (PP-MCG-ILS and PP-MSD-ILS) generate a sequence $\{\bm{\rho}_{i}\}$ whose limit points are first-order stationary with a zero descent direction ${\bm{d}}_{i}\rightarrow{\bm{0}}$.
+Theorem 1 is included to guarantee the existence of a well-defined and bounded step size for the proposed ILS-based PA algorithms.
 Proof. The proof is given in Appendix C. $\blacksquare$
-Theorem 1 provides the convergence analysis for the proposed iterative algorithms.
-We now prove the Lipschitz gradient continuity of the objectivity functions in (26) and (56).
-Firstly, all CRLB-involved matrices, e.g., ${\bm{P}}(\rho)$, ${\bm{Y}}(\rho)$, and their inverses are smooth on ${\bm{\rho}}$ over the feasible domain $\Omega\subset\mathbb{R}^{N}$.
-Moreover, ${\bm{C}}_{L}({\bm{\rho}})$ is uniformly positive definite.
+We now establish the Lipschitz continuity of the gradients of the objective functions in (26) and (56) by showing that their second-order derivatives are uniformly bounded over the feasible set.
+Firstly, since $\Omega\subset\mathbb{R}^{N}$ is compact and all CRLB-related matrices, e.g., ${\bm{P}}(\rho)$, ${\bm{Y}}(\rho)$, and their inverses depend smoothly on $\bm{\rho}$}, their spectral norms and those of their derivatives are uniformly bounded over $\Omega$.
+As a result, the operator norm of the Hessian of $f_{l}({\bm{\rho}})$ is uniformly bounded.
+Moreover, we know that ${\bm{C}}_{L}({\bm{\rho}})$ is uniformly positive definite.
 It follows that ${\bm{C}}_{L}({\bm{\rho}})$ is continuously differentiable, and so is ${\rm tr}\{{\bm{C}}_{L}({\bm{\rho}})\}$.
@@
 Since the matrix inverse and trace are smooth on the set of positive definite matrices, $\nabla f_{l}({\bm{\rho}})$ is continuously differentiable.
-Moreover, all matrices involved (including ${\bm{C}}_{L}({\bm{\rho}})$) have bounded spectral norm, so the operator norm of the Hessian of $f_{l}({\bm{\rho}})$ is also bounded.
+Together with the uniform boundedness of the Hessian $\nabla^{2}f_{l}({\bm{\rho}})$ over $\Omega$, this implies that $\nabla f_{l}({\bm{\rho}})$ is Lipschitz continuous over $\Omega$.
 Consequently, $\nabla\tilde{f}_{l}({\bm{\rho}})$ is Lipschitz continuous.
 A similar argument applies to $\nabla\tilde{f}_{v}({\bm{\rho}})$.
-On the other hand, $f_{\min,n}({\bm{\rho}})$, $f_{\max,n}({\bm{\rho}})$ and ${f}_{\rm s}({\bm{\rho}})$ are piecewise quadratic, and their gradients are Lipschitz continuous.
+Furthermore, $f_{\min,n}({\bm{\rho}})$, $f_{\max,n}({\bm{\rho}})$ and ${f}_{\rm s}({\bm{\rho}})$ are piecewise quadratic with finitely many breakpoints, and their gradients are globally Lipschitz continuous over $\Omega$.
 Hence, $\nabla_{\bm{\rho}}\alpha({\bm{\rho}})$ and $\nabla_{\bm{\rho}}\alpha_{\rm s}({\bm{\rho}})$ are Lipschitz continuous over $\Omega$, and the same applies to $\nabla_{\bm{\rho}}L({\bm{\rho}})$ and $\nabla_{\bm{\rho}}L_{\rm s}({\bm{\rho}})$.
-Additionally, since ${\bm{\Theta}}$ is a constant matrix with $\|{\bm{\Theta}}\|_{2}=1$, and $\nabla_{\bm{\rho}}f({\bm{\rho}})$ is Lipschitz continuous, the composition $\nabla\mathcal{L}({\bm{\varpi}})={\bm{\Theta}}\nabla L({\bm{\rho}})$ is also Lipschitz continuous.
-In other words, there exists a constant $C>0$ such that
+Finally, since $\bm{\Theta}$} is a constant matrix with $\|{\bm{\Theta}}\|_{2}=1$, and $\nabla_{\bm{\rho}}f({\bm{\rho}})$ is Lipschitz continuous, the composition $\nabla\mathcal{L}({\bm{\varpi}})={\bm{\Theta}}\nabla L({\bm{\rho}})$ is also Lipschitz continuous.
+In other words, there exists a constant $C>0$, independent of $\bm{\varpi}$}, such that
 $\small\|\nabla_{{\bm{\varpi}}}\mathcal{L}({\bm{\varpi}}_{1})-\nabla_{{\bm{\varpi}}}\mathcal{L}({\bm{\varpi}}_{2})\|_{2}\leq C\|{\bm{\varpi}}_{1}-{\bm{\varpi}}_{2}\|_{2},\forall{\bm{\varpi}}_{1},{\bm{\varpi}}_{2}\in\Omega_{\bm{\varpi}}$
@@
 Therefore, a more meaningful comparison of the overall computational cost focuses on the number of ILS steps performed per iteration.
-This number is influenced by the step size threshold, given by the right-hand side of (C), the initial step size $\upsilon_{0,i}$ in (48), and the scaling factor.
-At a given iterate ${\bm{\rho}}_{i}$, a larger search direction norm $\|{\bm{d}}_{i}\|$ generally leads to a smaller $\upsilon_{0,i}$ and a reduced step size threshold.
+This number is influenced by the step size threshold, given by the right-hand side of (C), the initial step size $\upsilon_{i,0}$ in (48), and the scaling factor.
+At a given iterate ${\bm{\rho}}_{i}$, a larger search direction norm $\|{\bm{d}}_{i}\|$ generally leads to a smaller $\upsilon_{i,0}$ and a reduced step size threshold.
 As a consequence, the number of ILS steps per iteration is expected to remain comparable between the PP-MCG-ILS and PP-MSD-ILS algorithms.
-Accordingly, the primary metrics for evaluating the relative performance of these two algorithms are their convergence rate and the quality of their steady-state solutions.
-We now qualitatively compare the computational complexity between algorithms employing ILS and those without ILS.
-As previously discussed, the number of ILS steps significantly impacts the number of evaluations of the objective function.
-Thus, the computational complexity for step size selection of the benchmark algorithms can be considerably lower at each iteration compared to the PP-MCG-ILS and PP-MSD-ILS algorithms.
-However, all four algorithms have similar complexity in direction update due to the inherently intensive gradient calculations involved.
-Moreover, the overall computational complexity heavily depends on the convergence rates, or equivalently, on the total number of iterations required for convergence.
-As will be demonstrated in the subsequent simulation results, the proposed PP-MCG-ILS algorithm exhibits a notably higher convergence rate compared to its counterparts, with or without ILS.
-This superior convergence performance significantly reduces the overall computational effort, particularly in the sense that it benefits demanding gradient computations.
+Therefore, their relative performance is primarily characterized by their convergence rates and the quality of the converged solutions.
+We next compare algorithms with and without ILS.
+As discussed above, the number of ILS steps directly affects the number of objective-function evaluations.
+Accordingly, benchmark algorithms without ILS incur a lower step size selection cost per iteration compared with the PP-MCG-ILS and PP-MSD-ILS algorithms.
+However, all four algorithms exhibit similar computational complexity in the direction update, since gradient evaluations dominate this stage.
+Finally, the overall computational complexity is strongly influenced by the convergence rate, or equivalently, by the total number of iterations required to reach convergence.
+As demonstrated in the simulation results, the proposed PP-MCG-ILS algorithm converges significantly faster than the benchmark methods, both with and without ILS.
+This faster convergence substantially reduces the overall computational effort, particularly by alleviating the repeated evaluation of computationally intensive gradients.
 VII Simulation Results
@@
 We employ the orthogonal chirp division multiplexing (OCDM) [31] waveform with a single Gaussian pulse shaping.
-The carrier frequency is $3$GHz.
+The carrier frequency is 33GHz.
 The lowpass equivalent is formulated by
@@
 VII-A CRLB Validation
-We first verify the CRLB in comparison with the MLE results.
+To verify the approximate CRLB, we first compare it with the mean squared error (MSE) of MLE results across the considered SCNR range.
 To maintain consistency with existing works [16, 17, 3], we use the signal energy-to-noise ratio (SENR), defined as $\tilde{\delta}_{\rm cn}\triangleq P/\sigma_{\rm cn}^{2}$, instead of the SCNR $\delta_{\rm cn}$ in evaluating the CRLBs.
@@
 Fig.
-3 compares the mean squared error (MSE) of the MLE with the derived CRLB.
-It is observed that the derived CRLBs closely match the MSEs for SENR values above $-20$ dB.
+3 compares the MSE of the MLE with the derived CRLB.
+It is observed that the derived CRLBs closely match the MSEs for SENR values above 20-20 dB.
 The small discrepancies in the low-SENR regime are attributed to the suboptimal behavior of the likelihood function under noise-dominated conditions.
@@
 To ensure communication quality, a larger number of APs are designated for transmitting ISAC signals, while a smaller subset is used for receiving target-reflected signals.
-Without loss of generality, we consider a CF MIMO-ISAC system comprising 10 transmitters and 2 receivers, as illustrated in Fig.
+Without loss of generality, we consider a CF MIMO-ISAC system comprising 10 transmitters and 2 receivers in Fig.
 2.
@@
 The PA budget is ${\bm{\rho}}_{\max}=0.3\times{\bm{1}}_{N}$.
-The SENR is set to be $-10$ dB.
+The SENR is set to be 10-10 dB.
 Within the coherence time for RCS, a typical modulus-squared RCS matrix is given by $\begin{bmatrix}0.37&0.70&1.38&0.65&2.40&0.18&0.25&0.82&0.14&0.35\\ 0.40\!&\!2.05&1.14&0.42&0.03&0.11&2.02&1.65&1.24&3.65\end{bmatrix}^{\rm T}$.
@@
 Figure 3: The MSE versus the CRLB under different SENRs.
-The PA vector is initialized as the uniform PA vector ${\bm{\rho}_{1}}={\bm{1}}_{N}/N$, unless noted otherwise.
+The PA vector is initialized as the uniform PA vector ${\bm{\rho}_{0}}={\bm{1}}_{N}/N$, unless noted otherwise.
 The initial penalty factor is set to be $\mu_{1}=10^{4}$.
@@
 The other simulation conditions are given in Table I.
+TABLE I: Simulation parameters
+  Algorithms $P$-MCG-ILS $P$-MSD-ILS $P$-NCG $P$-NSD
+  Fig. 4 $\{\epsilon_{\rm th},\sigma_{\upsilon}\}=\{1e-11,0.5\}$ same as PP-MCG-ILS $\upsilon\in\{4e-5,2e-5\}$ same as PP-NCG
+Figs. 5 and 6 same as Fig. 4 $\upsilon=2e-5$
+Figs. 7 and 8 $\epsilon_{\rm th}=1e-12$, $\sigma_{\upsilon}=0.7$ same as PP-MCG-ILS $\upsilon=2e-5$ same as PP-NCG
+Figs. 9 and 10 $\{\epsilon_{\rm th},\sigma_{\upsilon}\}=\{1e-11,0.7\}$
+Figure 4: SINR convergence behavior.
+Figure 5: PA: (a) $P$-MCG-ILS; (b) $P$-NCG with $\upsilon=2\times 10^{-5}$.
+Figure 6: CRLBs: (a) $P$-MCG-ILS; (b) $P$-NCG with $\upsilon=2\times 10^{-5}$.
 We begin by evaluating the convergence and steady-state performance of the proposed algorithms.
@@
 Furthermore, the benchmarks without ILS display noticeable numerical fluctuations, particularly under large fixed step sizes, suggesting potential instability and the risk of divergence.
-To ensure convergence reliability, we refrain from further increasing the step sizes in these simulations.
-[…]
-TABLE I: Simulation parameters
-  Algorithms PP-MCG-ILS PP-MSD-ILS PP-NCG PP-NSD
-  Fig. 4 $\{\epsilon_{\rm th},\sigma_{\upsilon}\}=\{1e-11,0.5\}$ same as PP-MCG-ILS $\upsilon\in\{4e-5,2e-5\}$ same as PP-NCG
-Figs. 5 and 6 same as Fig. 4 $\upsilon=2e-5$
-Figs. 7 and 8 $\epsilon_{\rm th}=1e-12$, $\sigma_{\upsilon}=0.7$ same as PP-MCG-ILS $\upsilon=2e-5$ same as PP-NCG
-Figs. 9 and 10 $\{\epsilon_{\rm th},\sigma_{\upsilon}\}=\{1e-11,0.7\}$
-Figure 4: SINR convergence behavior.
-Figure 5: PA coefficient: (a) PP-MCG-ILS; (b) PP-NCG with $\upsilon=2\times 10^{-5}$.
-Figure 6: CRLB values: (a) PP-MCG-ILS; (b) PP-NCG with $\upsilon=2\times 10^{-5}$.
-The sequential penalty-based algorithms above leverage the optimal solution from the previous iteration loop as the initial point for the current loop under an updated penalty factor.
-To further explore the impact of initialization, we consider fixing the penalty factor and comparing the ISAC performance under different PA initialization strategies.
-In addition to the commonly used uniform PA initialization, we introduce a heuristic PA scheme, where the initial allocation is proportional to the communication channel gain, given by ${\bm{\rho}}_{1}={\bm{g}}/\sum_{n}|g_{n}|^{2}$.
-The results are illustrated in Figs.
-7 and 8.
-When comparing to sequential penalty-based counterparts in Fig.
-4, the PP-MCG-ILS and PP-NCG algorithms initialized with uniform PA exhibit similar convergence behavior.
-However, the PP-MSD-ILS and PP-NSD algorithms show degraded convergence, highlighting the superior convergence properties of CG-based algorithms over SD-based methods under identical conditions.
-When employing the heuristic PA initialization, all algorithms demonstrate significantly better convergence rates compared to the uniform PA case.
-Moreover, both PP-NCG and PP-NSD achieve noticeable performance gains relative to their sequential penalty-based counterparts, particularly PP-NCG.
-In contrast, the PP-MCG-ILS algorithm shows robust and superior convergence behavior, and PP-MSD-ILS exhibits a slight degradation, relative to their sequential penalty-based counterparts.
+Therefore, we refrain from further increasing the step sizes in these simulations.
+The evolutions of the PA coefficients and the corresponding CRLB values over iterations are illustrated in Fig.
+5 and Fig.
+6, respectively.
+Overall, the results demonstrate that the PP-MCG-ILS algorithm achieves significantly faster convergence than the PP-NCG algorithm.
+As shown in Fig.
+5, all algorithms eventually converge to the same PA vector, given by ${\bm{\rho}}=[0.01,\ 0.30,\ 0.0638,\ 0.01,\ \dots,\ 0.01,\ 0.30,\ 0.2762]^{\rm T}$, which satisfies the power constraints.
+Notably, transmitters 2, 9, and 11 are allocated more power due to their superior communication channel gains.
+Furthermore, Fig.
+6 dipicts the exploration process of the CRLB values under threshold-based constraints, in particular for the location CRLB, thereby validating the effectiveness of the proposed optimization framework.
+The sequential penalty-based algorithms use the optimal solution from the previous loop as the initialization for the next loop with an updated penalty factor.
+To further examine the impact of initialization, we fix the penalty factor and compare the ISAC performance under different PA initialization strategies.
+In addition to uniform initialization, a heuristic scheme is considered, where the initial PA is set proportional to the communication channel gain, i.e., ${\bm{\rho}}_{0}={\bm{g}}/\sum_{n}|g_{n}|^{2}$.
+The results in Figs.
+7 and 8 show that a smaller penalty factor ($\mu=10^{5}$) yields faster convergence than $\mu=10^{6}$, especially for the benchmark algorithms.
+Moreover, heuristic PA initialization significantly accelerates convergence for all methods compared to uniform initialization.
+When comparing the fixed-penalty schemes with uniform initialization (Fig.
+7) to the sequential penalty-based schemes (Fig.
+4), the PP-MCG-ILS and PP-NCG algorithms exhibit similar convergence behavior, whereas the PP-MSD-ILS and PP-NSD algorithms converge more slowly.
+With heuristic initialization (Fig.
+8), PP-NCG and PP-NSD achieve noticeable performance gains, while PP-MCG-ILS maintains robust and superior convergence and PP-MSD-ILS shows a slight degradation.
 Figure 7: SINR with initial uniform PA
 Figure 8: SINR with initial heuristic PA
-Figure 9: Steady-state SINRs and CRLBs versus the location CRLB thresholds: (a) SINR; (b) CRLB.
-Figure 10: Steady-state SINRs and CRLBs versus the velocity CRLB thresholds: (a) SINR; (b) CRLB.
-To summarize, the proposed algorithms exhibit superior convergence behavior under the sequential penalty framework, consistently outperforming the benchmark methods.
-While the benchmark algorithms can benefit from specific conditions, such as an empirically tuned penalty factor and a well-initialized starting point, their performance is less robust and generalizable.
-In contrast, the PP-MCG-ILS algorithm demonstrates consistently strong convergence and steady-state accuracy across a wide range of settings, irrespective of initialization or penalty factor selection.
+To summarize, while the benchmark algorithms can benefit from specific conditions, such as an empirically tuned penalty factor and a well-initialized starting point, their performance is less robust and generalizable.
+In contrast, the PP-MCG-ILS algorithm exhibits superior convergence behavior, consistently outperforming the benchmark methods, irrespective of initialization or penalty factor selection.
 We now investigate the impact of the CRLB thresholds on the performance of the proposed PP-MCG-ILS algorithm.
 As shown in Fig.
-9, relaxing the location CRLB thresholds leads to a gradual increase in the steady-state SINR, which eventually stabilizes.
-This trend aligns with the behavior observed in Fig.
-9, where the achieved location CRLB value closely approaches its threshold and then converge.
-These trends reveal the trade-off between the communication SINR and the location CRLBs.
-Furthermore, it is noteworthy that the velocity CRLB remains largely unaffected by variations in the location CRLB thresholds.
-In contrast, Fig.
-10 shows that the performance of the PP-MCG-ILS algorithm exhibits negligible variation with the increased velocity CRLB threshold.
-The observed CRLB saturation with respect to the corresponding CRLB threshold may be an inherent outcome of the introduced power budget constraints.
+9, relaxing the location CRLB threshold results in a gradual increase in the steady-state SINR, which eventually saturates.
+This behavior is consistent with the trend observed in Fig.
+9, where the achieved location CRLB closely approaches its prescribed threshold and then converges.
+These results illustrate the inherent trade-off between communication SINR and location CRLB.
+Figure 9: Steady-state SINRs and CRLBs: (a) SINR; (b) CRLB.
+Figure 10: Steady-state SINRs and CRLBs: (a) SINR; (b) CRLB.
+However, the velocity CRLB in Fig.
+9 is largely insensitive to variations in the location CRLB threshold.
+Similarly, Fig.
+10 shows that changing the velocity CRLB threshold has a negligible impact on the performance of the PP-MCG-ILS algorithm.
+The observed saturation and difference of the CRLB with respect to its threshold is likely a consequence of the nonlinear structure of the CRLB formulation and the compact feasible region imposed by the power constraints.
 Figure 11: Learning results for pure sensing: (a) total power; (b) CRLBs.
 VII-C Power Allocation for Pure Sensing
-We now verify the effectiveness of the proposed P-MCG-ILS algorithm for total power minimization in pure sensing scenario with joint location and velocity estimation.
+We now evaluate the proposed P-MCG-ILS algorithm on total power minimization in pure sensing scenario.
 The initial penalty factor is set to be $\mu_{1}=1$, $\sigma_{\upsilon}=0.2$ and the penalty threshold $\epsilon_{\mu}=10^{-6}$.
@@
 In contrast, in the ISAC scenario (refer to Fig.
-6), the velocity CRLB matches its threshold loosely.
+6 and Fig.
+10), the velocity CRLB matches its threshold loosely.
 This difference arises because, in pure sensing, any enhancement in CRLB performance directly benefits from increased sensing power.
@@
 VIII Conclusions
-This paper presents, for the first time, a comprehensive investigation of the power allocation problem in a CF MIMO-ISAC framework, where joint location and velocity estimation were considered as the sensing task.
-We employed newly derived approximate CRLBs as sensing performance metrics, alongside user receive SINR as the communication metric.
-To address the inherent nonlinearity and non-convexity of the resulting optimization problem, we propose the PP-MCG-ILS algorithm.
-Additionally, the P-NCG-ILS algorithm was further proposed to handle the power minimization problem in pure sensing scenarios.
-We devoted into the detailed analysis of the convergence behavior and qualitatively compare the computational complexity of all proposed methods.
-Simulation results confirmed the accuracy of the derived CRLBs and demonstrate the efficacy of the proposed power allocation strategies in enhancing pure sensing and overall ISAC performance.
-Our findings offer valuable insights into resource-efficient design for next-generation CF MIMO-ISAC networks and lay the groundwork for future extensions involving more complex sensing tasks and dynamic resource management.
+This paper presented a comprehensive investigation of the PA problem in a CF MIMO-ISAC framework, where joint location and velocity estimation were considered as sensing tasks.
+Newly derived approximate CRLBs were employed as sensing performance metrics, together with the user receive SINR as the communication metric.
+To address the inherent nonlinearity and nonconvexity of the resulting optimization problem, the PP-MCG-ILS algorithm was proposed.
+In addition, the P-NCG-ILS algorithm was developed to handle the power minimization problem in pure sensing scenarios.
+We further conducted a detailed convergence analysis and provided a qualitative comparison of the computational complexity of the proposed methods.
+Simulation results confirmed the accuracy of the derived CRLBs and demonstrated the effectiveness of the proposed PA strategies in enhancing both pure sensing and overall ISAC performance.
+These findings provide valuable insights into resource-efficient design for next-generation CF MIMO-ISAC networks and lay the groundwork for future extensions involving more complex sensing tasks and dynamic resource management.
 Appendix A Relevant parameters of the derived CRLBs
@@
 where $S$ denotes the number of samples.
-In addition, the remaining parameters, including $\beta_{n,k}$, $\zeta_{n,k}$, $\xi_{n,k}$, $\varrho_{n,k}$, $\eta_{n,k}$, and $\kappa_{n,k}$, represent geometric spreading factors, given respectively by
-$[\beta_{n,k},\zeta_{n,k}]^{\rm T}=$ $\dfrac{|\alpha_{n,k}|}{c}\left(\dfrac{{\bm{l}}-{\bm{l}}_{n}}{\|{\bm{l}}_{n}-{\bm{l}}\|_{2}}+\dfrac{{\bm{l}}-{\bm{l}}_{k}}{\|{\bm{l}}_{k}-{\bm{l}}\|_{2}}\right),$ (62a)
-$[\xi_{n,k},\varrho_{n,k}]^{\rm T}=$ $\dfrac{f_{c}|\alpha_{n,k}|}{c}\left(\dfrac{{\bm{l}}_{n}-{\bm{l}}}{\|{\bm{l}}_{n}-{\bm{l}}\|_{2}}+\dfrac{{\bm{l}}_{k}-{\bm{l}}}{\|{\bm{l}}_{k}-{\bm{l}}\|_{2}}\right),$ (62b)
-$[\eta_{n,k},\kappa_{n,k}]^{\rm T}=$ $\dfrac{f_{c}|\alpha_{n,k}|}{c}\bigg{(}({\bm{l}}_{n}-{\bm{l}})\dfrac{{\bm{v}}^{\rm T}({\bm{l}}_{n}-{\bm{l}})}{\|{\bm{l}}_{n}-{\bm{l}}\|_{2}^{3}}-\dfrac{{\bm{v}}}{\|{\bm{l}}_{n}-{\bm{l}}\|_{2}}$
+The remaining parameters, including $\beta_{n,k}$, $\zeta_{n,k}$, $\xi_{n,k}$, $\varrho_{n,k}$, $\eta_{n,k}$, and $\kappa_{n,k}$, represent geometric spreading factors, given respectively by
+$[\beta_{n,k},\zeta_{n,k}]^{\rm T}=\$$\dfrac{|\alpha_{n,k}|}{c}\left(\dfrac{{\bm{l}}-{\bm{l}}_{n}}{\|{\bm{l}}_{n}-{\bm{l}}\|_{2}}+\dfrac{{\bm{l}}-{\bm{l}}_{k}}{\|{\bm{l}}_{k}-{\bm{l}}\|_{2}}\right),$ (62a)
+$[\xi_{n,k},\varrho_{n,k}]^{\rm T}=\$$\dfrac{f_{c}|\alpha_{n,k}|}{c}\left(\dfrac{{\bm{l}}_{n}-{\bm{l}}}{\|{\bm{l}}_{n}-{\bm{l}}\|_{2}}+\dfrac{{\bm{l}}_{k}-{\bm{l}}}{\|{\bm{l}}_{k}-{\bm{l}}\|_{2}}\right),$ (62b)
+$[\eta_{n,k},\kappa_{n,k}]^{\rm T}=\$$\dfrac{f_{c}|\alpha_{n,k}|}{c}\bigg(({\bm{l}}_{n}-{\bm{l}})\dfrac{{\bm{v}}^{\rm T}({\bm{l}}_{n}-{\bm{l}})}{\|{\bm{l}}_{n}-{\bm{l}}\|_{2}^{3}}-\dfrac{{\bm{v}}}{\|{\bm{l}}_{n}-{\bm{l}}\|_{2}}$
 $+({\bm{l}}_{k}-{\bm{l}})\dfrac{{\bm{v}}^{\rm T}({\bm{l}}_{k}-{\bm{l}})}{\|{\bm{l}}_{k}-{\bm{l}}\|_{2}^{3}}-\dfrac{{\bm{v}}}{\|{\bm{l}}_{k}-{\bm{l}}\|_{2}}\bigg).$ (62c)
@@
 Appendix C Proof of Theorem 1
-In light of (42), (44) and (63), the ILS condition requires
+From (42), (63), and (44), the ILS condition yields
 $\upsilon_{i}{\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i})+\upsilon_{i}^{2}/2{\bm{d}}_{i}^{\rm T}\nabla^{2}L({\bm{\rho}}_{i}){\bm{d}}_{i}\leq\epsilon_{L}\upsilon_{i}{\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i})$
@@
 where positive ${\bm{d}}_{i}^{\rm T}\nabla^{2}L({\bm{\rho}}_{i}){\bm{d}}_{i}$ is considered.
-Assuming the Lipschitz gradient continuity of $L({\bm{\rho}})$, $\|\nabla^{2}L({\bm{\rho}}_{i})\|$ is upper bounded.
-This indicates a positive and finite ${\bm{d}}_{i}^{\rm T}\nabla^{2}L({\bm{\rho}}_{i}){\bm{d}}_{i}$ and thus non-zero $\upsilon_{i}$.
+With the Lipschitz gradient continuity of $L({\bm{\rho}})$, $\|\nabla^{2}L({\bm{\rho}}_{i})\|$ is upper bounded.
+This indicates a positive and finite ${\bm{d}}_{i}^{\rm T}\nabla^{2}L({\bm{\rho}}_{i}){\bm{d}}_{i}$ and thus nonzero and bounded $\upsilon_{i}$.
 Furthermore, in light of (44), we have
 $\small\Delta^{2}L({\bm{\rho}}_{i+1})\leq\epsilon_{L}^{2}\upsilon_{i}^{2}({\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i}))^{2}.$ (65)
-Since $L({\bm{\rho}})$ is a monotonously decreasing and bounded function regarding ${\bm{\rho}}_{i}$, the (conjugate) gradient-based descent algorithm converges to $\Delta^{2}L({\bm{\rho}}_{i+1})\rightarrow 0$ when $i>I$ for a large constant $I$.
-However, the deflection factor computed by (41) or the descent direction vector in (49) can guarantee a strict descent ${\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i})<0$ at each iteration $i$ unless ${\bm{d}}_{i}={\bm{0}}$.
+Since $L({\bm{\rho}})$ is monotonously decreasing and bounded, we have $\Delta^{2}L({\bm{\rho}}_{i+1})\rightarrow 0$ when $i>I$ for a large constant $I$.
+This implies ${\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i})\rightarrow 0$.
+However, a strict descent is satisfied with ${\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i})<0$ at each iteration $i$ unless ${\bm{d}}_{i}={\bm{0}}$.
 Consequently, sequence $\{L({\bm{\rho}}_{i})\}$ will converge to a stationary point with a zero descent direction ${\bm{d}}_{i}\rightarrow{\bm{0}}$ when it has Lipschitz gradient continuity.
-According to Lemma 1, we have $\|\Delta{\bm{\rho}}_{i+1}\|\rightarrow 0$ when ${\bm{d}}_{i}\rightarrow{\bm{0}}$.
-On the other hand, we assume the objective function is not Lipschitz gradient continuous.
-In this case, if the sequence $\{{\bm{\rho}}_{i}\}$ pass a Lipschitz noncontinuous point ${\bm{\rho}}_{I}$ where ${\bm{d}}_{I}^{\rm T}\nabla^{2}L({\bm{\rho}}_{I}){\bm{d}}_{I}$ is extremely large, to satisfy (C), the step size should meet $\upsilon_{I}\rightarrow 0$.
-According to Lemma 1, we have $\|\Delta{\bm{\rho}}_{i+1}\|\rightarrow 0$ and $\Delta L({\bm{\rho}}_{i+1})\rightarrow 0$ when $\upsilon_{i}\rightarrow 0$ for $i\geq I$, i.e., the sequences $\{{\bm{\rho}}_{i}\}$ and $\{L({\bm{\rho}}_{i})\}$ converge to a Lipschitz gradient noncontinuous point.
+According to Lemma 1, we further have $\|\Delta{\bm{\rho}}_{i+1}\|\rightarrow 0$ when ${\bm{d}}_{i}\rightarrow{\bm{0}}$.
 $\blacksquare$